### v1

# Discrete Fourier Transform versus Discrete Chi-Square Method

Abstract
We compare two time series analysis methods, the Discrete Fourier Transform (DFT) and our Discrete Chi-square Method (DCM). DCM is designed for detecting many signals superimposed on an unknown trend. The solution for the non-linear DCM model is an ill-posed problem. The backbone of DCM is the Gauss-Markov theorem that the least squares fit is the best unbiased estimator for linear regression models. DCM is a simple numerical time series analysis method that performs a massive number of linear least squares fits. We show that our numerical solution for the DCM model fulfills the three conditions of a well-posed problem: existence, uniqueness and stability. The Fisher-test is used to identify the best DCM model from all alternative tested DCM models. The correct DCM model must also pass our Predictivity-test. Our analyses of seven different simulated data samples expose the weaknesses of DFT and the efficiency of DCM. The DCM signal and trend detection depend only on the sample size and the accuracy of data. DCM is an ideal forecasting method because the time span of observations is irrelevant. We recommend fast sampling of large high quality datasets and the analysis of those datasets using numerical DCM parallel computation Python code.
keywords:
Ill-posed problems, Time series analysis, Non-linear models, Discrete Chi-square Method (DCM), Discrete Fourier Transform (DFT)
1 Introduction
[9] defined the three conditions for a well-posed problem. The solution for the problem determines these conditions.
  • $\mathrm{C}_{1}$ Existence: A solution exists.
  • $\mathrm{C}_{2}$ Uniqueness: The solution is unique.
  • $\mathrm{C}_{3}$ Stability: Small changes in the input data cause small changes in the solution.
Stability means that the solution behaves predictably as the input varies. In other words, there exists a continuous mapping from the input space to the solution space.
If any of these conditions are not satisfied, the problem is classified as ill-posed. Ill-posed problems are encountered in various fields of science [22]. For example, Piskunov et al. [24] solves the stellar surface imaging inverse problem using the [29] regularization technique. Ill-posed problems are also encountered in time series analysis of astronomical data [30, 1]. The problem is ill-posed because the unknown signal frequencies are free parameters of non-linear time series analysis models. The Lomb-Scargle periodogram is the most frequently applied time series analysis method in astronomy because the observations are often unevenly spaced in time [21, 27]. Several extentions of this method have been published [35, 32]. The DFT pre-whitening technique can detect many signals from unevenly spaced data [e.g., 26, 34]. The data must be detrended before applying this technique, which detects one frequency at a time, until it detects no new frequency. In this study, artificial simulated data having equal weights is analysed using the DFT version published by Horne and Baliunas [11]. We compare the performance of this DFT version to the performance of our DCM.
[19] formulated the least squares method. [7] connected this method to the principles of probability and the normal distribution. A few years later, he showed that the least-squares method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [8]. The extended Gauss-Markov theorem states that the least squares method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [33]. This Gauss-Markov theorem is the backbone of our DCM.
We present the DCM formulation in Section 2.1. Data samples are simulated using the DCM model (Sect. 2.2). The four DFT limitations are discussed in Section 2.3 (Equations 26-29). DCM and DFT time series analysis of the simulated data samples shows that DCM does not suffer from the DFT limitations (Sects. 3.1-3.7). We then present the identification of the best DCM model for the data (Sect. 3.8), the DCM significance estimates (Sect. 3.9) and the DCM predictions (Sect. 3.10). Finally, we discuss our results (Sect. 4).
2 Methods
The observations are $y_{i}=y(t_{i})\pm\sigma(t_{i})$, where $t_{i}$ are the observing times and $\sigma_{i}$ are the errors (i=1,2,,n)(i=1,2,...,n). The time span of data is $\Delta T=t_{n}-t_{1}$. The mid point is $t_{\mathrm{mid}}=t_{1}+\Delta T/2$. The mean and the standard deviation of all $y_{i}$ values are denoted with $m$ and $s$.
2.1 Discrete Chi-square method (DCM)
[13] formulated DCM. The model of this method is
$g(t)=g(t,K_{1},K_{2},K_{3})=h(t)+p(t),$ (1)
where the integer values $K_{1}$, $K_{2}$ and $K_{3}$ are called the model orders [14]. The notation $g_{K_{1},K_{2},K_{3}}(t)$ is used to specify these orders. The DCM model is a sum of two functions. These functions are the periodic function
$h(t)$ $=$ $h(t,K_{1},K_{2})=\sum_{i=1}^{K_{1}}h_{i}(t,f_{i})$ (2)
$h_{i}(t,f_{i})$ $=$ $\sum_{j=1}^{K_{2}}B_{i,j}\cos{(2\pi jf_{i}t)}$
$+$ $C_{i,j}\sin{(2\pi jf_{i}t)}$
and the aperiodic function
[…]
where
$p_{k}(t)$ $=$ $M_{k}\left[{{2(t-t_{\mathrm{mid}}})\over{\Delta T}}\right]^{k}.$ (5)
For $k>1$, the $p_{k}(t)$ function full range is
$\begin{cases}2|M_{k}|,&\text{if $k$ is odd}\\ |M_{k}|,&\text{if $k$ is even.}\end{cases}$ (6)
The free parameters of $g(t)$ model are
$\bm{\beta}$ $=$ $[\beta_{1},\beta_{2},...,\beta_{\eta}]$
$=$ $[B_{1,1},C_{1,1},f_{1},...,B_{K_{1},K_{2}},C_{K_{1},K_{2}},f_{K_{1}},M_{0},...,M_{K_{3}}].$
The number of free parameters is
$\eta=K_{1}\times(2K_{2}+1)+K_{3}+1.$ (8)
We divide the free parameters $\bm{\beta}$ into two groups $\bm{\beta}_{I}$ and $\bm{\beta}_{II}$. The first group are the frequencies
$\bm{\beta}_{I}$ $=$ $[f_{1},...,f_{K_{1}}].$ (9)
Due to this group, all free parameters are not eliminated from all partial derivatives $\partial g/\partial\beta_{i}$. This makes the $g(t)$ model non-linear. If the $\bm{\beta}_{I}$ frequencies have constant values, the multipliers $2\pi jf_{i}$ in Equation 2.1 become constants. In this case, the model becomes linear because the partial derivatives $\partial g/\partial\beta_{i}$ no longer contain any free parameters. The solution for the second group of remaining free parameters
$\bm{\beta}_{II}$ $=$ $[B_{1,1}C_{1,1},...,B_{K_{1},K_{2}},C_{K_{1},K_{2}},M_{0},...,M_{K_{3}}]$ (10)
becomes unique. This solution passes the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem.
Let us assume that we search for periods between $f_{\mathrm{min}}$ and $f_{\mathrm{max}}$. The non-linear $g(t)$ model becomes linear, if the tested $\bm{\beta}_{I}$ frequencies are fixed to any constant values. The sum $h(t)$ of signals $h_{1}(t,f_{1})$, $h_{2}(t,f_{2})$ … and $h_{K_{1},f_{K_{1}}}(t)$ does not depend on the order in which these signals are added. For example, the two signal $K_{1}=2$ model symmetry is $h(t)=h_{1}(t,f_{1})+h_{2}(t,f_{2})=h_{1}(t,f_{2})+h_{2}(t,f_{1})$. Both $(f_{1},f_{2})$ and $(f_{2},f_{1})$ combinations give the same value for the $g(t)$ model. Since this symmetry applies to any $K_{1}$ number of signals, we compute the linear $g(t)$ models only for all tested frequency combinations
$f_{\mathrm{max}}\geq f_{1}>f_{2}>...>f_{K_{1}}\geq f_{\mathrm{min}}.$ (11)
The DCM model residuals
$\epsilon_{i}=y(t_{i})-g(t_{i})=y_{i}-g_{i}.$ (12)
give the sum of squared residuals
$R=\sum_{i=1}^{n}\epsilon_{i}^{2},$ (13)
and the Chi-square
$\chi^{2}=\sum_{i=1}^{n}{{\epsilon_{i}^{2}}\over{\sigma_{i}^{2}}}.$ (14)
For every tested $\bm{\beta}_{I}=[f_{1},f_{2},...,f_{K_{1}}]$ frequency combination, the DCM test statistic
$z$ $=$ $z(f_{1},f_{2},...,f_{K_{1}})=\sqrt{R/n}$ (15)
$z$ $=$ $z(f_{1},f_{2},...,f_{K_{1}})=\sqrt{\chi^{2}/n}$ (16)
is computed from a linear model least squares fit. Equation 15 or Equation 16 is applied for unknown or known $\sigma_{i}$ errors, respectively. The value of $z$ is unique.
In the preliminary long search, we test an evenly spaced grid of $n_{\mathrm{L}}$ frequencies between $f_{\mathrm{min}}$ and $f_{\mathrm{max}}$. This search gives the best frequency candidates $f_{\mathrm{1,mid}},...f_{\mathrm{K_{1},mid}}$.
In the final short search, we test a denser grid of $n_{\mathrm{S}}$ frequencies within an interval
$[f_{\mathrm{i,mid}}-a,f_{\mathrm{i,mid}}+a],$ (17)
where $i=1,..,K_{1}$, $a=c(f_{\mathrm{min}}-f_{\mathrm{max}})/2$ and $c=0.05$.
The total number of all tested long and short search frequency combinations is
$\binom{n_{f}}{K_{1}}={{n_{f}!}\over{K_{1}!(n_{f}-K_{1})!}},$ (18)
where $n_{f}=n_{L}$ and $n_{f}=n_{S}$, respectively.
The global periodogram minimum
$z_{\mathrm{min}}=z(f_{\mathrm{1,best}},f_{\mathrm{2,best}},...,f_{\mathrm{K_{1},best}})$ (19)
is at the tested frequencies $f_{\mathrm{1,best}},$ $f_{\mathrm{2,best}},$..., $f_{\mathrm{K_{1},best}}$. This tested frequency combination gives the best linear model for the data. The periodogam value $z$ is a scalar, which is computed from $K_{1}$ frequency values. It is possible to plot the $K_{1}=2$ two signal periodogram $z(f_{1},f_{2})$ as a map, where $f_{1}$ and $f_{2}$ are the coordinates, and $z=z(f_{1},f_{2})$ is the height. For more than two signals, there is no direct graphical $z$ plot because that requires more than three dimensions. Our solution for this dimensional problem is simple. We plot only the following one-dimensional slices of the full periodogram
$z_{1}(f_{1})$ $=$ $z(f_{1},f_{\mathrm{2,best}},...,f_{\mathrm{K_{1},best}})$ (20)
$z_{2}(f_{2})$ $=$ $z(f_{\mathrm{1,best}},f_{2},f_{\mathrm{3,best}},...,f_{\mathrm{K_{1},best}})$
$z_{3}(f_{3})$ $=$ $z(f_{\mathrm{1,best}},f_{\mathrm{2,best}},f_{3},f_{\mathrm{4,best}},...,f_{\mathrm{K_{1},best}})$~
$z_{4}(f_{4})$ $=$ $z(f_{\mathrm{1,best}},f_{\mathrm{2,best}},f_{\mathrm{3,best}},f_{4},f_{\mathrm{5,best}},f_{\mathrm{K_{1},best}})$
$z_{5}(f_{5})$ $=$ $z(f_{\mathrm{1,best}},f_{\mathrm{2,best}},f_{\mathrm{3,best}},f_{\mathrm{4,best}},f_{5},f_{\mathrm{K_{1},best}})$
$z_{6}(f_{6})$ $=$ $z(f_{\mathrm{1,best}},f_{\mathrm{2,best}},f_{\mathrm{3,best}},f_{\mathrm{4,best}},f_{\mathrm{5,best}},f_{6}).$
In the above-mentioned $K_{1}=2$ map, the slice $z_{1}(f_{1})$ would represent the height $z$ at the location $(f_{1},f_{\mathrm{2,best}})$ when moving along the straight constant line $f_{2}=f_{\mathrm{2,best}}$ that crosses the global minimum $z_{\mathrm{min}}$ (Equation 19) at the coordinate point $(f_{\mathrm{1,best}},f_{\mathrm{2,best}})$.
The short search gives the best $f_{\mathrm{1,best}},$..., $f_{\mathrm{K_{1},best}}$ frequencies for the data. These frequencies are the unique initial values for the first group of free parameters $\bm{\beta}_{\mathrm{I,initial}}=$ $f_{\mathrm{1,best}},$..., $f_{\mathrm{K_{1},best}}]$ (Equation 9). The linear model for these constant $f_{\mathrm{1,best}},...,f_{\mathrm{K_{1},best}}]$ frequencies gives the unique initial values for the second group $\bm{\beta}_{\mathrm{II,initial}}$ of free parameters (Equation 10). The non-linear iteration
$\bm{\beta}_{\mathrm{initial}}=[\bm{\beta}_{\mathrm{I,initial}},\bm{\beta}_{\mathrm{II,initial}}]\rightarrow\bm{\beta}_{\mathrm{final}}$ (21)
gives the final free parameter values $\bm{\beta}_{\mathrm{final}}$.
DCM determines the signal parameters
  • $P_{i}=1/f_{i}=$ Period
  • $A_{i}=$ Peak to peak amplitude
  • $t_{\mathrm{i,min,1}}=$ Deeper primary minimum
  • $t_{\mathrm{i,min,2}}=$ Secondary minimum (if present)
  • $t_{\mathrm{i,max,1}}=$ Higher primary maximum
  • $t_{\mathrm{i,max,2}}=$ Secondary maximum (if present)
and the trend parameters
  • $M_{k}$ = Polynomial coeffients.
We use the bootstrap technique to estimate the DCM model parameter errors [5]. The tested bootstrap frequencies are the same as in the short search. Our TSPA- and CPS-methods applied the same bootstrap tehnique [15, 20]. We select a random sample $\bm{\epsilon}^{*}$ from the residuals $\bm{\epsilon}$ of the DCM model (Equation 12). Any $\epsilon_{i}$ value can be chosen to the $\bm{\epsilon}^{*}$ sample as many times as the random selection happens to favour it. We create numerous artificial bootstrap data samples
$\bm{y}^{*}=\bm{g}+\bm{\epsilon}^{*}.$ (22)
Each $\bm{y}^{*}$ sample gives one estimate for every DCM model parameter. For each particular model parameter, the bootstrap error estimate is the standard deviation of all estimates obtained from all $\bm{y}^{*}$ samples.
There are, of course, totally wrong DCM models for the data. For example, DCM can be forced search for too few or too many $K_{1}$ signals, or the selected $p(t)$ trend order $K_{3}$ can be wrong, as shown in Figures 5-10 by [13]. Such DCM models are unstable and we denote them with $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”, like in [14] and Jetsu [^1] Jetsu L. “Do planets cause the sunspot cycle?”, submitted to Scientific Reports. August 13th, 2025. (2025). These unstable models have three signatures
  • Intersecting frequencies
  • Dispersing amplitudes
  • Leaking periods
Intersecting frequencies occur when the signal frequencies in the data are very close to each other. We give the following example of how this instability can arise in the two signal model. If the frequency $f_{1}$ approaches the frequency $f_{2}$, both $h_{1}(t)$ and $h_{2}(t)$ signals become essentially one and the same signal. The least squares fit fails because it makes no sense to model the same signal twice.
Dispersing amplitudes instability can occur, if the two signal frequencies are too close to each other. The least squares fit finds a model, where two high amplitude signals nearly cancel out each other. The low amplitude signal, the sum of these two high amplitude signals, fits to the data.
There are DCM models where the detected frequency $f$ is outside the tested frequency interval between $f_{\mathrm{min}}$ and $f_{\mathrm{max}}$. This leaking periods instability may indicate that the chosen tested period range is wrong.
Table 1: Model 1: DCM analysis between $P_{\mathrm{min}}=0.63$ and $P_{\mathrm{max}}=5.70$. (1) Simulated $P_{1}$, $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,max,1}}$ and $M_{0}$ values. (2-4) Detected values for different $n$ and $\mathrm{SN}$} combinations. Two lowest lines specify electronic data files and DCM analysis control files.
(1) (2) (3) (4)
$n=50$ $n=50$ $n=100$
Model 1 ${\mathrm{SN}}=10$ ${\mathrm{SN}}=50$ ${\mathrm{SN}}=100$
$P_{1}=1.9$ $1.58\pm 0.21$ $1.822\pm 0.042$ $1.863\pm 0.017$
$A_{1}=2.0$ $1.64\pm 0.33$ $1.894\pm 0.072$ $1.941\pm 0.025$
$t_{\mathrm{1,min,1}}=1.35$ $1.20\pm 0.11$ $1.312\pm 0.021$ $1.3321\pm 0.0088$
$t_{\mathrm{1,max,1}}=0.40$ $0.4098\pm 0.0058$ $0.4007\pm 0.0013$ $0.40056\pm 0.00078$
$M_{0}=1.0$ $1.20\pm 0.18$ $1.061\pm 0.036$ $1.030\pm 0.014$
Data file Model1n50SN10.dat Model1n50SN50.dat Model1n100SN100.dat
Control file dcmModel1n50SN10.dat dcmModel1n50SN50.dat dcmModel1n100SN100.dat
[…]
Refer to caption Refer to caption
                 (a)               (c)
                (b)               (d)
Refer to caption Refer to caption
                 (e)               (f)
Figure 1: Model 1 (Table 1: $n=50$, ${\mathrm{SN}}=50$ simulation). (a) DCM long search periodogram $z_{1}(f_{1})$ gives best period at 1.843 (diamond). (b) DCM short search periodogram $z_{1}(f_{1})$ gives best period at 1.822 (diamond). (c) DCM model $g(t)$ (black continuous line), DCM trend $p(t)$ (black dashed line) and data $y_{i}$ (black dots). (d) DCM model detrended g(t)p(t)g(t)-p(t) (black continuous line), DCM signal $h_{1}(t)$ (red thick continuous line), detrended data $y(t_{i})-p(t_{i})$ (black dots) and DCM model residuals $y(t_{i})-g(t_{1})$ (blue dots) offset to -0.65 level (blue dotted line). (e) DFT periodogram $z_{\mathrm{DFT}}(f)$ gives best period at 1.190 (Diamond). (f) DFT model $g_{\mathrm{DFT}}(t)$ (black continuous line), DFT trend $p_{\mathrm{DFT}}(t)$ (black dashed line) and data $y_{i}$ (black dots). (g) DFT model detrended $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (black contiuous line), DFT pure sine $s_{\mathrm{DFT}}(t)$ (red thick continuous line), detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and DFT model residuals (blue dots) offset to -1.5 level (blue dotted line).
2.2 Simulated DCM model data samples
We use seven different $g(t)$ models to simulate 21 artificial data samples (Sections 3.1-3.7). The time interval of all simulated samples is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$. The first and last time point values are then modified to $t_{1}^{\star}=0$ and $t_{n}^{\star}=\Delta T$. Hence, the distance between independent frequencies [18, 27, 25] is always
$f_{0}=1/\Delta T=1.$ (23)
The $n^{\star}$ residuals $\epsilon^{\star}(t_{i}^{\star})$ of simulated model are drawn from a random normal distribution $N(0,\sigma,n^{\star})$, where $\sigma$ is the accuracy of simulated data. The simulated data are
$y^{\star}(t_{i}^{\star})$ $=$ $g(t_{i}^{\star})+\epsilon^{\star}(t_{i}^{\star})$
$=$ $h(t_{i}^{\star})+p(t_{i}^{\star})+\epsilon^{\star}(t_{i}^{\star}).$
The peak to peak amplitudes of all simulated signals is $A=2.$ Our definition for the signal to noise ratio is
${\mathrm{SN}}=(A/2)/\sigma.$ (24)
2.3 Dicrete Fourier Transform (DFT)
We search for signals in the simulated data using the DFT version formulated by [11]. This method searches for the best pure sine model for the data. The notations for the DFT model are
$g_{\mathrm{DFT}}(t)=s_{\mathrm{DFT}}(t)+p_{\mathrm{DFT}}(t),$ (25)
where $s_{\mathrm{DFT}}(t)$ is the pure sine signal and $p_{\mathrm{DFT}}(t)$ is the trend. The pure sine signals for the $y_{i}$ data and the $\epsilon_{i}$ residuals are denoted with $s_{y,\mathrm{DFT}}(t)$ and $s_{\epsilon,\mathrm{DFT}}(t)$, respectively. Our DFT analysis may fail for four main reasons.
The data sample is “too short” [27, 28]
$P>\Delta T.$ (26)
The data contain “a trend” [3, 17]
$p_{\mathrm{DFT}}(t)\neq 0.$ (27)
The signal frequencies are “too close” [18, 23].
$|f_{1}-f_{2}|<f_{0}={\Delta T}^{-1}.$ (28)
The signals are not “pure sines” [16, 2].
$K_{2}\neq 1.$ (29)
We search for many pure sine signals using the DFT pre-whitening technique [26, 34]. The $p_{i}=p_{\mathrm{DFT}}(t_{i})$ trend is first removed from the data $y_{i}=y(t_{i})$. In the search for the first signal, we apply the DFT to the detrended $y_{i}-p_{i}$ data. When searching for the second signal, we apply the DTF to the original model residuals $\epsilon_{i}=y_{i}-g_{\mathrm{y,DFT}}(t_{i})$.
3 Results
We apply DCM and DFT to seven different simulated data samples (Sections 3.1 - 3.7). DCM analysis succeeds for all samples. DFT analysis fails for every sample.
3.1 Model 1
Our first simulation model is the one signal model
$g(t)=(A_{1}/2)\cos{\left[{{2\pi(t-t_{\mathrm{1,max,1}})}\over{P_{1}}}\right]}+M_{0}.$ (30)
We give the $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$ and $M_{0}$ values in Table 1 (Column 1). This sample is “too short” (Equation 26) because $P_{1}=1.9\Delta T$. The constant mean level $M_{0}$ is unknown (Equation 27). We perform the DCM and DFT time series analysis between $P_{\mathrm{min}}=P_{1}/3=0.63$ and $P_{\mathrm{max}}=3P_{1}=5.70$.
Model 1 is a DCM model, where $K_{1}=1$, $K_{2}=1$, $K_{3}=0$, $B_{1,1}=$ $(A_{1}/2)\cos{(2\pi f_{1}t_{\mathrm{1,max,1}})}$ and $C_{1,1}=$ $(A_{1}/2)$ $\sin{(2\pi f_{1}t_{\mathrm{1,max,1}})}.$ We give the DCM analysis results for three samples having different $n$ and $\mathrm{SN}$} combinations (Table 1: Columns 2-4). For each sample, this table specifies the electronic data file and the electronic DCM control file. [^2] We publish all our data files and all DCM control files in https://zenodo.org/uploads/17018676 The detected $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$ and $M_{0}$ values are correct and accurate even for the combination $n=50$ and ${\mathrm{SN}}=50$. These values become more accurate when $n$ and $\mathrm{SN}$} increase.
A graphical presentation of DCM analysis results is shown for Model 1 simulated data sample, where $n=50$ and ${\mathrm{SN}}=50$ (Figures 1a-d). The DCM long search $z_{1}(f_{1})$ periodogram minimum is at $P_{1}=1.843$ (Figure 1a: diamond). The DCM short search periodogram $z_{1}(f_{1})$ gives the best period at $P_{1}=1.822$ (Figure 1b: diamond). The continuous black line denoting the DCM model $g(t)$ fits perfectly to the black dots denoting the data $y_{i}$ (Figure 1c). The mean $p(t)=M_{0}=1.061\pm 0.036$ is correct (Figure 1c: dashed black line). The detrended model g(t)p(t)g(t)-p(t) (black continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots) and the pure sine signal $h_{1}(t)$ (red thick continuous line) are shown in Figure 1d. Note that the thick continuous red line stays under the thin continuous black line because $h_{1}(t)=g(t)-p(t)$. The DCM residuals (blue dots) are offset to the level of -0.65 (blue dotted line). These residuals are stable and display no trends.
The DFT detects the wrong period $P_{1}=1.190$ (Figure 1e: Diamond). The DFT mean level estimate $M_{0}=1.527$ is also wrong (Figure 1f: Dashed black line). The black dots denoting the data $y_{i}$ deviate from the continuous black line denoting the DFT model $g_{\mathrm{DFT}}(t)$. The detrended model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (continuous black line), the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and the signal $s_{\mathrm{y,DFT}}(t)$ (continuous thick red line) are shown in Figure 1g. The thin black line covers the thick red line because $s_{\mathrm{y,DFT}}(t)=g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$. The DFT residuals (blue dots) offset to the level of -1.5 (blue dotted line) display obvious trends, especially at the end of analysed sample.
For the simulated data of Model 1, the DCM analysis succeeds, but the DFT analysis fails.
Table 2: Model 2. DCM analysis between $P_{\mathrm{min}}=0.63$ and $P_{\mathrm{max}}=4.70$. Notations as in Table 1.
(1) (2) (3) (4)
n=1000n=1~000 n=10000n=10~000 n=10000n=10~000
Model 2 ${\mathrm{SN}}=100$ ${\mathrm{SN}}=100$ ${\mathrm{SN}}=200$
$P_{1}=1.9$ $1.98\pm 0.34$ $1.852\pm 0.055$ $1.933\pm 0.040$
$A_{1}=2.0$ $2.4\pm 2.8$ $1.84\pm 0.20$ $2.13\pm 0.16$
$t_{\mathrm{1,min,1}}=1.35$ $1.39\pm 0.17$ $1.326\pm 0.027$ $1.367\pm 0.020$
$t_{\mathrm{1,max,1}}=0.40$ $0.4015\pm 0.0020$ $0.40007\pm 0.00072$ $0.40017\pm 0.00075$
$M_{0}=1.0$ $0.8\pm 1.4$ $1.079\pm 0.096$ $0.937\pm 0.080$
$M_{1}=0.25$ $0.29\pm 0.21$ $0.229\pm 0.026$ $0.266\pm 0.020$
$M_{2}=0.50$ $0.62\pm 0.52$ $0.451\pm 0.064$ $0.540\pm 0.049$
Data file Model2n1000SN100.dat Model2n10000SN100.dat Model2n10000SN200.dat
Control file dcmModel2n1000SN100.dat dcmModel2n10000SN100.dat dcmModel2n10000SN200.dat
[…]
Refer to caption Refer to caption
                 (a)               (c)
                 (b)               (d)
Refer to caption Refer to caption
                 (e)               (f)
Figure 2: Model 2 (Table 2: n=10000n=10~000, ${\mathrm{SN}}=100$ simulation). (c) Colour of $g(t)$ line has been changed from black to white. (d) Colour of $g(t)$ line has been changed from black to white. Colour of offset level -0.3 dotted line has been changed from blue to white. Locations of best periods (diamonds) are explained in text (Section 3.2). Otherwise, notations are as in Figure 1.
Table 3: Model 3. DCM analysis between $P_{\mathrm{min}}=0.053$ and $P_{\mathrm{max}}=0.480$. Notations as in Table 1.
(1) (2) (3) (4)
$n=50$ $n=50$ $n=100$
Model 3 ${\mathrm{SN}}=10$ ${\mathrm{SN}}=50$ ${\mathrm{SN}}=100$
$P_{1}=0.16$ $0.15898\pm 0.00068$ $0.15989\pm 0.00014$ $0.159950\pm 0.000077$
$A_{1}=2.0$ $1.85\pm 0.10$ $1.987\pm 0.036$ $1.994\pm 0.015$
$t_{\mathrm{1,min,1}}=0.11$ $122\pm 0.0020 0.110$$80\pm 0.00088 0.110$$1004\pm 0.00030$
$t_{\mathrm{1,max,1}}=0.03$ $0328\pm 0.0022 0.0$$3086\pm 0.00094 0.0$$03007\pm 0.00033$
$P_{2}=0.17$ $0.17025\pm 0.00066$ $0.17018\pm 0.00024$ $0.16996\pm 0.00080$
$A_{2}=2.0$ $2.001\pm 0.090$ $1.962\pm 0.046$ $2.002\pm 0.012$
$t_{\mathrm{2,min,1}}=0.135$ $0.1312\pm 0.0022$ $0.13461\pm 0.00087$ $0.13495\pm 0.00031$
$t_{\mathrm{2,max,1}}=0.05$ $0461\pm 0.0024 0.0$$4952\pm 0.00099 0.0$$04997\pm 0.000038$
$M_{0}=1.0$ $0.996\pm 0.011$ $0.9985\pm 0.0032$ $0.9995\pm 0.0012$
Data file Model3n50SN10.dat Model3n50SN50.dat Model3n100SN100.dat
Control file dcmModel3n50SN10.dat dcmModel3n50SN50.dat dcmModel3n100SN100.dat
[…]
Refer to caption Refer to caption
                (a)               (c)
                (b)               (d)
Refer to caption Refer to caption
                (e)               (g)
                 (f)               (h)
Figure 3: Model 3 (Table 3: $n=50$, ${\mathrm{SN}}=50$ simulation). (a) DCM long search periodograms $z_{1}(f_{1})$ (red) and $z_{2}(f)$ (blue) give best periods at 0.160 and 0.168 (diamonds). (b) DCM short search periodograms $z_{1}(f_{1})$ (red) and $z_{2}(f)$ (blue) give best periods at 0.160 and 0.170 (diamonds). (c) DCM model $g(t)$ (black continuous line), DCM trend $p(t)$ (black dashed line) and data $y_{i}$ (black dots). (d) DCM model detrended g(t)p(t)g(t)-p(t) (black continuous line), DCM signal $h_{1}(t)$ (red thick continuous line), DCM signal $h_{2}(t)$ (blue continuous thin line), detrended data $y(t_{i})-p(t_{i})$ (black dots) and DCM model residuals $y(t_{i})-g(t_{1})$ (blue dots) offset to -3.0 level (blue dotted line) (e) DFT periodogram $z_{\mathrm{DFT}}(f)$ for the original data gives best period at 0.163 (diamond). (f) DFT periodogram $z_{\mathrm{DFT}}(f)$ for the sine model residuals gives best period at 0.182 (diamond). (g) DFT model $g_{\mathrm{DFT}}(t)$ (black continuous line), DFT trend $p_{\mathrm{DFT}}(t)$ (black dashed line) and data $y_{i}$ (black dots). (h) DFT model detrended $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (black continuous line), DFT pure sine model for original data $s_{y,\mathrm{DFT}}(t)$ (red thick continuous line), DFT pure sine model for first residuals $s_{\epsilon,\mathrm{DFT}}(t)$ (blue continuous thin line), detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and DFT model residuals (blue dots) offset to -3.0 level (blue dotted line).
Table 4: Model 4. DCM analysis between $P_{\mathrm{min}}=0.053$ and $P_{\mathrm{max}}=0.480$. Notations as in Table 1.
(1) (2) (3) (4)
$n=50$ $n=50$ $n=100$
Model 4 ${\mathrm{SN}}=10$ ${\mathrm{SN}}=50$ ${\mathrm{SN}}=100$
$P_{1}=0.16$ $0.1606\pm 0.0084$ $0.15982\pm 0.00023$ $0.16001\pm 0.000058$
$A_{1}=2.0$ $1.97\pm 0.15$ $1.958\pm 0.051$ $2.012\pm 0.014$
$t_{\mathrm{1,min,1}}=0.11$ $134\pm 0.0036 0.110$$1059\pm 0.00088$ $0.10977\pm 0.00033$
$t_{\mathrm{1,max,1}}=0.03$ $0331\pm 0.0040 0.0$$3068\pm 0.00097 0.0$$02976\pm 0.00035$
$P_{2}=0.17$ $0.1744\pm 0.0012$ $0.17009\pm 0.00026$ $0.169958\pm 0.000092$
$A_{2}=2.0$ $1.74\pm 0.17$ $1.982\pm 0.044$ $2.018\pm 0.017$
$t_{\mathrm{2,min,1}}=0.135$ $0.1321\pm 0.0044$ $0.1342\pm 0.0011$ $0.13528\pm 0.00042$
$t_{\mathrm{2,max,1}}=0.05$ $0460\pm 0.0048 0.0$$492\pm 0.0012 0.0$$05031\pm 0.00034$
$M_{0}=1.0$ $1.015\pm 0.024$ $0.9997\pm 0.0030$ $1.0012\pm 0.0016$
$M_{1}=0.25$ $0.219\pm 0.023$ $0.2565\pm 0.0066$ $0.2495\pm 0.0019$
$M_{2}=0.5$ $0.485\pm 0.050$ $0.511\pm 0.011$ $0.4983\pm 0.0033$
Data file Model4n50SN10.dat Model4n50SN50.dat Model4n100SN100.dat
Control file dcmModel4n50SN10.dat dcmModel4n50SN50.dat dcmModel4n100SN100.dat
[…]
Refer to caption Refer to caption
                 (a)               (c)
                 (b)               (d)
Refer to caption Refer to caption
                 (e)               (g)
                 (f)               (h)
Figure 4: Model 4 (Table 4: $n=50$, ${\mathrm{SN}}=50$ simulation). Notations as in Figure 3. Best periods (diamonds) are explained in Section 3.4.
3.2 Model 2
Our next one signal data simulation model is
$g(t)$ $=$ $(A_{1}/2)\cos{\left[{{2\pi(t-t_{\mathrm{1,max,1}})}\over{P_{1}}}\right]}$
$+$ $M_{0}+M_{1}T+M_{2}T^{2}$
where $T=[2(t-t_{\mathrm{mid}})]/\Delta T$. We give the $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values in Table 2. As a DCM model, the orders of Model 2 are $K_{1}=1$, $K_{2}=1$ and $K_{3}=2$. The simulated data sample is “too short” because the period $P_{1}$ is $1.9\times\Delta T$ (Equation 26). The parabolic trend $p(t)$ is unknown (Equation 27). Again, we use the DCM and DFT time series analysis methods to search for periods between $P_{\mathrm{min}}=P_{1}/3=0.63$ and $P_{\mathrm{max}}=3P_{1}=4.70$.
The DCM analysis results are given in Table 2. These results are not very accurate for the $n=100$ and ${\mathrm{SN}}=100$ combination, but they definitely improve for larger $n$ and $\mathrm{SN}$} values. For Model 2, the DCM analysis results are illustrated for the n=10000n=10~000 and ${\mathrm{SN}}=100$ combination (Figures 2a-d). The DCM long search $z_{1}(f_{1})$ periodogram minimum is at $P_{1}=1.843$ (Figure 2a: diamond). The DCM short search gives the value $P_{1}=1.852$ (Figure 2b: diamond). The DCM model $g(t)$ is so good that it’s continuous black line is totally covered by the black dots representing the $y_{i}$ data (Figure 2c). Therefore, the colour of this $g(t)$ line has been changed from black to white. The results for the parabolic trend coefficients $M_{0}=1.079\pm 0.096$, $M_{1}=0.229\pm 0.026$ and $M_{2}=0.451\pm 0,064$ of the dashed black $p(t)$ line are correct. In Figure 1d, the white continuous line shows the detrended model g(t)p(t)g(t)-p(t). The black dots show the detrended data $y(t_{i})-p(t_{i})$ and the red thick continuous line shows the pure sine signal $h_{1}(t)$. Note that the red thick line is under the white thin line because $h_{1}(t)=g(t)-p(t)$. The DCM residuals (blue dots) are offset to the level of -0.65. The colour of dotted line, which denotes this offset level, has been changed from blue to white. The distribution of these DCM model residuals is stable, as expected for a random normal distribution.
The wrong period $P_{1}=0.697$ is detected by DFT (Figure 2e: Diamond). The DFT estimates for the trend $p(t)$ coefficients, $M_{0}=1.92$, $M_{1}=-0.15$ and $M_{2}=-0.56$ are also wrong (Figure 2f: dashed black line). The data $y_{i}$ (black dots) deviate from the DFT model $g_{\mathrm{DFT}}(t)$ (black continuous line), especially in the end of the sample (Figure 2g). For the detrended DFT model, the thin black line covers the thick red line because $s_{\mathrm{y,DFT}}(t)=g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (Figure 2g). The $s_{\mathrm{y,DFT}}(t)$ sine curve peak to peak amplitude is far below the correct $A_{1}=2.0$ value. The DFT residuals (blue dots) are offset to the level of -1.5 (blue dotted line). The trends of these residuals confirm that the DFT analysis fails.
Only DCM (not DFT) succeeds in the analysis of Model 2 simulated data.
3.3 Model 3
Our third data simulation model is the two signal model
$g(t)$ $=$ $(A_{1}/2)\cos{\left[{{2\pi(t-t_{\mathrm{1,max,1}})}\over{P_{1}}}\right]}$
$+$ $(A_{2}/2)\cos{\left[{{2\pi(t-t_{\mathrm{2,max,1}})}\over{P_{2}}}\right]}+M_{0}.$
We give the $P_{1}$, $P_{2}$, $A_{1}$$A_{2}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{2,max,1}}$ and $M_{0}$ values in Table 3 (Column 1). The DCM orders of Model 3 are $K_{1}=2$, $K_{2}=1$ and $K_{3}=0$. The simulated data sample is not “too short” (Equation 26) because $P_{1}<P_{2}<\Delta T$. The $p(t)$ trend mean level $M_{0}$ is unknown (Equation 27). The two frequencies are “too close” because $\Delta f=f_{1}-f_{2}=f_{0}/2.72$ (Equation 28). We apply the DCM and DFT time series analyses to search for periods between $P_{\mathrm{min}}=P_{1}/3=0.053$ and $P_{\mathrm{max}}=3P_{1}=0.480$.
The DCM analysis results for different $n$ and $\mathrm{SN}$} combinations are given in Table 3 (Columns 2-4). The results are surprisingly accurate even for the small $n=50$ sample having a low ${\mathrm{SN}}=10$. The DCM results for Model 3 are illustrated in Figures 1a-d ($n=50$ and ${\mathrm{SN}}=50$ combination). The DCM long search $z_{1}(f_{1})$ periodogram (red) and $z_{2}(f)$ periodogram (blue) minima are at the best periods $P_{1}=0.159$ and $P_{2}=0.168$ (Figure 3a: diamonds). The DCM short search gives $P_{1}=0.160$ and $P_{2}=0.170$ (Figure 3b: diamonds). The black continuous DCM model $g(t)$ curve crosses trough the black dots of $y_{i}$ data (Figure 1c). The result $M_{0}=0.9985\pm 0.0032$ for the dashed black $p(t)$ trend line is correct. Our Figure 3d shows the detrended DCM model g(t)p(t)g(t)-p(t) (black continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the pure sine signal $h_{1}(t)$ (red thick continuous line) and the pure sine signal $h_{2}(t)$ (blue thin continuous line). The DCM residuals (blue dots) are offset to the level of -3.0 (blue dotted line). These residuals are small and their level is stable.
The DFT detects the wrong period $P_{1}=1.163$ for the original data (Figure 3e: diamond). This is an expected result because the detected period should be close to $(P_{1}+P_{2})/2$ when the peak to peak amplitudes of the simulated data, $A_{1}=A_{2}$, are equal (Jetsu 2025). The two DFT periodogram $z_{\mathrm{DFT}}(f)$ peaks at frequencies $1/P_{1}$ and $1/P_{2}$ overlap and merge into one peak. The period $P_{2}=0.182$ detected for the residuals is also wrong (Figure 3f: diamond). The DFT trend $p_{\mathrm{PDF}}(t)$ estimate $M_{0}=0.880$ fails (Figure 3f: dashed black line). The black dots $y_{i}$ show minor deviations from the continuous black DFT model $g_{\mathrm{DFT}}(t)$ line (Figure 3g). The detrended model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (continuous black line), the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots), the pure sine signal $s_{\mathrm{y,DFT}}(t)$ for the data (continuous thick red line) and the pure sine signal $s_{\mathrm{\epsilon,DFT}}(t)$ for the residuals (continuous blue thin line) are shown in Figure 3h. Note that the $s_{\mathrm{y,DFT}}(t)$ and $s_{\mathrm{\epsilon,DFT}}(t)$ signal amplitudes are far from equal. The DFT residuals (blue dots) offset to the level of -3.0 (blue dotted line) are not stable.
The DCM analysis of Model 3 simulated data succeeds, but the DFT analysis does not.
3.4 Model 4
The next data simulation model is
$g(t)$ $=$ $(A_{1}/2)\cos{\left[{{2\pi(t-t_{\mathrm{1,max,1}})}\over{P_{1}}}\right]}$
$+$ $(A_{2}/2)\cos{\left[{{2\pi(t-t_{\mathrm{2,max,1}})}\over{P_{2}}}\right]}$
$+$ $M_{0}+M_{1}T+M_{2}T^{2},$
where $T=[2(t-t_{\mathrm{mid}})]/\Delta T$. In this model, two signals are superimposed on an unknown parabolic trend. The $P_{1}$, $P_{2}$, $A_{1}$$A_{2}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{2,max,1}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values are given in Table 4. This simulated sample is not “too short” (Equation 26). The polynomial trend $p(t)$ is unknown (Equation 27). The $\Delta f=f_{1}-f_{2}=f_{0}/2.72$ difference means that the frequencies are “too close” (Equation 28). The DCM and DFT time series analysis methods are used to search for periods between $P_{\mathrm{min}}=P_{1}/3=0.053$ and $P_{\mathrm{max}}=3P_{1}=0.480$.
Model 4 is a DCM model having orders $K_{1}=2$, $K_{2}=1$ and $K_{3}=2$. Our DCM analysis results for different $n$ and $\mathrm{SN}$} combinations are given in Table 4 (Columns 2-4). The DCM detects the correct $P_{1}$, $P_{2}$, $A_{1}$, $A_{2}$, $t_{\mathrm{max,1}}$, $t_{\mathrm{max,2}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values even for the lowest $n=50$ and ${\mathrm{SN}}=10$ combination. Our Figures 4a-d illustrate the DCM analysis results for one sample of simulated data (Model 4: $n=50$ and ${\mathrm{SN}}=50$). The DCM long search best periods are at $P_{1}=0.159$ and $P_{2}=0.168$ (Figure 4a: diamonds). The DCM short search values are $P_{1}=0.160$ and $P_{2}=0.170$ (Figure 4b: diamonds). The continuous DCM model $g(t)$ black line crosses through all black dots representing the $y_{i}$ data (Figure 4c). The DCM detects the correct polynomial trend $p(t)$ coefficients $M_{0}=0.9997\pm 0.0030$, $M_{1}=0.2565\pm 0.0066$ and $M_{2}=0.511\pm 0.011$ (Figure 4c: dashed black line). The detrended DCM model g(t)p(t)g(t)-p(t) (black continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the pure sine signal $h_{1}(t)$ (red thick continuous line) and the pure sine signal $h_{2}(t)$ (blue thin continuous line) are shown in Figure 4d. The DCM residuals (blue dots) offset to the level of -3.0 show no trends and are extremely stable.
Since the peak to peak amplitudes of the simulated data signals are equal, $A_{1}=A_{2}$, the expected result for the DFT analysis of original data is $(P_{1}+P_{2})/2$, where $P_{1}$ and $P_{2}$ are the simulated signal periods (Jetsu 2025). The DFT detects this expected wrong period $P_{1}=1.165$ for the original data (Figure 4e: diamond). A wrong period $P_{2}=0.147$ is also detected for the residuals (Figure 4f: diamond). The DFT estimate $M_{0}=0.962$ for the $p(t)$ trend is close to the correct value $M_{0}=1$, but the $M_{1}=-0.017$ and $M_{2}=0.903$ estimates are wrong (Figure 4g: dashed black line). The continuous black DFT model $g_{\mathrm{DFT}}(t)$ line deviates from the black dots of data $y_{i}$, especially in the beginning and end of the sample. In our Figure 4h, the black dots are the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ and the continuous black line is the detrended DFT model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$. The continuous thick red line is the pure sine signal $s_{\mathrm{y,DFT}}(t)$ for the original data and the continuous thinner blue line is the pure sine signal $s_{\mathrm{\epsilon,DFT}}(t)$ for the residuals. The $s_{\mathrm{\epsilon,DFT}}(t)$ signal amplitude is far below the simulated correct value $A_{2}=2.0$. The blue dots representing the DFT model residuals are offset to the level of -3.0 and show clear trends.
Our Model 4 simulated data analysis succeeds for the DCM and fails for the DFT.
Table 5: Model 5. DCM analysis between $P_{\mathrm{min}}=0.47$ and $P_{\mathrm{max}}=4.20$. Notations as in Table 1.
(1) (2) (3) (4)
n=10000n=10~000 n=10000n=10~000 n=100000n=100~000
Model 5 ${\mathrm{SN}}=1~000$ ${\mathrm{SN}}=10~000$ ${\mathrm{SN}}=10~000$
$P_{1}=1.4$ $1.404\pm 0.046$ $1.393\pm 0.014$ $1.4019\pm 0.0035$
$A_{1}=2.0$ $2.27\pm 0.45$ $1.97\pm 0.13$ $1.994\pm 0.029$
$t_{\mathrm{1,min,1}}=1.1$ $129\pm 0.023 1.1$$1030\pm 0.0028$ $1.09766\pm 0.00082$
$t_{\mathrm{1,max,1}}=0.4$ $0.4267\pm 0.0012$ $0.4066\pm 0.0059$ $0.3967\pm 0.0020$
$P_{2}=1.9$ $5.24\pm 0.83$ $2.023\pm 0.094$ $1.854\pm 0.023$
$A_{2}=2.0$ $5\pm 24$ $2.35\pm 0.34$ $1.918\pm 0.074$
$t_{\mathrm{2,min,1}}=1.55$ $3.23\pm 0.50$ $1.599\pm 0.030$ $1.5292\pm 0.0066$
$t_{\mathrm{2,max,1}}=0.6$ $0.61\pm 0.34$ $0.587\pm 0.030$ $0.6024\pm 0.0064$
$M_{0}=1.0$ $2\pm 12$ $0.81\pm 0.16$ $1.055\pm 0.035$
$M_{1}=0.25$ $1.27\pm 0.53$ $0.230\pm 0.013$ $0.2590\pm 0.0029$
$M_{2}=0.5$ $3.8\pm 1.8$ $0.596\pm 0.084$ $0.470\pm 0.020$
Data file Model5n10000SN1000.dat Model5n10000SN10000.dat Model5n100000SN10000.dat
Control file dcmModel5n10000SN1000.dat dcmModel5n10000SN10000.dat dcmModel5n100000SN10000.dat
[…]
Refer to caption Refer to caption
                 (a)               (c)
                 (b)               (d)
Refer to caption Refer to caption
                 (e)               (g)
                 (f)               (h)
Figure 5: Model 5 (Table 5: n=10000n=10~000, ${\mathrm{SN}}=10~000$ simulation). Notations as in Figure 3. Best periods (diamonds) are explained in Section 3.5.
Table 6: Model 6. DCM analysis between $P_{\mathrm{min}}=0.053$ and $P_{\mathrm{max}}=0.480$. Notations as in Table 1.
(1) (2) (3) (4)
$n=50$ $n=50$ $n=100$
Model 6 ${\mathrm{SN}}=10$ ${\mathrm{SN}}=50$ ${\mathrm{SN}}=100$
$P_{1}=0.16$ $0.15966\pm 0.00020$ $0.160022\pm 0.000021$ $0.160024\pm 0.000019$
$A_{1}=2.0$ $1.983\pm\pm 0.041$ $2.0031\pm 0.0086$ $2.0013\pm 0.0032$
$t_{\mathrm{1,min,1}}=0.0979$ $09868\pm 0.00057 0.0$$97933\pm 0.000096 0.0$$097934\pm 0.000079$
[…]
$t_{\mathrm{1,max,1}}=0.1421$ $0.14241\pm 0.00064$ $0.142099\pm 0.000079$ $0.142100\pm 0.000077$
$t_{\mathrm{1,max,2}}=0.0575$ $05827\pm 0.00057 0.0$$5761\pm 0.00011 0.0$$057448\pm 0.000074$
$M_{0}=0$ $003\pm 0.012 0.0$$055\pm 0.0026 0.0$$00176\pm 0.00088$
Data file Model6n50SN10.dat Model6n50SN50.dat Model6n100SN100.dat
Control file dcmModel6n50SN10.dat dcmModel6n50SN50.dat dcmModel6n100SN100.dat
[…]
Refer to caption Refer to caption
                 (a)               (c)
                 (b)               (d)
Refer to caption Refer to caption
                 (e)               (g)
                 (f)               (h)
Figure 6: Model 6 (Table 6: $n=50$, ${\mathrm{SN}}=50$ simulation). Notations as in Figure 3. Best periods (diamonds) are explained in Section 3.6.
Table 7: Model 7. DCM analysis between $P_{\mathrm{min}}=0.4$ and $P_{\mathrm{max}}=3.6$. Notations as in Table 1.
(1) (2) (3) (4)
$n=100$ n=1000n=1~000 n=10000n=10~000
Model 7 ${\mathrm{SN}}=5~000~000$ ${\mathrm{SN}}=1~000~000$ ${\mathrm{SN}}=1~000~000$
$P_{1}=1.2$ $1.19917\pm 0.00070$ $1.20000\pm 0.00063$ $1.20000\pm 0.00024$
$A_{1}=2.0$ $2.20\pm 0.20$ $1.94\pm 0.14$ $1.990\pm 0.057$
$t_{\mathrm{1,min,1}}=0.6892$ $0.6835\pm 0.0050$ $0.6900\pm 0.0043$ $0.6888\pm 0.0019$
$t_{\mathrm{1,min,2}}=0.1134$ $0.1235\pm 0.0098$ $0.1160\pm 0.0067$ $0.1128\pm 0.0026$
$t_{\mathrm{1,max,1}}=1.0195$ $1.0253\pm 0.0052$ $1.0176\pm 0.0039$ $1.0188\pm 0.0018$
$t_{\mathrm{1,max,2}}=0.3779$ $0.3705\pm 0.0082$ $0.3804\pm 0.0050$ $0.3780\pm 0.0020$
$P_{2}=1.4$ $1.4050\pm 0.0048$ $1.3986\pm 0.00034$ $1.3997\pm 0.0014$
$A_{2}=2.0$ $1.871\pm 0.098$ $2.049\pm 0.095$ $2.009\pm 0.040$
$t_{\mathrm{2,min,1}}=0.9262$ $0.938\pm 0.012$ $0.9231\pm 0.0090$ $0.9252\pm 0.0036$
$t_{\mathrm{2,min,2}}=0.2766$ $0.260\pm 0.016$ $0.281\pm 0.012$ $0.2771\pm 0.0046$
$t_{\mathrm{2,max,1}}=1.3109$ $1.3165\pm 0.0049$ $1.3105\pm 0.0039$ $1.3101\pm 0.0014$
$t_{\mathrm{2,max,2}}=0.5864$ $0.5817\pm 0.0037$ $0.5860\pm 0.0042$ $0.5865\pm 0.0015$
$M_{0}=1.0$ $0.950\pm 0.050$ $1.011\pm 0.035$ $1.002\pm 0.013$
$M_{1}=0.25$ $0.2447\pm 0.0067$ $0.2548\pm 0.0039$ $0.2509\pm 0.0018$
$M_{2}=0.5$ $0.551\pm 0.050$ $0.488\pm 0.035$ $0.498\pm 0.014$
Data file Model7n100SN5000000.dat Model7n1000SN1000000.dat Model7n10000SN1000000.dat
Control file dcmModel7n100SN5000000.dat dcmModel7n1000SN1000000.dat dcmModel7n10000SN1000000.dat
[…]
Refer to caption Refer to caption
                 (a)               (c)
                 (b)               (d)
Refer to caption Refer to caption
                 (e)               (g)
                 (f)               (h)
Figure 7: Model 7 (Table 7: $n=100$, ${\mathrm{SN}}=5~000~000$ simulation). otations as in Figure 3. Best periods (diamonds) are explained in Section 3.7.
3.5 Model 5
The mathematical equation
$g(t)$ $=$ $(A_{1}/2)\cos{\left[{{2\pi(t-t_{\mathrm{1,max,1}})}\over{P_{1}}}\right]}$
$+$ $(A_{2}/2)\cos{\left[{{2\pi(t-t_{\mathrm{2,max,1}})}\over{P_{2}}}\right]}$
$+$ $M_{0}+M_{1}T+M_{2}T^{2}$
for our fifth model for simulated data is the same for Model 4 (Equation 3.4). However, this new Model 5 differs from the earlier Model 4 because we use totally different $P_{1}$, $P_{2}$, $A_{1}$$A_{2}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{2,max,1}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values (Table 5, Column 1). The simulated data sample is “too short” for both $P_{1}$ and $P_{2}$ periods (Equation 26). There is “a trend” $p(t)$ (Equation 27). The signal frequencies $f_{1}$ and $f_{2}$ are “too close” (Equation 28). The three main reasons that can cause the failure of DFT analysis are present. We perform the DCM and DFT time series analysis between $P_{\mathrm{min}}=P_{1}/3=0.47$ and $P_{\mathrm{max}}=3P_{1}=4.20$.
Model 5 has DCM orders $K_{1}=2$, $K_{2}=1$ and $K_{3}=2$. We give the DCM analysis results for different $n$ and $\mathrm{SN}$} combinations in Table 5 (Columns 2-4). The DCM fails to detect the correct $P_{1}$, $P_{2}$, …, $M_{1}$ and $M_{2}$ values for the the lowest n=10000n=10~000 and ${\mathrm{SN}}=1~000$ combination (Table 5, Column 2). This shows that the DCM can fail, just like any other time series analysis method, if the quality of data is too low. The DCM results for higher $n$ and $\mathrm{SN}$} combinations are correct (Table 5, Columns 3-4).
We show the DCM analysis results for one sample of Model 5 simulated data (Figures 5a-d: n=10000n=10~000 and ${\mathrm{SN}}=10~000$). The long and short DCM searches give $P_{1}=1.407$ and $P_{2}=1.917$ (Figure 5a: diamonds), and $P_{1}=1.393$ and $P_{2}=2.024$ (Figure 5b: Diamonds). The continuous line denoting the model $g(t)$ is totally covered by the black dots of $y_{i}$ data (Figure 5c). Therefore, we use white colour to highlight this DCM model $g(t)$ line. The detected polynomial trend $p(t)$ coefficients $M_{0}=0.81\pm 0.16$, $M_{1}=0.230\pm 0.013$ and $M_{2}=0.596\pm 0.084$ are correct (Figure 5c: dashed black line). We show the detrended DCM model g(t)p(t)g(t)-p(t) (white continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the signal $h_{1}(t)$ (red thick continuous line) and the signal $h_{2}(t)$ (blue thin continuous line) in Figure 5d. The DCM residuals (blue dots) are offset to the level of -1.5. These blue dots appear black because there are 10 000 of them. For obvious reasons, the -1.5 level of these residuals is highlighted by a white dotted line. The DCM model residuals are stable and show no trends.
The DFT detects the wrong periods for the original data (Figure 5e: diamond at $P_{1}=0.592$) and for the residuals (Figure 5f: diamond at $P_{2}=1.012$). The DFT estimates for the $p(t)$ trend, $M_{0}=2.758$, $M_{1}=0.091$ and $M_{2}=-2.120$, are also wrong (Figure 5g: dashed black line). The DFT model $g_{\mathrm{DFT}}(t)$ black continuous line deviates from the black dots of data $y_{i}$, especially in the end of the simulated data sample. The black dots denoting the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ and the continuous black line denoting the detrended DFT model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ are shown in Figure 5h. The DFT model $g_{\mathrm{DFT}}(t)$ gives very low amplitudes for the pure sine signal $s_{\mathrm{y,DFT}}(t)$ (continuous thick red line) and pure sine signal $s_{\mathrm{\epsilon,DFT}}(t)$ (continuous thinner blue line). These amplitudes are far below the correct simulated values $A_{1}=A_{2}=2.$ The offset level for the DFT model residuals (blue dots) is -0.5. These residuals show strong trends.
The DCM analysis succeeds for Model 5 simulated data, but the DFT analysis fails.
3.6 Model 6
Our sixth data simulation model is
$g(t)$ $=$ $c_{1}\cos{\left[{{2\pi t}\over{P_{1}}}\right]}$
$+$ $c_{2}\cos{\left[{{4\pi(t-c_{3})}\over{P_{1}}}\right]}+M_{0},$
where $P=0.16$ and $M_{0}=0$. The coefficients $c_{1}=0.3655$, $c_{2}=0.7310$ and $c_{3}=0.3000$ determine the $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,min,2}}$, $t_{\mathrm{1,max,1}}$ and $t_{\mathrm{1,max,2}}$ values given in Table 6 (Column 1). This simulated data sample is not “too short” because $P_{1}<\Delta T$ (Equation 26). There is no trend because $p(t)=M_{0}=0$ (Equation 27). These simulated data contain only one signal (Equation 28). However, this simulation Model 6 is not a $K_{2}=1\texteuro$ pure sine model (Equation 29). This double wave simulation model has two unequal minima and maxima. Its DCM orders are $K_{1}=1$, $K_{2}=2$ and $K_{3}=0$. We perform the DCM and DFT time series analysis between $P_{\mathrm{min}}=P_{1}/3=0.053$ and $P_{\mathrm{max}}=3P_{1}=0.480$.
The DCM time series analysis results for different $n$ and $\mathrm{SN}$} combinations are given in Table 6 (Columns 2-4). The DCM detects the correct $P_{1}$, $A_{1}$, $t_{\mathrm{1,min,1}}$ $t_{\mathrm{1,min,2}}$ $t_{\mathrm{1,max,1}}$ $t_{\mathrm{1,max,2}}$ and $M_{0}$ values even for the lowest $n=50$ and ${\mathrm{SN}}=10$ combination. We demonstrate the DCM analysis results for simulated data having $n=50$ and ${\mathrm{SN}}=50$ (Figures 6a-d). The long and short searches give $P_{1}=0.159$ (Figure 6a: diamond) and $P_{1}=0.160$ (Figure 6b: diamond). The DCM model $g(t)$ black line covers the black dots denoting the data $y_{i}$ (Figure 6c). The "trend" at $p(t)=M_{0}=-0.0055\pm 0.0026$ is correct. The detrended model g(t)p(t)g(t)-p(t) (black continuous line), the pure sine signal $h_{1}(t)$ (red thick continuous line) and the detrended data $y(t_{i})-p(t_{i})$ (black dots) are displayed in Figure 6d. The thick continuous red line stays under the thin continuous black line because $h_{1}(t)=g(t)-p(t)$. The DCM residuals (blue dots) are offset to the level of -1.8 (blue dotted line). These residuals show no trends and their level is stable.
The DFT detects the wrong period $P_{1}=0.080$ (Figure 6e: Diamond). This results is exactly half of the correct simulated value $P_{1}=0.160$. The reason for this “detection” is that the double wave dominates because $c_{2}=2c_{1}$ in Model 6 (Equation 3.7). The DFT mean level estimate $M_{0}=-0.172$ fails (Figure 6f: Dashed black line). The DFT analysis of the residuals $\epsilon_{i}=$ $y(t_{i})-[s_{\mathrm{y,DFT}}(t_{i})+p_{\mathrm{DFT}}(t_{i})]$ gives $P_{2}=0.161$, which is nearly equal to the correct simulated $P_{1}=0.160$ value. The black dots denoting the data $y_{i}$ show minor deviations from continuous black line denoting the DFT model $g_{\mathrm{DFT}}(t)$ (Figure 6g). The detrended model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (continuous black line), the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and the signal $s_{\mathrm{y,DFT}}(t)$ (continuous thick red line) are shown in Figure 6h. The DFT analysis residuals (blue dots) are offset to the level of -1.5 (blue dotted line). These residuals display trends.
We conclude that the DFT “detects” the $P_{1}/2$ and $P_{1}$ periods, where $P_{1}$ is the correct simulated period value. However, the DFT two signal model is not the correct model for these Model 6 simulated data, which contain only one signal. If the correct period is $P$ and the correct model is a double wave $(K_{2}=2)$, the DCM pure sine model $(K_{1}=1)$ analysis can also give the values $P/2$ and $P$ [14].
The DCM analysis succeeds for Model 6 simulated data. The DFT analysis fails.
3.7 Model 7
Our seventh simulation model is
$g(t)$ $=$ $c_{1}\cos{\left[{{2\pi t}\over{P_{1}}}\right]}+c_{2}\cos{\left[{{4\pi(t-c_{3})}\over{P_{1}}}\right]}$
$+$ $c_{4}\cos{\left[{{2\pi t}\over{P_{2}}}\right]}+c_{5}\cos{\left[{{4\pi(t-c_{6})}\over{P_{2}}}\right]}$
$+$ $M_{0}+M_{1}T+M_{2}T^{2}$
where $T=[2(t-t_{\mathrm{mid}})]/\Delta T$, $P_{1}=1.2$, $P_{2}=1.4$, $M_{0}=1$ $M_{1}=0.25$ and $M_{2}=0.5$. The coefficients $c_{1}=0.3687$, $c_{2}=0.7374$, $c_{3}=0.4000$, $c_{4}=0.3708$, $c_{5}=0.7416$ and $c_{6}=0.6000$ determine the $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,min,2}}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{1,max,2}}$, $A_{2}$, $t_{\mathrm{2,min,1}}$, $t_{\mathrm{2,min,2}}$, $t_{\mathrm{2,max,1}}$ and $t_{\mathrm{2,max,2}}$ values given in Table 7 (Column 1). This simulated data sample is “too short” because $\Delta T<P_{1}<P_{2}$ (Equation 26). The parabolic $p(t)$ represents “a trend” (Equation 27). The signal frequencies $f_{1}$ and $f_{2}$ are “too close” because $\Delta f=f_{1}-f_{2}=0.12<f_{0}=1$ (Equation 28). The two $h_{1}(t)$ and $h_{2}(t)$ signals are not “pure sines” (Equation 29). These signals are double waves having two unequal minima and maxima. All four main reasons that can cause the failure of DFT analysis are present (Equations 26-29). Our DCM and DFT time series analysis of Model 7 simulated data is performed between $P_{\mathrm{min}}=P_{1}/3=0.4$ and $P_{\mathrm{max}}=3P_{1}=3.6$.
The DCM orders of Model 7 are $K_{1}=2$, $K_{2}=2$ and $K_{3}=2$. This model has $\eta=13$ free parameters (Equation 8). We give the DCM analysis results for different $n$ and $\mathrm{SN}$} combinations in Table 7 (Columns 2-4). The DCM analysis results are displayed for one sample of Model 7 simulated data (Figures 7a-d: $n=100$ and ${\mathrm{SN}}=5~000~000$). Since there are $\eta=13$ free model parameters, this $n=100$ sample is quite small for time series analysis. The long and short DCM searches give $P_{1}=1.200$ and $P_{2}=1.385$ (Figure 7a: diamonds), and $P_{1}=1.199$ and $P_{2}=1.405$ (Figure 7b: Diamonds). The black $g(t)$ model line goes through all black $y_{i}$ data dots (Figure 7c). The DCM detects the correct polynomial trend $p(t)$ coefficients $M_{0}=0.950\pm 0.050$, $M_{1}=0.2447\pm 0.0067$ and $M_{2}=0.551\pm 0.050$ (Figure 7c: dashed black line). The detrended DCM model g(t)p(t)g(t)-p(t) (white continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the signal $h_{1}(t)$ (red thick continuous line) and the signal $h_{2}(t)$ (blue thin continuous line) are displayed in Figure 7d. The $n=100$ residuals (blue dots) are offset to the level of -1.8 (dotted blue line). These DCM model residuals are small and their level is stable. These results confirm that the DCM time series analysis method can detect complex non-linear models $(\eta=13)$ from very small data samples $(n=100)$, if these data are extremely accurate $({\mathrm{SN}}=5~000~000).$
The DFT time series analysis gives the wrong periods for the original data (Figure 7e: diamond at $P_{1}=0.542$ and for the residuals (Figure 7f: diamond at $P_{2}=1.059$. The DFT also gives wrong $p(t)$ trend coefficients $M_{0}=0.670$, $M_{1}=-0.025$ and $M_{2}=0.406$ (Figure 7g: dashed black line). The DFT model $g_{\mathrm{DFT}}(t)$ (black continuous line) deviates from the data $y_{i}$ (black dots). This deviation is largest at the beginning and the end of the simulated data sample. In our Figure 7h, the black dots denote the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$, the continuous black line denotes the detrended DFT model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ and the continuous thick red line denotes pure sine signal $s_{\mathrm{y,DFT}}(t)$ detected from the original data. The DFT model $g_{\mathrm{DFT}}(t)$ gives very low amplitude for the pure sine signal $s_{\mathrm{\epsilon,DFT}}(t)$ detected from the residuals (continuous thinner blue line). The correct simulated peak to peak amplitude for this second signal is much larger, $A_{2}=2.$ The DFT model residuals (blue dots) are offset to -1.5 level and show strong trends.
The DCM time series analysis succeeds for Model 7 simulated data. The DFT time series analysis fails, as predicted by Equations 26-29.
Table 8: Fisher tests. DCM analysis of Model 1 simulated data ($n=100,{\mathrm{SN}}=100$, data file Model1n100SN100). (1) Simple model: Model number $\mathcal{M}$, model $g_{K_{1},K_{2},K_{3}}$, number of free parameters $\eta$, chi-square $\chi^{2}$ and control file name. (2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.
[…]
22footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
3.8 Best model
The $K_{1}$, $K_{2}$ and $K_{3}$ orders of the best model are not necessarily known when some time series analysis method is applied to the real data. We know a priori the best DCM and DFT model orders for the simulated data of Models 1-7 (Sections 3.1-3.7). It could be argued that our DCM analysis succeeds only for this reason.
All alternative $K_{1}$, $K_{2}$ and $K_{3}$ order models are nested. For example, the simple one signal $K_{1}=1$ model $g_{1}(t)$ is a special case of the complex two signal $K_{1}=2$ model $g_{2}(t)$ having $A_{2}=0$. We use the Fisher-test to compare any pair of simple $g_{1}(t)$ and complex $g_{2}(t)$ models. The model parameters (Equation 13: $R_{1},R_{2}$ ), (Equation 14: $\chi_{1}$$\chi_{2}$) and (Equation 8: $\eta_{1}<\eta_{2}$) give the test statistic of Fisher-test
$F_{R}$ $=$ $\left({{R_{1}}\over{R_{2}}}-1\right)\left({{n-\eta_{2}-1}\over{\eta_{2}-\eta_{1}}}\right)$ (37)
$F_{\chi}$ $=$ $\left({{\chi_{1}^{2}}\over{\chi_{2}^{2}}}-1\right)\left({{n-\eta_{2}-1}\over{\eta_{2}-\eta_{1}}}\right).$ (38)
The Fisher-test is based on the null hypothesis
  • $H_{\mathrm{0}}$: “The complex model $g_{2}(t)$ does not provide a significantly better fit to the data than the simple model $g_{\mathrm{1}}(t)$.”
Under this hypothesis, the $F_{R}$ and $F_{\chi}$ test statistic parameters have an $F$ distribution with $\nu_{1}=\eta_{2}-\eta_{1}$ and $\nu_{2}=n-\eta_{2}$ degrees of freedom [4]. The critical level $Q_{F}=P(F_{R}\geq F)$ or $Q_{F}=P(F_{\chi}\geq F)$ is the probability that $F_{R}$ or $F_{\chi}$ exceeds the numerical value $F$. If
$Q_{F}<\gamma_{F}=0.001,$ (39)
we reject the $H_{0}$ hypothesis, which means that the complex $g_{2}(t)$ model is better than the simple $g_{1}(t)$ model. The pre-assigned significance level $\gamma_{F}=0.001$ represents the probability that we falsely reject the $H_{0}$ hypothesis when it is in fact true.
Larger $F_{R}$ or $F_{\chi}$ values have smaller $Q_{F}$ critical levels. Hence, the probability for the $H_{0}$ hypothesis rejection increases when $F_{R}$ or $F_{\chi}$ increases. If the number of complex model free parameters $\eta_{2}$ increases, the $R_{2}$ or $\chi^{2}_{2}$ values decrease. This increases the $F_{R}$ or $F_{\chi}$ values because the terms $(R_{1}/R_{2}-1)$ or $(\chi^{2}_{1}/\chi^{2}_{2}-1)$ (Equations 37 and 38: first terms). At the same time, the $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty term decreases (Equations 37 and 38: second terms), and this decreases the $F_{R}$ or $F_{\chi}$ values. This penalty term prevents the use of too high $\eta_{2}$ values (too complex models).
[…]
Table 9: Fisher tests for Model 1 $n=100,{\mathrm{SN}}=100$ simulated data (electronic data file Model1n100SN100). (1) Polynomial model. (2) DCM model $g_{1,1,0}$. Note that $\mathcal{M}$=1-3  polynomials represent simple models, and $\mathcal{M}$=5-8  polynomials represent complex models. Otherwise as in Table 8.
(1) (2)
$g_{1,1,0}$
Polynomial $\eta=4,\chi^{2}=84.7422$
$\mathcal{M}$=1 $\uparrow$
$g(t)=p(t,K_{3}=0)$ $F=63218$
$\eta=1$, $\chi^{2}=169260$ $Q_{\mathrm{F}}<\!10^{-16}$
$\mathcal{M}$=2 $\uparrow$
$g(t)=p(t,K_{3}=1)$ $F=54366$
$\eta=2$, $\chi^{2}=97076$ $Q_{\mathrm{F}}<\!10^{-16}$
$\mathcal{M}$=3 $\uparrow$
$g(t)=p(t,K_{3}=2)$ $F=1945$
$\eta=3$, $\chi^{2}=1820$ $Q_{\mathrm{F}}<\!10^{-16}$
$\mathcal{M}$=4 $\uparrow$
$g(t)=p(t,K_{3}=3)$ No test11footnotemark: 1
$\eta=4$, $\chi^{2}=469$
$\mathcal{M}$=5 $\uparrow$
$g(t)=p(t,K_{3}=4)$ F=0.3718F=-0.3718
$\eta=5$, $\chi^{2}=85.0787$ $Q_{\mathrm{F}}=1$
$\mathcal{M}$=6 $\uparrow$
$g(t)=p(t,K_{3}=5)$ $F=0.3403$
$\eta=6$, $\chi^{2}=84.1265$ $Q_{\mathrm{F}}=0.712$
$\mathcal{M}$=7 $\uparrow$
$g(t)=p(t,K_{3}=6)$ $F=0.3524$
$\eta=7$, $\chi^{2}=83.7794$ $Q_{\mathrm{F}}=0.787$
$\mathcal{M}$=8 $\uparrow$
$g(t)=p(t,K_{3}=7)$ $F=0.9394$
$\eta=8$, $\chi^{2}=81.3819$ $Q_{\mathrm{F}}=0.445$
[…]
33footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.
In the above example, the Fisher-test finds the correct number of signals $(K_{1})$ and the correct trend $(K_{3})$ for the pure sine signal alternative $(K_{2}=1)$. We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large. One example of testing the $K_{2}=1$ and $K_{2}=2$ signal models against each other can be found in Jetsu [14, Table 4].
We conclude that the best model for the real data can be found by applying the Fisher-test to any arbitrary number of different nested DCM or DFT models.
3.9 Significance estimates
The Fisher-test gives the critical level $Q_{\mathrm{F}}$ for rejecting the $H_{0}$ hypothesis (Equation 39). After the detection of the first strongest signal, the values $Q_{\mathrm{F}}$ of these critical levels increase for the next detected weaker signals. Typical examples can be found in Jetsu 2025 (Tables S5-S16). No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected.
The Fisher test critical level $Q_{\mathrm{F}}$ is the significance estimate for each signal detected after the first signal. However, this Fisher-test process gives no significance estimate for the first detected signal.
We estimate the critical level $Q_{\mathrm{F}}$ for the first detected signal $h_{1}(t)$ by testing the alternative that there are no signals in the data. The “no signal hypothesis” is
$h(t)=0\Rightarrow h_{1}(t)=0\Rightarrow g(t)=p(t).$ (40)
In Model 1 simulated data combination $n=100$ and ${\mathrm{SN}}=100$, the correct model is $g_{1,1,0}(t)$. We use the Fisher-test to compare this correct $g_{1,1,0}(t)$ model to different polynomial models $g(t)=p(t)$ having $K_{3}=0,1,..,7$ (Table 9).
The $\mathcal{M}$=1, 2 and 3 models have less free parameters than the correct $g_{1,1,0}(t)$ model, but the $\chi^{2}$ values of these three polynomial models are so large that the Fisher-test is merely a formality. The $\mathcal{M}$=4 model has the same number of free parameters as the correct $g_{1,1,0}(t)$ model, but the comparison of $\chi^{2}$ values reveals that this third order $p(t)$ polynomial is not the correct model for the data. The next model $\mathcal{M}$=5 has more free parameters than the correct $g_{1,1,0}(t)$ model. This fourth order $p(t)$ polynomial $\mathcal{M}$=5 model must be rejected because it has a larger $\chi^{2}$ value than the correct $g_{1,1,0}(t)$ model. The critical levels $Q_{\mathrm{F}}\gg\gamma=0.001$ for the remaining $\mathcal{M}$=6, 7 and 8 polynomial models are so large that these models must also be rejected. The results in Table 9 confirm the “no signal hypothesis” (Equation 40) rejection. The analysed Model 1 simulated data must contain at least the $h_{1}(t)$ pure sine signal. For the constant, linear or parabolic $p(t)$ trend alternatives, the critical level for this $h_{1}(t)$ signal is $Q_{\mathrm{F}}<\!10^{-16}$.
There is no need to discuss the DFT significance estimates [11, their Equation 22] because this method fails to detect the correct Model 1 period.
Table 10: DCM time series analysis between $P_{\mathrm{min}}=0.53$ and $P_{\mathrm{max}}=4.80$ for predictive data (Model = 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data: first forty observations). (1) Model. (2) Predictive data test statistic $z$ (Equation 16). (3) Predicted data test statistic $z_{\mathrm{pred}}$ (Equation 3.10). (4) Data file. (5) Control file. (6) Figure where prediction is shown.
(1) (2) (3) (4) (5) (6)
Model $z$ $z_{\mathrm{pred}}$ Data file Control file Figure
$g_{1,1,0}$ 2.27 8.61 Model3n40SN10.dat dcmModel3K110.dat 8a
$g_{2,1,0}$ 0.76 0.88 Model3n40SN10.dat dcmModel3K210.dat 8b
$g_{3,1,0}$ 0.70 0.91 Model3n40SN10.dat dcmModel3K310.dat 8c
[…]
Figure 8: DCM model prediction (Model 3 simulated data for combination $n=50$ and ${\mathrm{SN}}=10$: first forty observations are predictive data). (a) One signal model $g_{1,1,0}$ results. Predictive data are $t_{i}$, $y_{i}$ and $\sigma_{i}$ ($n=40$ black dots) and predicted data are ten last observations $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$ ($n=10$ open black dots). Continuous black line of predictive data model $g(t,\bm{\beta}_{\mathrm{pred}})$ ends to vertical dotted blue line, where dotted black line of predicted model $g^{\prime}(t,\bm{\beta}_{\mathrm{pred}})$ begins. Dotted red line denotes $\pm 3\sigma$ errors of both models. Residuals of predictive data model $\epsilon_{i}=y_{i}-g_{i}$ (blue dots), residuals of predicted data model $\epsilon_{i}=^{\prime}y_{i}^{\prime}-g_{i}^{\prime}$ (open blue dots) and $\pm 3\sigma$ errors of both models (red dotted line) are offset to level -1 (blue dotted line). (b) Two signal model $g_{2,1,0}$ results. Otherwise as in “a”. (c) Three signal model $g_{3,1,0}$ results. Otherwise as in “a”.
3.10 Prediction
There are numerous techniques for predicting (forecasting) a time series [e.g., 10, 12]. The DCM model $g(t_{i},\bm{\beta})$ can be used to predict. We divide all data into the predictive data and the predicted data. The time points, the observations and the errors of these samples are
  • $n$ predictive data values $t_{i}$, $y_{i}$ and $\sigma_{i}$
  • $n^{\prime}$ predicted data values $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$
The DCM gives the best predictive data model
$g_{i}=g(t_{i},\bm{\beta}_{\mathrm{pred}}),$ (41)
where $\bm{\beta}_{\mathrm{pred}}$ are the free parameter values. The $t_{\mathrm{mid,pred}}\!=\!(t_{n}\!+\!t_{1})/2$ and $\Delta T_{\mathrm{pred}}\!=\!t_{n}\!-\!t_{1}$ values are computed from the predictive data time points $t_{i}$.
The predicted data model values are
$g^{\prime}_{i}$ $=$ $g(t^{\prime}_{i},\bm{\beta}_{\mathrm{pred}}),$ (42)
where $t_{\mathrm{mid}}=t_{\mathrm{mid,pred}}$ and $\Delta T=\Delta T_{\mathrm{pred}}$. We do not compute “new” $t_{\mathrm{mid}}$ and $\Delta T$ values from the predicted data time points $t^{\prime}_{i}$ because the correct $g^{\prime}_{i}$ values are obtained only from the $\bm{\beta}_{\mathrm{pred}}$, $t_{\mathrm{mid,pred}}$ and $\Delta T_{\mathrm{pred}}$ combination of Equation 41. The $n^{\prime}$ predicted data model residuals
$\epsilon^{\prime}_{i}=y^{\prime}_{i}-g^{\prime}_{i}$
give the predicted data test statistic
$z_{\mathrm{pred}}$ $=$ z test statistic for predicted
data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ (Equation 15 or 16).
This parameter $z_{\mathrm{pred}}$ measures how well the prediction (Equation 42) obtained from the predictive data works for the predicted data. If the DCM detects a new signal from the predictive data, there are two alternatives:
  • If this new signal is real, the $z_{\mathrm{pred}}$ value of predicted data decreases.
  • If this new signal is unreal, the $z_{\mathrm{pred}}$ value of predicted data increases.
This “Prediction-test” technique revealed at least five real signals in the sunspot record (Jetsu 2025).
We compute the $z_{\mathrm{pred}}$ parameter value from the known predicted data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ values (Equation 3.10) . Predictions are possible even if all $t^{\prime}_{i}$, $y^{\prime}_{i}$ or $\sigma^{\prime}_{i}$ predicted data values are unknown. In this case, the $t_{i}$, $y_{i}$ and $\sigma_{i}$ values of all data can be used as predictive data. The best DCM model $g(t_{i},\bm{\beta}_{\mathrm{pred}})$ for all data determines the correct $\bm{\beta}_{\mathrm{pred}}$, $t_{\mathrm{mid,pred}}=(t_{n}+t_{1})/2$ and $\Delta T_{\mathrm{pred}}=t_{n}-t_{1}$ combination. The $n^{\prime}$ predicted data time points $t_{i}^{\prime}$ can be created for any arbitrary chosen time interval $\Delta T^{\prime}=t_{n}^{\prime}-t_{i}^{\prime}$. The $n^{\prime}$ predicted model $g_{i}^{\prime}=g^{\prime}_{i}(t_{i}^{\prime},\bm{\beta}_{\mathrm{pred}})$ values are obtained from Equation 42. These $g^{\prime}_{i}$ values can be used, for example, to compute the predicted mean level
$m_{\mathrm{pred}}={{1}\over{n^{\prime}}}\sum_{i=1}^{n^{\prime}}g^{\prime}_{i}$ (44)
during the chosen time interval $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$. Jetsu (2025) used this $m_{\mathrm{pred}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720 [31]. Since the known mean level of all sunspot data was $m$, Jetsu (2025) used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
  • If DCM detects a real new signal in all data, the $m_{\mathrm{pred}}$ value decreases.
  • If DCM detects many real signals in all data, the $m_{\mathrm{pred}}$ value falls below $m$.
  • If DCM detects an unreal new signal in all data, the $m_{\mathrm{pred}}$ value increases.
We use Model 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data (Table 3, column 2) to illustrate the DCM prediction technique. The black dots in Figures 8a-c are the first $n=40$ predictive data values $t_{i}$, $y_{i}$ and $\sigma_{i}$. The open black dots denote the $n^{\prime}=10$ predicted data values $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$. The continuous black line is the predictive model $g(t)$ and the dotted black line is the predicted model $g^{\prime}(t)$. The red dotted line shows the $\pm 3\sigma$ errors of both models. The blue dots are the predictive data $\epsilon_{i}=y_{i}-g_{i}$ residuals and the open blue dots are the predicted data $\epsilon_{i}^{\prime}=y_{i}^{\prime}-g_{i}^{\prime}$ residuals.
Model 3 is the sum of two $P_{1}=0.16$ and $P_{2}=0.17$ pure sine signals $(K_{1}=2,K_{2}=1)$ superimposed on the constant mean level $M_{0}=1$ $(K_{3}=0)$. We compute the $z$ and $z_{\mathrm{pred}}$ values for the $g_{1,1,0}$, $g_{2,1,0}$ and $g_{3,1,0}$ models (Table 10). These three models have the same $K_{2}=1$ and $K_{3}=0$ orders as Model 3, but their signal numbers $K_{1}=1,2$ or 3 are different, the $g_{2,1,0}$ model being the correct simulation Model 3.
The correct $g_{2,1,0}$ model gives the smallest $z_{\mathrm{pred}}=0.88$ value (Table 10). Therefore, it is a better predictive model than the $g_{1,1,0}$ and $g_{3,1,0}$ models. The one signal $g_{1,1,0}$ model prediction fails because the blue open circles denoting the predicted data residuals $\epsilon_{i}^{\prime}=y_{i}^{\prime}-g_{i}^{\prime}$ show large deviations from to the blue dotted line offset level of $\epsilon_{i}^{\prime}=-1$ (Figure 8a). The two signal $g_{2,1,0}$ model prediction succeeds because all open blue dots denoting the predicted data residuals stay close to the blue dotted line offset level $\epsilon_{i}^{\prime}=-1$, as well as inside the red dotted $\pm 3\sigma$ model error limits (Figure 8b).
The three signal model $g_{3,1,0}$ periods are $P_{1}=0.0643\pm 0.0018$, $P_{2}=0.1592\pm 0.0012$ and $P_{3}=0.1716\pm 0.0012$. The amplitudes of these pure sine signals are $A_{1}=0.092\pm 0.025$, $A_{2}=1.879\pm 0.16$ and $A_{3}=1.82\pm 0.15$. The periods and the amplitudes of the two strongest $P_{2}$ and $P_{3}$ signals are correct because they are the same as in Model 3 simulation (Table 3, column 1). Therefore, the $g_{3,1,0}$ prediction appears nearly as good as the $g_{2,1,0}$ prediction because the residuals $\epsilon_{i}^{\prime}$ are close to the offset level of $\epsilon_{i}^{\prime}=-1$ (blue open dots), and these residuals also stay inside the red dotted $\pm 3\sigma$ model error limits (Figure 8c). The amplitude $A_{1}=0.092$ of the third $P_{1}=0.0643$ signal is very low. Due this weak “unreal” $P_{1}$ signal, the $g_{3,1,0}$ model has a larger $z_{\mathrm{pred}}=0.91$ value than the correct $g_{2,1,0}$ model (Table 10).
The $n=40$ predictive data parameters $\eta_{1}=7$, $\eta_{2}=10$, $\chi_{1}=22.946$ and $\chi_{2}=19.790$ for the simple $g_{2,1,0}$ model and the complex $g_{3,1,0}$ model give Fisher-test values $F_{\chi}=1.32$ and $Q_{\mathrm{F}}=0.28>\gamma_{\mathrm{F}}=0.001$. The $H_{0}$ hypothesis is not rejected. For the predictive data, the Fisher-test also confirms that the $g_{2,1,0}$ model is better than the $g_{3,1,0}$ model. Finally, we note that the three signal model $g_{3,1,0}$ is not unstable ($\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”) although the simulated data contains only two signals.
For the $g_{2,1,0}$ model, the DCM detects the correct Model 3 signal frequencies (Table 3), but these frequencies are “too close” for the correct DFT detection (Equation 28). Therefore, the DFT prediction can not succeed.
4 Discussion
The solution for the non-linear DCM model is an ill-posed problem (Equations 1-5). We present a numerical solution that fulfills the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem. The DCM model is just one special case of a non-linear model. Our technique can be applied to solve other non-linear models: Use the free parameters that make the model non-linear (Equation 9: $\bm{\beta}_{I}$) for solving the remaining other free parameters (Equation 10: $\bm{\beta}_{II}$).
4.1 Existence ( $\mathrm{C}_{1}$)
[6] transformed the original function into the frequency domain. The modern DFT time series analysis method transforms the original data into the frequency domain and gives the best frequency for a pure sine model. [8] presented the least squares method, which minimizes the differences between the data and the linear model. Our DCM does the same by testing a large number of linear models. For every chosen DCM model, the number of tested linear models is
$\binom{n_{\mathrm{L}}}{K_{1}}+(1+n_{\mathrm{B}})\times\binom{n_{\mathrm{S}}}{K_{1}}$, (45)
where the number of tested long and short search frequencies is $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$, respectively. The number of bootstrap samples is $n_{\mathrm{B}}$. In other words, we solve this ill-posed problem by using brute numerical force. If the data contain only zero mean white noise, a linear least squares fit solution exists for every tested model.
4.2 Uniqueness ( $\mathrm{C}_{2}$)
We make the most of the Gauss-Markov theorem. When the numerical values of the tested frequencies $\bm{\beta}_{I}$ (Equation 9) are fixed, the model becomes linear and the solution for the other free parameters $\bm{\beta}_{II}$ is unique (Equation 10). All possible $\bm{\beta}_{I}$ frequency combinations are tested (Equation 11). For every tested frequency combination $\bm{\beta}_{I}$, the linear model gives a unique value for the test statistic $z$ (Equations  15 and 16). From all tested frequency combinations, we select the best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ which minimises $z$. The linear model for this best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ gives unique values for the remaining other free parameters $\bm{\beta}_{\mathrm{II,best}}$. The only goal for the massive DCM search (Equation 45) is to find these unique initial free parameter values $\bm{\beta}_{\mathrm{initial}}=[\bm{\beta}_{\mathrm{I,best}},\bm{\beta}_{\mathrm{II,best}}]$ for the non-linear iteration that gives the unique final free parameter values $\bm{\beta}_{\mathrm{final}}$ (Equation 21).
We use the Fisher-test to compare many different non-linear DCM models against each other. The selection criterion for the best model is unique (Equation 39). The best DCM model is not necessarily the correct model, if this correct model is not among the compared models. The correct model must be able the predict the future and past data. We formulate the Predictability-test for alternative DCM models (Equation 3.10). The order of Fisher- and Predictability-tests can be reversed. However, the former uses all data, while the latter uses a subset of all data, the predictive data. In ideal cases, both tests identify the same best and correct model, like in Section 3.10.
4.3 Stability ( $\mathrm{C}_{3}$)
The artificial bootstrap data samples (Equation 22) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation). We routinely check the stability of these solutions (e.g. Jetsu 2025, Figure S5). The unstable models, where the model parameter changes are large, are rejected (Section 2.1: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}” models).
There are additional signatures of stability. The $z$ periodogram solution is unique for every tested $\bm{\beta}_{I}$ frequency combination. If these periodograms are continuous and their changes are not irregular (e.g., like in Figs. 1a-b), the DCM model solution is stable because it does not change by increasing the number of tested frequencies $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$. Furthermore, the solutions for all seven simulated data samples converge when the $n$ and $\mathrm{SN}$} values increase (Tables 1 - 7). The different $n$ and $\mathrm{SN}$} combinations give the same stable DCM model solution.
We conclude that our numerical DCM model solution fulfills the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
4.4 Performance
DCM does not suffer from the DFT limitations (Equations 26-29). The simulated Models 1-7 expose these DFT weaknesses, but in all these cases, DCM detects all periodic signals superimposed on an unknown trend (Sections 3.1 - 3.7). The correct signal(-s) and trend are detected. If there is enough $(n)$ very accurate $(\sigma_{i})$ data $(y_{i})$, DCM can predict the future or postdict the past by finding the correct DCM model for these data. We do admit that the $n$ and $\mathrm{SN}$} values of Models 2, 5 and 7 are extreme and unrealistic for most cases of real data. However, those model simulations are necessary for demonstrating that DCM can detect signals from short $\Delta T$ samples of high quality data. Those DCM signal detections depend only on the number of observations and their accuracy.
The above-mentioned Model 2, 5 and 7 simulations also reveal that the relative accuracy of amplitude and period estimates is lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 2, 5 and 7). This statistical effect is the same when DCM is applied to the data of any arbitrary phenomenon. This effect would, for example, explain why our solar cycle amplitude predictions are less accurate than our solar cycle minimum and maximum epoch predictions (Jetsu 2025). DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct. This would explain the inexact signal periods that DFT has detected earlier from the sunspot data [e.g., 34].
We show that the DCM signal detections depend only on $n$ and $\mathrm{SN}$}, not on the time span $\Delta T$. This dimensional effect makes sense. If the data are inside an area $\Delta T\times\Delta y$, the ratio $y_{i}/\sigma_{i}$ reveals or hides the $g_{i}$ model details. DCM is an ideal forecasting method because the $\Delta T$ time span of data is irrelevant. This may motivate fast sampling of large high quality datasets and the use of numerical DCM parallel Python compution code in the analysis of those datasets.
[…]
Acknowledgements This work has made use of NASA’s Astrophysics Data System (ADS).
[…]
Author contribution Lauri Jetsu created the simulated datasets, performed the DCM and DFT analysis and wrote the manuscript.
[…]
Consent for publication The author of this manuscript grants permission for the publication in Statistics and Computing.
[…]
Data availability All data files and all DCM control files are stored to
[…]
https://zenodo.org/uploads/17018676.
There is also a manual of how to use those codes.
Declarations
[…]

### v2

@@
 The solution for the non-linear DCM model is an ill-posed problem.
-The backbone of DCM is the Gauss-Markov theorem that the least squares fit is the best unbiased estimator for linear regression models.
+The backbone of DCM is the Gauß-Markov theorem that the least squares fit is the best unbiased estimator for linear regression models.
 DCM is a simple numerical time series analysis method that performs a massive number of linear least squares fits.
-We show that our numerical solution for the DCM model fulfills the three conditions of a well-posed problem: existence, uniqueness and stability.
+Hence, the data spacing, even or uneven, is irrelevant.
+We show that our numerical solution for the DCM model fulfils the three conditions of a well-posed problem: existence, uniqueness and stability.
 The Fisher-test is used to identify the best DCM model from all alternative tested DCM models.
@@
 DCM is an ideal forecasting method because the time span of observations is irrelevant.
+If the Gauß-Markov theorem is valid, DCM can not fail.
 We recommend fast sampling of large high quality datasets and the analysis of those datasets using numerical DCM parallel computation Python code.
@@
 1 Introduction
+[21] formulated the least squares method.
+[11] connected this method to the principles of probability and the normal distribution.
+A few years later, he showed that the least-squares method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [12].
+The extended Gauß-Markov theorem states that the least squares method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [30].
+The Gauß-Markov theorem is the backbone of our Discrete Chi-Square Method (DCM).
+This time series analysis method performs a massive number of linear least squares fits to find the best model for the data.
 [14] defined the three conditions for a well-posed problem. The solution for the problem determines these conditions.
@@
   • $\mathrm{C}_{3}$ Stability: Small changes in the input data cause small changes in the solution.
-Stability means that the solution behaves predictably as the input varies. In other words, there exists a continuous mapping from the input space to the solution space.
+Stability means that the solution behaves predictably as the input varies.
+In other words, there exists a continuous mapping from the input space to the solution space.
 If any of these conditions are not satisfied, the problem is classified as ill-posed.
-Ill-posed problems are encountered in various fields of science [22].
-For example, Piskunov et al.
-[24] solves the stellar surface imaging inverse problem using the [29] regularization technique.
-Ill-posed problems are also encountered in time series analysis of astronomical data [30, 1].
-The problem is ill-posed because the unknown signal frequencies are free parameters of non-linear time series analysis models.
-The Lomb-Scargle periodogram is the most frequently applied time series analysis method in astronomy because the observations are often unevenly spaced in time [21, 27].
-Several extentions of this method have been published [35, 32].
-The DFT pre-whitening technique can detect many signals from unevenly spaced data [e.g., 26, 34].
-The data must be detrended before applying this technique, which detects one frequency at a time, until it detects no new frequency.
-In this study, artificial simulated data having equal weights is analysed using the DFT version published by Horne and Baliunas [11].
-We compare the performance of this DFT version to the performance of our DCM.
-[19] formulated the least squares method.
-[7] connected this method to the principles of probability and the normal distribution.
-A few years later, he showed that the least-squares method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [8].
-The extended Gauss-Markov theorem states that the least squares method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [33].
-This Gauss-Markov theorem is the backbone of our DCM.
+[…]
+[…]
+However, no published solution seems to exist for the more complicated cases, where the signal periods are longer than $\Delta T$ and/or these signals are not pure sines.
+We will show that DCM can find a well-posed numerical solution even for such complicated cases.
+This method can detect an unknown number of periodic signals superimposed on an unknown aperiodic trend.
+The noise level can be unknown because the performance of DCM depends only on the sample size $(n)$ and data accuracy $(\sigma)$.
+The time span $(\Delta T)$ is irrelevant.
+The data spacing, even or uneven, is also irrelevant.
+In this study, we apply one published DFT version [17] to analyse artificial simulated unevenly spaced data having equal weights.
+The performance of this DFT version is compared to the performance of our DCM.
 We present the DCM formulation in Section 2.1.
@@
 The observations are $y_{i}=y(t_{i})\pm\sigma(t_{i})$, where $t_{i}$ are the observing times and $\sigma_{i}$ are the errors (i=1,2,,n)(i=1,2,...,n).
-The time span of data is $\Delta T=t_{n}-t_{1}$.
+The time span of data, which is also called the data window, is $\Delta T=t_{n}-t_{1}$.
 The mid point is $t_{\mathrm{mid}}=t_{1}+\Delta T/2$.
@@
 $g(t)=g(t,K_{1},K_{2},K_{3})=h(t)+p(t),$ (1)
-where the integer values $K_{1}$, $K_{2}$ and $K_{3}$ are called the model orders [14].
+where the integer values $K_{1}$, $K_{2}$ and $K_{3}$ are called the model orders.
 The notation $g_{K_{1},K_{2},K_{3}}(t)$ is used to specify these orders.
@@
 $\bm{\beta}$ $=$ $[\beta_{1},\beta_{2},...,\beta_{\eta}]$
-$=$ $[B_{1,1},C_{1,1},f_{1},...,B_{K_{1},K_{2}},C_{K_{1},K_{2}},f_{K_{1}},M_{0},...,M_{K_{3}}].$
+$=$ $[B_{1,1},C_{1,1},f_{1},...,B_{K_{1},K_{2}},C_{K_{1},K_{2}},f_{K_{1}},$
+$M_{0},...,M_{K_{3}}].$
 The number of free parameters is
@@
 The solution for the second group of remaining free parameters
-$\bm{\beta}_{II}$ $=$ $[B_{1,1}C_{1,1},...,B_{K_{1},K_{2}},C_{K_{1},K_{2}},M_{0},...,M_{K_{3}}]$ (10)
+$\bm{\beta}_{II}$ $=$ $[B_{1,1}C_{1,1},...,B_{K_{1},K_{2}},C_{K_{1},K_{2}},$
+$M_{0},...,M_{K_{3}}]$
 becomes unique. This solution passes the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem.
@@
 where $n_{f}=n_{L}$ and $n_{f}=n_{S}$, respectively.
+The even or uneven data spacing is irrelevant because the linear least squares fit result for every tested frequency combination does not depend on this spacing.
 The global periodogram minimum
@@
 These frequencies are the unique initial values for the first group of free parameters $\bm{\beta}_{\mathrm{I,initial}}=$ $f_{\mathrm{1,best}},$..., $f_{\mathrm{K_{1},best}}]$ (Equation 9).
-The linear model for these constant $f_{\mathrm{1,best}},...,f_{\mathrm{K_{1},best}}]$ frequencies gives the unique initial values for the second group $\bm{\beta}_{\mathrm{II,initial}}$ of free parameters (Equation 10).
+The linear model for these constant $f_{\mathrm{1,best}},...,f_{\mathrm{K_{1},best}}]$ frequencies gives the unique initial values for the second group $\bm{\beta}_{\mathrm{II,initial}}$ of free parameters (Equation 2.1).
 The non-linear iteration
@@
 and the trend parameters
-  • $M_{k}$ = Polynomial coeffients.
+  • $M_{k}$ = Polynomial coefficients.
 We use the bootstrap technique to estimate the DCM model parameter errors [9].
 The tested bootstrap frequencies are the same as in the short search.
-Our TSPA- and CPS-methods applied the same bootstrap tehnique [15, 20].
 We select a random sample $\bm{\epsilon}^{*}$ from the residuals $\bm{\epsilon}$ of the DCM model (Equation 12).
@@
 For example, DCM can be forced search for too few or too many $K_{1}$ signals, or the selected $p(t)$ trend order $K_{3}$ can be wrong, as shown in Figures 5-10 by [19].
-Such DCM models are unstable and we denote them with $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”, like in [14] and Jetsu [^1] Jetsu L.
+Such DCM models are unstable and we denote them with $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”, like in Jetsu [^1] Jetsu L.
 “Do planets cause the sunspot cycle?”, submitted to Scientific Reports.
@@
 (f) DFT model $g_{\mathrm{DFT}}(t)$ (black continuous line), DFT trend $p_{\mathrm{DFT}}(t)$ (black dashed line) and data $y_{i}$ (black dots).
-(g) DFT model detrended $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (black contiuous line), DFT pure sine $s_{\mathrm{DFT}}(t)$ (red thick continuous line), detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and DFT model residuals (blue dots) offset to -1.5 level (blue dotted line).
+(g) DFT model detrended $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (black continuous line), DFT pure sine $s_{\mathrm{DFT}}(t)$ (red thick continuous line), detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and DFT model residuals (blue dots) offset to -1.5 level (blue dotted line).
 2.2 Simulated DCM model data samples
@@
 The first and last time point values are then modified to $t_{1}^{\star}=0$ and $t_{n}^{\star}=\Delta T$.
-Hence, the distance between independent frequencies [18, 27, 25] is always
+Hence, the distance between independent frequencies [20] is always
 $f_{0}=1/\Delta T=1.$ (23)
@@
 ${\mathrm{SN}}=(A/2)/\sigma.$ (24)
-2.3 Dicrete Fourier Transform (DFT)
+2.3 Discrete Fourier Transform (DFT)
 We search for signals in the simulated data using the DFT version formulated by [17]. This method searches for the best pure sine model for the data. The notations for the DFT model are
@@
 Our DFT analysis may fail for four main reasons.
-The data sample is “too short” [27, 28]
+The data sample is “too short” [20]
 $P>\Delta T.$ (26)
-The data contain “a trend” [3, 17]
+The data contain “a trend” [23]
 $p_{\mathrm{DFT}}(t)\neq 0.$ (27)
-The signal frequencies are “too close” [18, 23].
+The signal frequencies are “too close” [20].
 $|f_{1}-f_{2}|<f_{0}={\Delta T}^{-1}.$ (28)
-The signals are not “pure sines” [16, 2].
+The signals are not “pure sines” [5].
 $K_{2}\neq 1.$ (29)
-We search for many pure sine signals using the DFT pre-whitening technique [26, 34].
-The $p_{i}=p_{\mathrm{DFT}}(t_{i})$ trend is first removed from the data $y_{i}=y(t_{i})$.
-In the search for the first signal, we apply the DFT to the detrended $y_{i}-p_{i}$ data.
+The $p_{i}=p_{\mathrm{DFT}}(t_{i})$ trend is first removed from the data $y_{i}=y(t_{i})$ [23].
+Then, the technique described by Ghaderpour et al.
+[13] is used to detect the pure sine signals.
+We apply the DFT to the detrended $y_{i}-p_{i}$ data.
 When searching for the second signal, we apply the DTF to the original model residuals $\epsilon_{i}=y_{i}-g_{\mathrm{y,DFT}}(t_{i})$.
@@
               (f)            (h)
-Figure 7: Model 7 (Table 7: $n=100$, ${\mathrm{SN}}=5~000~000$ simulation). otations as in Figure 3. Best periods (diamonds) are explained in Section 3.7.
+Figure 7: Model 7 (Table 7: $n=100$, ${\mathrm{SN}}=5~000~000$ simulation). Notations as in Figure 3. Best periods (diamonds) are explained in Section 3.7.
 3.5 Model 5
@@
 However, the DFT two signal model is not the correct model for these Model 6 simulated data, which contain only one signal.
-If the correct period is $P$ and the correct model is a double wave $(K_{2}=2)$, the DCM pure sine model $(K_{1}=1)$ analysis can also give the values $P/2$ and $P$ [14].
+If the correct period is $P$ and the correct model is a double wave $(K_{2}=2)$, the DCM pure sine model $(K_{1}=1)$ analysis can also give the values $P/2$ and $P$.
 The DCM analysis succeeds for Model 6 simulated data. The DFT analysis fails.
@@
 (1) Simple model: Model number $\mathcal{M}$, model $g_{K_{1},K_{2},K_{3}}$, number of free parameters $\eta$, chi-square $\chi^{2}$ and control file name.
-(2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.
+(2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.22footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
 […]
-22footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
 3.8 Best model
@@
 Otherwise as in Table 8.
+33footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.
 (1) (2)
@@
 […]
-33footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.
 In the above example, the Fisher-test finds the correct number of signals $(K_{1})$ and the correct trend $(K_{3})$ for the pure sine signal alternative $(K_{2}=1)$.
 We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large.
-One example of testing the $K_{2}=1$ and $K_{2}=2$ signal models against each other can be found in Jetsu [14, Table 4].
+One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu (2025, Table S7).
 We conclude that the best model for the real data can be found by applying the Fisher-test to any arbitrary number of different nested DCM or DFT models.
@@
 For the constant, linear or parabolic $p(t)$ trend alternatives, the critical level for this $h_{1}(t)$ signal is $Q_{\mathrm{F}}<\!10^{-16}$.
-There is no need to discuss the DFT significance estimates [11, their Equation 22] because this method fails to detect the correct Model 1 period.
+There is no need to discuss the DFT significance estimates [17, their Equation 22] because this method fails to detect the correct Model 1 period.
 Table 10: DCM time series analysis between $P_{\mathrm{min}}=0.53$ and $P_{\mathrm{max}}=4.80$ for predictive data (Model = 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data: first forty observations).
@@
 3.10 Prediction
-There are numerous techniques for predicting (forecasting) a time series [e.g., 10, 12].
+There are numerous techniques for predicting (forecasting) a time series [e.g., 16, 18].
 The DCM model $g(t_{i},\bm{\beta})$ can be used to predict.
@@
 during the chosen time interval $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$.
-Jetsu (2025) used this $m_{\mathrm{pred}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720 [31].
+Jetsu (2025) used this $m_{\mathrm{pred}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720.
 Since the known mean level of all sunspot data was $m$, Jetsu (2025) used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
@@
 The solution for the non-linear DCM model is an ill-posed problem (Equations 1-5).
-We present a numerical solution that fulfills the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem.
+We present a numerical solution that fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem.
 The DCM model is just one special case of a non-linear model.
-Our technique can be applied to solve other non-linear models: Use the free parameters that make the model non-linear (Equation 9: $\bm{\beta}_{I}$) for solving the remaining other free parameters (Equation 10: $\bm{\beta}_{II}$).
+Our technique can be applied to solve other non-linear models: Use the free parameters that make the model non-linear (Equation 9: $\bm{\beta}_{I}$) for solving the remaining other free parameters (Equation 2.1: $\bm{\beta}_{II}$).
 4.1 Existence ( $\mathrm{C}_{1}$)
@@
 The modern DFT time series analysis method transforms the original data into the frequency domain and gives the best frequency for a pure sine model.
-[8] presented the least squares method, which minimizes the differences between the data and the linear model.
+[12] presented the least squares method, which minimises the differences between the data and the linear model.
 Our DCM does the same by testing a large number of linear models.
+The data spacing, even or uneven, is irrelevant for these models.
 For every chosen DCM model, the number of tested linear models is
@@
 4.2 Uniqueness ( $\mathrm{C}_{2}$)
-We make the most of the Gauss-Markov theorem.
-When the numerical values of the tested frequencies $\bm{\beta}_{I}$ (Equation 9) are fixed, the model becomes linear and the solution for the other free parameters $\bm{\beta}_{II}$ is unique (Equation 10).
+We make the most of the Gauß-Markov theorem.
+When the numerical values of the tested frequencies $\bm{\beta}_{I}$ (Equation 9) are fixed, the model becomes linear and the solution for the other free parameters $\bm{\beta}_{II}$ is unique (Equation 2.1).
 All possible $\bm{\beta}_{I}$ frequency combinations are tested (Equation 11).
@@
 The different $n$ and $\mathrm{SN}$} combinations give the same stable DCM model solution.
-We conclude that our numerical DCM model solution fulfills the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
+We conclude that our numerical DCM model solution fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
 4.4 Performance
@@
 DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct.
-This would explain the inexact signal periods that DFT has detected earlier from the sunspot data [e.g., 34].
+The leakage of DFT spectral power [20, 13] would explain why correct signal periods have not been detected earlier from the sunspot data.
 We show that the DCM signal detections depend only on $n$ and $\mathrm{SN}$}, not on the time span $\Delta T$.
@@
 DCM is an ideal forecasting method because the $\Delta T$ time span of data is irrelevant.
-This may motivate fast sampling of large high quality datasets and the use of numerical DCM parallel Python compution code in the analysis of those datasets.
+This may motivate fast sampling of large high quality datasets and the use of numerical DCM parallel Python computation code in the analysis of those datasets.
+4.5 Methodological contribution
+What is our methodological contribution?
+DCM has already been published and DFT has been applied in innumerable studies.
+We would get the same results if DCM were compared to any other time series analysis method than DFT.
+Practical time series analysis applications, like digital signal processing, require fast algorithms [20].
+Our statistical numerical solution for the non-linear time series analysis model requires a lot of CPU.
+Yet, the CPU is beside the point if a well-posed numerical solution can be found.
+DCM proceeds through two stages.
+For every chosen alternative model, DCM performs a massive number of linear least squares fits (Equation 18) to find the unique initial free parameter values for the non-linear iteration (Equation 21).
+This first stage devours CPU.
+The second fast stage, the Fisher-test, then reveals the best model alternative (Equation 39).
+DCM is an extremely simple time series analysis method that works in the time domain and uses real numbers.
+The more elaborate spectral estimating techniques, like DFT, work in the frequency domain and use complex numbers.
+We also want to draw attention to our idea of frequency space symmetry.
+This symmetry reduces the number of tested frequency combinations (Equation 11).
+For example, were this symmetry not exploited, the four signal $K_{1}=4$ search would give $4!=24$ same solutions in four-dimensional frequency space.
+The identification of the best frequency combination would be a daunting exercise, let alone the CPU consumed in testing all possible frequency combinations inside a tesseract (a four-dimensional cube).
+Our one-dimensional periodogram slices can "see inside” such multi-dimensional structures (Equation 20).
+Generally, different time series analysis methods suffer from different limitations. We list some of those limitations below.
+  1. 1.
+    The required data spacing is even or uneven.
+  2. 2.
+    The number of signals is unknown.
+  3. 3.
+    The presence and shape of trend are unknown.
+  4. 4.
+    The degree of noise is unknown.
+  5. 5.
+    The sample window causes leakage.
+  6. 6.
+    The leakage weakens frequency resolution.
+  7. 7.
+    The signal frequencies are too close.
+  8. 8.
+    The signal periods exceed the data time span.
+  9. 9.
+    The signal shapes are not pure sinusoids.
+  10. 10.
+    The forecasts of non-linear models fail.
+DCM does not suffer from any of these limitations. This is our methodological contribution.
+We hope that our results are considered to represent advances in the field of time series analysis. Our DCM can not fail if the Gauß-Markov theorem is valid.
 […]

### v3

@@
-Discrete Fourier Transform versus Discrete Chi-Square Method
+Discrete Chi-Square Method beats Discrete Fourier Transform and other similar parametric time series analysis methods
 Abstract
 We compare two time series analysis methods, the Discrete Fourier Transform (DFT) and our Discrete Chi-square Method (DCM).
-DCM is designed for detecting many signals superimposed on an unknown trend.
-The solution for the non-linear DCM model is an ill-posed problem.
+DCM is designed for detecting signal(-s) superimposed on an unknown trend.
+The analytical solution for the non-linear DCM model is an ill-posed problem.
+We present a computational statistical well-posed solution for this problem.
 The backbone of DCM is the Gauß-Markov theorem that the least squares fit is the best unbiased estimator for linear regression models.
-DCM is a simple numerical time series analysis method that performs a massive number of linear least squares fits.
+DCM can not fail because this simple time series analysis method computes a massive number of linear least squares fits.
 Hence, the data spacing, even or uneven, is irrelevant.
-We show that our numerical solution for the DCM model fulfils the three conditions of a well-posed problem: existence, uniqueness and stability.
-The Fisher-test is used to identify the best DCM model from all alternative tested DCM models.
-The correct DCM model must also pass our Predictivity-test.
-Our analyses of seven different simulated data samples expose the weaknesses of DFT and the efficiency of DCM.
-The DCM signal and trend detection depend only on the sample size and the accuracy of data.
-DCM is an ideal forecasting method because the time span of observations is irrelevant.
-If the Gauß-Markov theorem is valid, DCM can not fail.
-We recommend fast sampling of large high quality datasets and the analysis of those datasets using numerical DCM parallel computation Python code.
+We use the Fisher-test to identify the best DCM model from all alternative tested DCM models.
+This best model must also pass our Predictivity-test.
+Our analyses of simulated complex data sets expose the weaknesses of DFT and the efficiency of DCM.
+The DFT and DCM are frequency-domain parametric methods.
+There are many other similar parametric methods.
+Just like DFT, those other methods suffer from their own particular application limitations.
+The list of those limitations is long.
+DCM suffers from none of those limitations.
+The performance of DCM depends only on the data set size and accuracy.
+DCM is an ideal forecasting method because the data set time span $(\Delta T)$ is irrelevant.
+It does not matter how long $(\Delta T)$ and/or complex the data set is because DCM will inevitably detect the signal(-s) and the trend when the data set size $(n)$ and/or accuracy $(\sigma)$ become adequate.
 keywords:
-Ill-posed problems, Time series analysis, Non-linear models, Discrete Chi-square Method (DCM), Discrete Fourier Transform (DFT)
+Computational statistics, Ill-posed problems, Non-linear time series analysis models, Discrete Chi-square Method, Discrete Fourier Transform
 1 Introduction
@@
 The extended Gauß-Markov theorem states that the least squares method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [34].
-The Gauß-Markov theorem is the backbone of our Discrete Chi-Square Method (DCM).
+The Gauß-Markov theorem is the backbone of our DCM.
 This time series analysis method performs a massive number of linear least squares fits to find the best model for the data.
@@
 If any of these conditions are not satisfied, the problem is classified as ill-posed.
-[…]
-[…]
-However, no published solution seems to exist for the more complicated cases, where the signal periods are longer than $\Delta T$ and/or these signals are not pure sines.
-We will show that DCM can find a well-posed numerical solution even for such complicated cases.
+Ill-posed problems arise in many fields of science [27, 30, 25].
+Typical examples are partial differential equation solutions [17], nonlinear parameter estimation [1] and regularisation of inverse problems [9, 33].
+Ill-posed problems are encountered in time series analysis [3, 5] because the models for the data are often non-linear [32, 31].
+The unknown free parameters of the model, namely the frequencies, are in the arguments of trigonometric functions representing the signals.
+This causes non-linearity because the model partial derivatives contain free parameters of the model.
+There are numerous techniques for solving the ill-posed problem of detecting one signal superimposed on a trend, especially for evenly spaced data [7, 28, 2].
+Kay and Marple [22] compared many spectral estimating techniques that can detect more than one signal from evenly spaced data.
+They discussed how the time span of data $(\Delta T)$ limits the frequency resolution for detecting many signals and causes leakage that shifts power from one spectral peak to another.
+The spectral estimating techniques fail if the data are non-stationary.
+Therefore, trends changing the sample mean and/or variance must be removed before applying spectral estimation [26].
+Ghaderpour et al.
+[15] presented different techniques for reducing spectral leakage.
+They also discussed techniques for detecting many signals in unevenly spaced data, if these signals are pure sines.
+However, no published solution seems to exist for the more complex data sets, where the signal periods are longer than $\Delta T$ and/or these signals are not pure sines.
+We will show that DCM can find a well-posed computational statistical solution even for such complex data sets.
 This method can detect an unknown number of periodic signals superimposed on an unknown aperiodic trend.
@@
 We present the DCM formulation in Section 2.1.
-Data samples are simulated using the DCM model (Sect.
+Data sets are simulated using the DCM model (Sect.
 2.2).
-The four DFT limitations are discussed in Section 2.3 (Equations 26-29).
-DCM and DFT time series analysis of the simulated data samples shows that DCM does not suffer from the DFT limitations (Sects.
+The four DFT application limitations are discussed in Section 2.3 (Equations 27-30).
+DCM and DFT time series analysis of the simulated data sets shows that DCM does not suffer from the DFT application limitations (Sects.
 3.1-3.7).
@@
 In the above-mentioned $K_{1}=2$ map, the slice $z_{1}(f_{1})$ would represent the height $z$ at the location $(f_{1},f_{\mathrm{2,best}})$ when moving along the straight constant line $f_{2}=f_{\mathrm{2,best}}$ that crosses the global minimum $z_{\mathrm{min}}$ (Equation 19) at the coordinate point $(f_{\mathrm{1,best}},f_{\mathrm{2,best}})$.
+These one-dimensional periodogram slices (Equation 20) allow us "to see inside” the multi-dimensional structure of the DCM test statistic $z$ (Equations 15 and 16).
+This kind of visualisation is important in time series analysis [29].
 The short search gives the best $f_{\mathrm{1,best}},$..., $f_{\mathrm{K_{1},best}}$ frequencies for the data.
@@
 gives the final free parameter values $\bm{\beta}_{\mathrm{final}}$.
-DCM determines the signal parameters
+Furlan and Mortarino [11] emphasise that the analytical error estimates for the non-linear model free parameters can be tricky.
+They compare different analytical free parameter error estimating methods by using the computational statistical bootstrap technique [10].
+They conclude that the analytical error estimates become less reliable if the number of free parameters increases.
+For our non-linear DCM model, we determine the $i$:th signal parameters
   • $P_{i}=1/f_{i}=$ Period
@@
   • $M_{k}$ = Polynomial coefficients.
-We use the bootstrap technique to estimate the DCM model parameter errors [9].
-The tested bootstrap frequencies are the same as in the short search.
+Of these model parameters, all $P_{i}=1/f_{i}$ and all $M_{k}$ estimates are among the free parameters $\bm{\beta}$.
+However, the $A_{i}$, $t_{\mathrm{i,min,1}}$, $t_{\mathrm{i,min,2}}$, $t_{\mathrm{i,max,1}}$ and $t_{\mathrm{i,max,2}}$ model parameters depend on $\bm{\beta}$ values.
+The analytical solutions for these model parameters are simple for pure sines $(K_{2}=1)$, but become quite complicated for double waves $(K_{2}=2)$.
+Clearly, the analytical solution for the errors of DCM model parameters would be tedious.
+The analytical solution for the model error
+$g(t)\pm\sigma_{g(t)}$ (22)
+would be practically impossible because the number of DCM model free parameters is large, especially if the data contains many signals.
+We use the computational statistical bootstrap technique [10] to solve these analytical problems.
+This approach gives the DCM model parameter error estimates, as well as the $\sigma_{g(t)}$ model error estimates (Equation 22).
+In our bootstrap, the tested frequencies are the same as in the short search.
 We select a random sample $\bm{\epsilon}^{*}$ from the residuals $\bm{\epsilon}$ of the DCM model (Equation 12).
 Any $\epsilon_{i}$ value can be chosen to the $\bm{\epsilon}^{*}$ sample as many times as the random selection happens to favour it.
-We create numerous artificial bootstrap data samples
+We create numerous artificial bootstrap data sets
 $\bm{y}^{*}=\bm{g}+\bm{\epsilon}^{*}.$ (23)
-Each $\bm{y}^{*}$ sample gives one estimate for every DCM model parameter.
-For each particular model parameter, the bootstrap error estimate is the standard deviation of all estimates obtained from all $\bm{y}^{*}$ samples.
+Each $\bm{y}^{*}$ sample gives one estimate for every DCM model parameter value and model $g(t)$ value.
+For each particular model parameter and each particular $g(t)$, the bootstrap error estimate is the standard deviation of all estimates obtained from all $\bm{y}^{*}$ samples.
 There are, of course, totally wrong DCM models for the data.
@@
 2.2 Simulated DCM model data samples
-We use seven different $g(t)$ models to simulate 21 artificial data samples (Sections 3.1-3.7).
-The time interval of all simulated samples is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$.
+We use seven different $g(t)$ models to simulate 21 artificial complex data sets (Sections 3.1-3.7).
+The time interval of all simulated data sets is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$.
 The first and last time point values are then modified to $t_{1}^{\star}=0$ and $t_{n}^{\star}=\Delta T$.
@@
 The pure sine signals for the $y_{i}$ data and the $\epsilon_{i}$ residuals are denoted with $s_{y,\mathrm{DFT}}(t)$ and $s_{\epsilon,\mathrm{DFT}}(t)$, respectively.
-Our DFT analysis may fail for four main reasons.
+Our DFT analysis may fail due to four application limitations.
 The data sample is “too short” [22]
@@
 3 Results
-We apply DCM and DFT to seven different simulated data samples (Sections 3.1 - 3.7). DCM analysis succeeds for all samples. DFT analysis fails for every sample.
+We apply DCM and DFT to seven different simulated data sets (Sections 3.1 - 3.7).
+As we proceed, these data sets become increasingly complex.
+DCM analysis succeeds for all data sets.
+DFT analysis fails for every data set.
 3.1 Model 1
@@
 3.7 Model 7
-Our seventh simulation model is
+In this section, we analyse our most complex data set. The seventh simulation model is
 $g(t)$ $=$ $c_{1}\cos{\left[{{2\pi t}\over{P_{1}}}\right]}+c_{2}\cos{\left[{{4\pi(t-c_{3})}\over{P_{1}}}\right]}$
@@
 These signals are double waves having two unequal minima and maxima.
-All four main reasons that can cause the failure of DFT analysis are present (Equations 26-29).
+All four main reasons that can cause the failure of DFT analysis are present (Equations 27-30).
+Therefore, this simulated data set is the most complex one of all analysed seven data sets.
 Our DCM and DFT time series analysis of Model 7 simulated data is performed between $P_{\mathrm{min}}=P_{1}/3=0.4$ and $P_{\mathrm{max}}=3P_{1}=3.6$.
@@
 These DCM model residuals are small and their level is stable.
-These results confirm that the DCM time series analysis method can detect complex non-linear models $(\eta=13)$ from very small data samples $(n=100)$, if these data are extremely accurate $({\mathrm{SN}}=5~000~000).$
+These results confirm that the DCM time series analysis method can detect complex non-linear models $(\eta=13)$ from very small data sets $(n=100)$, if these data are extremely accurate $({\mathrm{SN}}=5~000~000).$
 The DFT time series analysis gives the wrong periods for the original data (Figure 7e: diamond at $P_{1}=0.542$ and for the residuals (Figure 7f: diamond at $P_{2}=1.059$.
@@
 The correct simulated peak to peak amplitude for this second signal is much larger, $A_{2}=2.$ The DFT model residuals (blue dots) are offset to -1.5 level and show strong trends.
-The DCM time series analysis succeeds for Model 7 simulated data. The DFT time series analysis fails, as predicted by Equations 26-29.
+The DCM time series analysis succeeds for Model 7 simulated data. The DFT time series analysis fails, as predicted by Equations 27-30.
 Table 8: Fisher tests.
@@
 If the number of complex model free parameters $\eta_{2}$ increases, the $R_{2}$ or $\chi^{2}_{2}$ values decrease.
-This increases the $F_{R}$ or $F_{\chi}$ values because the terms $(R_{1}/R_{2}-1)$ or $(\chi^{2}_{1}/\chi^{2}_{2}-1)$ (Equations 37 and 38: first terms).
-At the same time, the $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty term decreases (Equations 37 and 38: second terms), and this decreases the $F_{R}$ or $F_{\chi}$ values.
+This increases the $F_{R}$ or $F_{\chi}$ values because the terms $(R_{1}/R_{2}-1)$ or $(\chi^{2}_{1}/\chi^{2}_{2}-1)$ increase (Equations 38 and 39: first terms).
+At the same time, the $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty term decreases (Equations 38 and 39: second terms), and this decreases the $F_{R}$ or $F_{\chi}$ values.
 This penalty term prevents the use of too high $\eta_{2}$ values (too complex models).
@@
 For the constant, linear or parabolic $p(t)$ trend alternatives, the critical level for this $h_{1}(t)$ signal is $Q_{\mathrm{F}}<\!10^{-16}$.
-There is no need to discuss the DFT significance estimates [17, their Equation 22] because this method fails to detect the correct Model 1 period.
+There is no need to discuss the DFT significance estimates [19, their Equation 22] because this method fails to detect the correct Model 1 period.
 Table 10: DCM time series analysis between $P_{\mathrm{min}}=0.53$ and $P_{\mathrm{max}}=4.80$ for predictive data (Model = 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data: first forty observations).
@@
 Continuous black line of predictive data model $g(t,\bm{\beta}_{\mathrm{pred}})$ ends to vertical dotted blue line, where dotted black line of predicted model $g^{\prime}(t,\bm{\beta}_{\mathrm{pred}})$ begins.
-Dotted red line denotes $\pm 3\sigma$ errors of both models.
+Dotted red line denotes $\pm 3\sigma_{g(t)}$ errors of both models (Equation 22).
 Residuals of predictive data model $\epsilon_{i}=y_{i}-g_{i}$ (blue dots), residuals of predicted data model $\epsilon_{i}=^{\prime}y_{i}^{\prime}-g_{i}^{\prime}$ (open blue dots) and $\pm 3\sigma$ errors of both models (red dotted line) are offset to level -1 (blue dotted line).
@@
 3.10 Prediction
-There are numerous techniques for predicting (forecasting) a time series [e.g., 16, 18].
+There are numerous techniques for predicting (forecasting) a time series [e.g., 18, 20, 23].
 The DCM model $g(t_{i},\bm{\beta})$ can be used to predict.
@@
 Therefore, the DFT prediction can not succeed.
+The analyses of all seven complex data sets indicate that the relative accuracy of amplitude and period estimates are lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 1 - 7).
+This statistical effect is the same when DCM is applied to the data of any arbitrary phenomenon.
+This effect would, for example, explain why our solar cycle amplitude predictions are less accurate than our solar cycle minimum and maximum epoch predictions (Jetsu 2025).
+DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct.
+The leakage of DFT spectral power [22, 15] would explain why correct signal periods have not been detected earlier from the sunspot data.
 4 Discussion
 The solution for the non-linear DCM model is an ill-posed problem (Equations 1-5).
-We present a numerical solution that fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem.
+We present a computational statistical solution that fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem.
 The DCM model is just one special case of a non-linear model.
@@
 The number of bootstrap samples is $n_{\mathrm{B}}$.
-In other words, we solve this ill-posed problem by using brute numerical force.
-If the data contain only zero mean white noise, a linear least squares fit solution exists for every tested model.
+In other words, we solve this ill-posed problem by using brute computational force.
+If the data contain only zero mean white noise, the Gauß-Markov theorem ensures that a linear least squares fit solution exists for every tested frequency combination.
 4.2 Uniqueness ( $\mathrm{C}_{2}$)
@@
 4.3 Stability ( $\mathrm{C}_{3}$)
-The artificial bootstrap data samples (Equation 22) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation).
+The artificial bootstrap data sets (Equation 23) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation).
 We routinely check the stability of these solutions (e.g.
@@
 1a-b), the DCM model solution is stable because it does not change by increasing the number of tested frequencies $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$.
-Furthermore, the solutions for all seven simulated data samples converge when the $n$ and $\mathrm{SN}$} values increase (Tables 1 - 7).
+Furthermore, the solutions for all seven simulated data sets converge when the $n$ and $\mathrm{SN}$} values increase (Tables 1 - 7).
 The different $n$ and $\mathrm{SN}$} combinations give the same stable DCM model solution.
-We conclude that our numerical DCM model solution fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
+We conclude that our computational statistical DCM model solution fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
 4.4 Performance
-DCM does not suffer from the DFT limitations (Equations 26-29).
-The simulated Models 1-7 expose these DFT weaknesses, but in all these cases, DCM detects all periodic signals superimposed on an unknown trend (Sections 3.1 - 3.7).
+DCM does not suffer from the DFT limitations (Equations 27-30).
+The simulated complex data sets of Models 1-7 expose these DFT weaknesses, but in all these cases, DCM detects all periodic signals superimposed on an unknown trend (Sections 3.1 - 3.7).
 The correct signal(-s) and trend are detected.
@@
 Those DCM signal detections depend only on the number of observations and their accuracy.
-The above-mentioned Model 2, 5 and 7 simulations also reveal that the relative accuracy of amplitude and period estimates is lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 2, 5 and 7).
-This statistical effect is the same when DCM is applied to the data of any arbitrary phenomenon.
-This effect would, for example, explain why our solar cycle amplitude predictions are less accurate than our solar cycle minimum and maximum epoch predictions (Jetsu 2025).
-DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct.
-The leakage of DFT spectral power [20, 13] would explain why correct signal periods have not been detected earlier from the sunspot data.
-We show that the DCM signal detections depend only on $n$ and $\mathrm{SN}$}, not on the time span $\Delta T$.
-This dimensional effect makes sense.
-If the data are inside an area $\Delta T\times\Delta y$, the ratio $y_{i}/\sigma_{i}$ reveals or hides the $g_{i}$ model details.
-DCM is an ideal forecasting method because the $\Delta T$ time span of data is irrelevant.
-This may motivate fast sampling of large high quality datasets and the use of numerical DCM parallel Python computation code in the analysis of those datasets.
-4.5 Methodological contribution
-What is our methodological contribution?
-DCM has already been published and DFT has been applied in innumerable studies.
-We would get the same results if DCM were compared to any other time series analysis method than DFT.
-Practical time series analysis applications, like digital signal processing, require fast algorithms [20].
-Our statistical numerical solution for the non-linear time series analysis model requires a lot of CPU.
-Yet, the CPU is beside the point if a well-posed numerical solution can be found.
+Our DCM computes the $R$ (Equation 13) or $\chi^{2}$ (Equation 14) values for a massive number of linear models (Equation 46).
+The Gauß-Markov theorem ensures that the model having the lowest $R$ or $\chi^{2}$ is eventually always found, and the result for the non-linear iteration is unique (Equation 21).
+We could have simulated more complex data sets by increasing the values of $K_{1}$ (number of signals), $K_{2}$ (signal shape) and $K_{3}$ (polynomial order of trend).
+DCM would eventually detect those signals and the trend when the sampled data sets get larger $(n)$ and/or the signal to noise ratio becomes higher $({\mathrm{SN}})$.
+In this sense, DCM has no limit for the complexity of the data sets.
+In the real world, it may be impossible to eliminate noise $({\mathrm{SN}})$, but it is possible increase the sample size $(n)$.
+Since the time span $(\Delta T)$ is irrelevant, the best solution is fast sampling of accurate data sets.
+For example, many observers could measure the same phenomenon simultaneously.
+We recommend such fast sampling combined to DCM parallel computation Python code analysis.
+We want to draw attention to three aspects of DCM computational performance.
+  1. 1.
+    Practical time series analysis applications, like digital signal processing, require fast computational algorithms [22].
+Our computational statistical solution for the non-linear DCM model requires a lot of CPU.
+Yet, the CPU used is beside the point if a well-posed computational solution can be found to any challenging scientific non-linear problem.
 DCM proceeds through two stages.
@@
 The second fast stage, the Fisher-test, then reveals the best model alternative (Equation 40).
-DCM is an extremely simple time series analysis method that works in the time domain and uses real numbers.
-The more elaborate spectral estimating techniques, like DFT, work in the frequency domain and use complex numbers.
-We also want to draw attention to our idea of frequency space symmetry.
-This symmetry reduces the number of tested frequency combinations (Equation 11).
-For example, were this symmetry not exploited, the four signal $K_{1}=4$ search would give $4!=24$ same solutions in four-dimensional frequency space.
-The identification of the best frequency combination would be a daunting exercise, let alone the CPU consumed in testing all possible frequency combinations inside a tesseract (a four-dimensional cube).
-Our one-dimensional periodogram slices can "see inside” such multi-dimensional structures (Equation 20).
-Generally, different time series analysis methods suffer from different limitations. We list some of those limitations below.
+  2. 2.
+    Our frequency space symmetry idea dramatically reduces CPU (Equation 11).
+For example, were this symmetry not exploited, the four signal $K_{1}=4$ search would give $4!=24$ exactly the same solutions from four-dimensional tested grid in frequency space.
+This would cause unnecessary use of CPU in the preliminary testing all possible frequency combinations inside a tesseract (a four-dimensional cube).
+After one necessary iteration of Equation 21, and 23 unnecessary ones, the identification of the best frequency combination would be a daunting exercise.
+  3. 3.
+    The DCM signal detections depend only on the sample size $n$ and the signal to noise ratio $\mathrm{SN}$}, but not on the time span $\Delta T$.
+This dimensional effect makes sense.
+If the $t_{i}$ and $y_{i}$ data are inside a rectangle $\Delta T\times\Delta y=(t_{n}-t_{1})\times[{\mathrm{max}}(y_{i})-{\mathrm{min}}(y_{i})]$, then the $y_{i}/\sigma_{i}$ ratio reveals or hides the $g_{i}$ model details.
+Since the time span of data $\Delta T$ is irrelevant, DCM is an ideal forecasting method.
+We have been deluded into unnecessary waiting for the repetition of the signal(-s).
+DCM can perceive the future and the past much “earlier” (from shorter $\Delta T$) than previously thought, like in Equations 27 and 29.
+This should motivate fast sampling of large high quality data sets and their analysis using the DCM parallel Python code.
+The invention of computers has made it possible to use this computational statistical method, which is probably the simplest possible one.
+The analytical solution for the non-linear DCM model is ill-posed (Equation 1).
+Regardless of the data set complexity, DCM can find a well-posed computational statistical solution for this problem (Equation 21).
+So far, we have compared DCM only to DFT.
+Both methods are parametric and operate in the frequency-domain.
+There are innumerable similar parametric time series analysis methods [e.g.
+22, 15].
+Each one of these other time series analysis methods suffers from their own particular application limitations.
+We list some of those application limitations below.
   1. 1.
@@
     The forecasts of non-linear models fail.
-DCM does not suffer from any of these limitations. This is our methodological contribution.
-We hope that our results are considered to represent advances in the field of time series analysis. Our DCM can not fail if the Gauß-Markov theorem is valid.
+DCM does not suffer from any of these application limitations. This simple computational statistical method can fail if, and only if, the Gauß-Markov theorem is not valid. That is impossible.
 […]
@@
 […]
-Consent for publication The author of this manuscript grants permission for the publication in Statistics and Computing.
+Consent for publication The author of this manuscript grants permission for the publication in Computational Statistics.
 […]

### v4

@@
-Discrete Chi-Square Method beats Discrete Fourier Transform and other similar parametric time series analysis methods
+Discrete Chi-Square Method can model and forecast complex time series
 Abstract
-We compare two time series analysis methods, the Discrete Fourier Transform (DFT) and our Discrete Chi-square Method (DCM).
-DCM is designed for detecting signal(-s) superimposed on an unknown trend.
+We show how intensive, large and accurate time series can allow us to see through time.
+Many phenomena have aperiodic and periodic components.
+An ideal time series analysis method would detect such trend and signal(-s) combinations.
+The widely-used Discrete Fourier Transform (DFT) and other frequency-domain parametric time series analysis methods have many application limitations constraining the trend and signal(-s) detection.
+We show that none of those limitations constrains our Discrete Chi-square Method (DCM) which can detect signal(-s) superimposed on an unknown trend.
+Our simulated time series analyses ascertain the revolutionary Window Dimension Effect (WD-effect): “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.” The simulations also expose the DFT’s weaknesses and the DCM’s efficiency.
+The DCM’s backbone is the Gauß-Markov theorem that the Least Squares (LS) is the best unbiased estimator for linear regression models.
+DCM can not fail because this simple method is based on the computation of a massive number of linear model LS fits.
+The Fisher-test gives the signal significance estimates and identifies the best DCM model from all alternative tested DCM models.
 The analytical solution for the non-linear DCM model is an ill-posed problem.
-We present a computational statistical well-posed solution for this problem.
-The backbone of DCM is the Gauß-Markov theorem that the least squares fit is the best unbiased estimator for linear regression models.
-DCM can not fail because this simple time series analysis method computes a massive number of linear least squares fits.
-Hence, the data spacing, even or uneven, is irrelevant.
-We use the Fisher-test to identify the best DCM model from all alternative tested DCM models.
-This best model must also pass our Predictivity-test.
-Our analyses of simulated complex data sets expose the weaknesses of DFT and the efficiency of DCM.
-The DFT and DCM are frequency-domain parametric methods.
-There are many other similar parametric methods.
-Just like DFT, those other methods suffer from their own particular application limitations.
-The list of those limitations is long.
-DCM suffers from none of those limitations.
-The performance of DCM depends only on the data set size and accuracy.
-DCM is an ideal forecasting method because the data set time span $(\Delta T)$ is irrelevant.
-It does not matter how long $(\Delta T)$ and/or complex the data set is because DCM will inevitably detect the signal(-s) and the trend when the data set size $(n)$ and/or accuracy $(\sigma)$ become adequate.
+We present a computational well-posed solution.
+The DCM can forecast complex time series.
+The best DCM model must be correct if it passes our Forecast-test.
+Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series.
 keywords:
-Computational statistics, Ill-posed problems, Non-linear time series analysis models, Discrete Chi-square Method, Discrete Fourier Transform
+Time series analysis, Non-linear and non-stationary processes, Ill-posed problems, Computational statistics, Forecasting, Discrete Chi-square Method, Discrete Fourier Transform
 1 Introduction
-[24] formulated the least squares method.
+Cheng et al.
+[7] stated, “Forecasting the evolution of complex systems is noted as one of the 10 grand challenges of modern science.” We will show that the DCM can detect the signal(-s) and the trend in complex time series, and can also forecast the time series evolution.
+Our DCM is based on the LS method, which was originally formulated by [28].
 [14] connected this method to the principles of probability and the normal distribution.
-A few years later, he showed that the least-squares method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [14].
-The extended Gauß-Markov theorem states that the least squares method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [34].
+A few years later, he showed that the LS method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [15].
+The extended Gauß-Markov theorem states that the LS method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [43].
 The Gauß-Markov theorem is the backbone of our DCM.
-This time series analysis method performs a massive number of linear least squares fits to find the best model for the data.
+This time series analysis method performs a massive number of LS fits to find the best model for the data.
 [18] defined the three conditions for a well-posed problem. The solution for the problem determines these conditions.
@@
 This causes non-linearity because the model partial derivatives contain free parameters of the model.
-There are numerous techniques for solving the ill-posed problem of detecting one signal superimposed on a trend, especially for evenly spaced data [7, 28, 2].
-Kay and Marple [22] compared many spectral estimating techniques that can detect more than one signal from evenly spaced data.
-They discussed how the time span of data $(\Delta T)$ limits the frequency resolution for detecting many signals and causes leakage that shifts power from one spectral peak to another.
+Our DCM is a frequency-domain parametric time series analysis method.
+Here, we discuss shortly such methods in the order of increasing model complexity.
+Numerous methods can be applied if the time series is stationary (constant mean and variance).
+The Fast Fourier Transform for evenly spaced data [8] and the DFT for unevenly spaced data [30, 36] are the most widely-used methods.
+The FFT and DFT assume that the correct model for the time series is a pure sine.
 The spectral estimating techniques fail if the data are non-stationary.
 Therefore, trends changing the sample mean and/or variance must be removed before applying spectral estimation [33].
+There are numerous methods for detecting one signal superimposed on a trend, especially for evenly spaced data [6, 37, 2].
+Kay and Marple [25] compared many spectral estimating techniques that can detect more than one signal from evenly spaced data.
+They discussed how the sample window $(\Delta T)$ limits the frequency resolution for detecting many signals and causes leakage that shifts power from one spectral peak to another.
+Exactly correct frequency values may not be detected because the leakage can shift the periodogram peaks.
 Ghaderpour et al.
@@
 They also discussed techniques for detecting many signals in unevenly spaced data, if these signals are pure sines.
-However, no published solution seems to exist for the more complex data sets, where the signal periods are longer than $\Delta T$ and/or these signals are not pure sines.
-We will show that DCM can find a well-posed computational statistical solution even for such complex data sets.
-This method can detect an unknown number of periodic signals superimposed on an unknown aperiodic trend.
-The noise level can be unknown because the performance of DCM depends only on the sample size $(n)$ and data accuracy $(\sigma)$.
-The time span $(\Delta T)$ is irrelevant.
-The data spacing, even or uneven, is also irrelevant.
-In this study, we apply one published DFT version [19] to analyse artificial simulated unevenly spaced data having equal weights.
-The performance of this DFT version is compared to the performance of our DCM.
-We present the DCM formulation in Section 2.1.
-Data sets are simulated using the DCM model (Sect.
-2.2).
-The four DFT application limitations are discussed in Section 2.3 (Equations 27-30).
-DCM and DFT time series analysis of the simulated data sets shows that DCM does not suffer from the DFT application limitations (Sects.
-3.1-3.7).
-We then present the identification of the best DCM model for the data (Sect.
-3.8), the DCM significance estimates (Sect.
-3.9) and the DCM predictions (Sect.
-3.10).
-Finally, we discuss our results (Sect.
-4).
+There are more complex modelling problems.
+For example, the signals are not necessarily pure sines or the sample window $\Delta T$ is shorter than the signal period(-s).
+Reliable forecasting is the ultimate test for any time series analysis method [7].
+Forecasting can work for linear and stationary processes, but it is challenging for non-linear or non-stationary ones.
+Other forecasting challenges are overfitting and forecast error estimation [32, 27, 16, 34].
+Our DCM is a frequency-domain parametric time series analysis method.
+Such methods have their own particular limitations that constrain their applications.
+Here is our list of those application limitations (AL).
+  1. 1.
+    Data errors (level of noise) are unknown.
+  2. 2.
+    Data error information is not utilised.
+  3. 3.
+    Data must be evenly spaced.
+  4. 4.
+    Model parameter errors are unknown.
+  5. 5.
+    Model and forecast errors are unknown.
+  6. 6.
+    Sample window is shorter than signal period(s).
+  7. 7.
+    Presence and shape of trend are unknown.
+  8. 8.
+    Sample window causes leakage.
+  9. 9.
+    Leakage weakens frequency resolution.
+  10. 10.
+    Signal shapes are not pure sines.
+  11. 11.
+    Number of signals is unknown.
+  12. 12.
+    Correct model alternative is unknown.
+  13. 13.
+    Signal significances are unknown.
+  14. 14.
+    Model solution is ill-posed.
+  15. 15.
+    Complex non-linear model forecasts fail.
+This study proceeds through following stages.
+We present the DCM formulation (Section 2.1).
+The WD-effect is discussed (Section 2.2).
+Complex time series are simulated using the DCM model (Sect.
+2.3).
+The DFT and its ALs are presented (Section 2.4).
+We compare our DCM to the renowned DFT.
+The time series analysis of simulated data sets shows that the ALs of DFT do not constrain DCM (Sections 3.1-3.8).
+Then, we present the identification of the best DCM model for the data (Section 3.9), the DCM significance estimates (Section 3.10), the well-posed computational DCM model solution (Section 3.11) and the DCM forecasts (Section 3.12).
+Finally, we discuss our results (Section 4).
 2 Methods
 The observations are $y_{i}=y(t_{i})\pm\sigma(t_{i})$, where $t_{i}$ are the observing times and $\sigma_{i}$ are the errors (i=1,2,,n)(i=1,2,...,n).
-The time span of data, which is also called the data window, is $\Delta T=t_{n}-t_{1}$.
+The sample window is $\Delta T=t_{n}-t_{1}$.
 The mid point is $t_{\mathrm{mid}}=t_{1}+\Delta T/2$.
@@
 2.1 Discrete Chi-square method (DCM)
-[21] formulated DCM. The model of this method is
+[23] introduced the DCM which is a frequency-domain parametric time series analysis method. The primary objective of this current paper is to show that the DCM outperforms all other similar methods.
+The DCM model is
 $g(t)=g(t,K_{1},K_{2},K_{3})=h(t)+p(t),$ (1)
@@
 In this case, the model becomes linear because the partial derivatives $\partial g/\partial\beta_{i}$ no longer contain any free parameters.
-The solution for the second group of remaining free parameters
+The LS solution for the second group of remaining free parameters
 $\bm{\beta}_{II}$ $=$ $[B_{1,1}C_{1,1},...,B_{K_{1},K_{2}},C_{K_{1},K_{2}},$
@@
 The sum $h(t)$ of signals $h_{1}(t,f_{1})$, $h_{2}(t,f_{2})$ … and $h_{K_{1},f_{K_{1}}}(t)$ does not depend on the order in which these signals are added.
-For example, the two signal $K_{1}=2$ model symmetry is $h(t)=h_{1}(t,f_{1})+h_{2}(t,f_{2})=h_{1}(t,f_{2})+h_{2}(t,f_{1})$.
+For example, the two signal $K_{1}=2$ model symmetry is
+$h(t)$ $=$ $h_{1}(t,f_{1})+h_{2}(t,f_{2})$
+$=$ $h_{1}(t,f_{2})+h_{2}(t,f_{1})$
 Both $(f_{1},f_{2})$ and $(f_{2},f_{1})$ combinations give the same value for the $g(t)$ model.
@@
 $f_{\mathrm{max}}\geq f_{1}>f_{2}>...>f_{K_{1}}\geq f_{\mathrm{min}}.$ (12)
+This frequency space symmetry idea reduces CPU consumption dramatically.
+For example, were this symmetry not exploited, the four signal $K_{1}=4$ search would give $4!=24$ exactly the same solutions from four-dimensional tested grid in frequency space.
+This would cause unnecessary use of CPU in the preliminary testing all possible frequency combinations inside a tesseract (a four-dimensional cube).
+The search for the best frequency combination from $4!=24$ different non-linear iterations would be a pointless exercise (see Equation 22).
 The DCM model residuals
@@
 $\chi^{2}=\sum_{i=1}^{n}{{\epsilon_{i}^{2}}\over{\sigma_{i}^{2}}}.$ (15)
-For every tested $\bm{\beta}_{I}=[f_{1},f_{2},...,f_{K_{1}}]$ frequency combination, the DCM test statistic
-$z$ $=$ $z(f_{1},f_{2},...,f_{K_{1}})=\sqrt{R/n}$ (15)
-$z$ $=$ $z(f_{1},f_{2},...,f_{K_{1}})=\sqrt{\chi^{2}/n}$ (16)
-is computed from a linear model least squares fit. Equation 15 or Equation 16 is applied for unknown or known $\sigma_{i}$ errors, respectively. The value of $z$ is unique.
+For every tested $\bm{\beta}_{I}=[f_{1},f_{2},...,f_{K_{1}}]$ frequency combination, the LS fit gives the DCM test statistic
+$z$ $=z(f_{1},f_{2},...,f_{K_{1}})=\sqrt{R/n},$ unknown $\sigma_{i}$ (16)
+$z$ $=z(f_{1},f_{2},...,f_{K_{1}})=\sqrt{\chi^{2}/n},$ known $\sigma_{i}$ (17)
+computed for a linear model.
+The value of $z$ is unique.
+The errors can be unknown in Equation 16 (AL1).
+For known errors, DCM uses this information in Equation 17 (AL2).
 In the preliminary long search, we test an evenly spaced grid of $n_{\mathrm{L}}$ frequencies between $f_{\mathrm{min}}$ and $f_{\mathrm{max}}$.
@@
 where $n_{f}=n_{L}$ and $n_{f}=n_{S}$, respectively.
-The even or uneven data spacing is irrelevant because the linear least squares fit result for every tested frequency combination does not depend on this spacing.
+The even or uneven data spacing is irrelevant because the LS fit result for every tested frequency combination does not depend on this spacing (AL3).
 The global periodogram minimum
@@
 In the above-mentioned $K_{1}=2$ map, the slice $z_{1}(f_{1})$ would represent the height $z$ at the location $(f_{1},f_{\mathrm{2,best}})$ when moving along the straight constant line $f_{2}=f_{\mathrm{2,best}}$ that crosses the global minimum $z_{\mathrm{min}}$ (Equation 20) at the coordinate point $(f_{\mathrm{1,best}},f_{\mathrm{2,best}})$.
-These one-dimensional periodogram slices (Equation 20) allow us "to see inside” the multi-dimensional structure of the DCM test statistic $z$ (Equations 15 and 16).
+These one-dimensional periodogram slices (Equation 21) allow us "to see inside” the multi-dimensional structure of the DCM test statistic $z$ (Equations 16 and 17).
 This kind of visualisation is important in time series analysis [38].
@@
 For our non-linear DCM model, we determine the $i$:th signal parameters
-  • $P_{i}=1/f_{i}=$ Period
-  • $A_{i}=$ Peak to peak amplitude
-  • $t_{\mathrm{i,min,1}}=$ Deeper primary minimum
-  • $t_{\mathrm{i,min,2}}=$ Secondary minimum (if present)
-  • $t_{\mathrm{i,max,1}}=$ Higher primary maximum
-  • $t_{\mathrm{i,max,2}}=$ Secondary maximum (if present)
+$P_{i}$ $=$ $1/f_{i}=\text{Period}$ (23)
+$A_{i}$ $=$ Peak to peak amplitude (24)
+$t_{\mathrm{i,min,1}}$ $=$ Deeper primary minimum (25)
+$t_{\mathrm{i,min,2}}$ $=$ Secondary minimum (if present) (26)
+$t_{\mathrm{i,max,1}}$ $=$ Higher primary maximum (27)
+$t_{\mathrm{i,max,2}}$ $=$ Secondary maximum (if present) (28)
 and the trend parameters
-  • $M_{k}$ = Polynomial coefficients.
+$M_{k}$ $=$ Polynomial coefficients. (29)
 Of these model parameters, all $P_{i}=1/f_{i}$ and all $M_{k}$ estimates are among the free parameters $\bm{\beta}$.
@@
 We use the computational statistical bootstrap technique [11] to solve these analytical problems.
-This approach gives the DCM model parameter error estimates, as well as the $\sigma_{g(t)}$ model error estimates (Equation 22).
 In our bootstrap, the tested frequencies are the same as in the short search.
@@
 Any $\epsilon_{i}$ value can be chosen to the $\bm{\epsilon}^{*}$ sample as many times as the random selection happens to favour it.
-We create numerous artificial bootstrap data sets
-$\bm{y}^{*}=\bm{g}+\bm{\epsilon}^{*}.$ (23)
-Each $\bm{y}^{*}$ sample gives one estimate for every DCM model parameter value and model $g(t)$ value.
-For each particular model parameter and each particular $g(t)$, the bootstrap error estimate is the standard deviation of all estimates obtained from all $\bm{y}^{*}$ samples.
+We create $J=1,2,...,n_{\mathrm{B}}$ residual random samples $\bm{\epsilon}^{*}_{J}$.
+Every $\bm{\epsilon}^{*}_{J}$ sample gives one artificial bootstrap data set
+$\bm{y}^{*}_{J}=\bm{g}+\bm{\epsilon}^{*}_{J}.$ (31)
+Each $\bm{y}^{*}_{J}$ sample gives one free parameter estimate $\bm{\beta}_{J}$.
+This bootstrap procedure gives $n_{\mathrm{B}}$ free parameter estimates $\bm{\beta}_{J}$.
+The standard deviations for all $n_{\mathrm{B}}$ estimates for $P_{i}=1/f_{i}$ and all $M_{k}$ give the errors of these model parameters.
+We use each $\bm{\beta}_{J}$ to compute the model for a dense grid of time points.
+This gives us $n_{\mathrm{B}}$ numerical $A_{i}$ $t_{\mathrm{i,min,1}}$, $t_{\mathrm{i,min,2}}$, $t_{\mathrm{i,max,1}}$ and $t_{\mathrm{i,max,2}}$ estimates.
+The standard deviation of those $n_{\mathrm{B}}$ estimates is the error of these model parameters.
+The $\bm{\beta}_{J}$ values give $n_{\mathrm{B}}$ estimates for $g(t)$ at any time $t$.
+The standard deviation of these $g(t)$ estimates gives the error $\sigma_{g(t)}$ of Equation 30.
+Note that these $\bm{\beta}_{J}$ values can also be used to compute the errors for $h(t),h_{i}(t),p(t)$ and $p_{k}(t)$.
+Our computational statistical bootstrap approach gives not only the errors estimates for the DCM model parameters (AL4), but also the model error inside $\Delta T$ and the forecast error outside $\Delta T$ (AL5: Equation 30).
 There are, of course, totally wrong DCM models for the data.
@@
 If the frequency $f_{1}$ approaches the frequency $f_{2}$, both $h_{1}(t)$ and $h_{2}(t)$ signals become essentially one and the same signal.
-The least squares fit fails because it makes no sense to model the same signal twice.
+The LS fit fails because it makes no sense to model the same signal twice.
 Dispersing amplitudes instability can occur, if the two signal frequencies are too close to each other.
-The least squares fit finds a model, where two high amplitude signals nearly cancel out each other.
+The LS fit finds a model, where two high amplitude signals nearly cancel out each other.
 The low amplitude signal, the sum of these two high amplitude signals, fits to the data.
@@
 (g) DFT model detrended $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (black continuous line), DFT pure sine $s_{\mathrm{DFT}}(t)$ (red thick continuous line), detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and DFT model residuals (blue dots) offset to -1.5 level (blue dotted line).
-2.2 Simulated DCM model data samples
-We use seven different $g(t)$ models to simulate 21 artificial complex data sets (Sections 3.1-3.7).
-The time interval of all simulated data sets is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$.
+The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [24, TSPA, Equation 1] and the Continuous Period Search [29, CPS, Equation 3].
+TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$.
+TSPA can detect pure sine and double wave signals ($K_{2}=1$ or 2).
+The extension of TSPA, the CPS method, tests three alternatives: $K_{2}=1$ or 2, or no signal at all.
+The DCM model is more sophisticated because it can have any arbitrary $K_{1}$, $K_{2}$ and $K_{3}$ combination.
+The DCM sum $g(t)$ =h(t)+p(t)=h(t)+p(t) of arbitrary periodic and aperiodic functions represents a universal model because the innumerable $K_{1}$, $K_{2}$ and $K_{3}$ combinations allow unlimited complexity.
+Even more complex signal shape $(K_{2})$ combinations could be used.
+For example, these shapes could be $K_{2,i=1}=3$ (3rd harmonic), $K_{2,i=2}=1$ (pure sine) and $K_{2,i=3}=2$ (double wave), where if $K_{2,i}$ denotes the shape of $i$:th signal.
+2.2 WD-effect: Spearhead of DCM
+In this section, we discuss what causes the WD-effect defined in our abstract:
+  • “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.”
+The consequences of this effect are also discussed.
+The $t_{i}$, $y_{i}$ and $\sigma_{i}$ data are inside the rectangle $\Delta T\times\Delta y=(t_{n}-t_{1})\times[{\mathrm{max}}(y_{i}+\sigma_{i})-{\mathrm{min}}(y_{i}-\sigma_{i})]$.
+The LS fit results do not depend on $\Delta T$.
+If the measuring time intervals $\delta t_{i}$ for each observation $y_{i}$ fulfil $\delta t_{i}\ll\Delta T$, the $R$ and $\chi^{2}$ values obtained from all LS (Equation 19) depend only on $y_{i}$ changes in $\Delta Y$ direction, but not on the $t_{i}$ changes in $\Delta T$ direction.
+For fixed $y_{i}$ and $\sigma_{i}$, the residuals $\epsilon_{i}$ determine the $R$ and $\chi^{2}$ values.
+These residuals measure only the $\Delta y$ direction.
+Hence, the periodogram $z$ values (Equation 16 or 17) do not depend on $\Delta T$.
+If $\Delta T$ decreases, the $R$ and $\chi^{2}$ estimates can measure the $g_{i}$ model details inside the $\Delta T\times\Delta y$ rectangle only if $n$ increases and/or $\sigma$ decreases.
+Better data reveal these model details.
+Plenty of concrete examples will confirm that the WD-effect is real (Sections 3.1-3.7).
+Our DCM can detect signals(-s) when the sample window $\Delta T$ is shorter than the period $P$ value(-s).
+The DCM surpasses AL6.
+The data window $\Delta T$ length is irrelevant.
+This means that the DCM can model an infinitesimal time series, as well as forecast its future and past, if the sample size $(n)$ and/or the data accuracy $(\sigma)$ are sufficient.
+That is a revolutionary achievement.
+2.3 Simulated DCM model time series samples
+We use seven different $g(t)$ models to simulate 21 artificial complex time series (Sections 3.1-3.7).
+The sample window of all simulated time series is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$.
 The first and last time point values are then modified to $t_{1}^{\star}=0$ and $t_{n}^{\star}=\Delta T$.
@@
 2.4 Discrete Fourier Transform (DFT)
-We search for signals in the simulated data using the DFT version formulated by [19]. This method searches for the best pure sine model for the data. The notations for the DFT model are
+The DFT is one of the most prevalent frequency-domain parametric time series analysis methods for unevenly spaced time series.
+It searches for the best pure sine model for the data.
+The equivalent DCM model has the orders $K_{1}=1,K_{2}=1$ and $K_{3}=0$.
+Any time series analysis method must be remarkable if it performs better than the distinguished DFT.
+Therefore, we will compare how DCM and DFT perform in the analyses of simulated time series (Sections 3.1-3.7).
+We search for signals in these simulated time series by using the frequently applied DFT version formulated by [21] [^2] This paper had over 2600 citations in December 2025.
+Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [21, Equation 1].
+The notations for the DFT model are
 $g_{\mathrm{DFT}}(t)=s_{\mathrm{DFT}}(t)+p_{\mathrm{DFT}}(t),$ (34)
-where $s_{\mathrm{DFT}}(t)$ is the pure sine signal and $p_{\mathrm{DFT}}(t)$ is the trend.
+where $s_{\mathrm{DFT}}(t)$ is the sum of pure sine signals and $p_{\mathrm{DFT}}(t)$ is the trend.
 The pure sine signals for the $y_{i}$ data and the $\epsilon_{i}$ residuals are denoted with $s_{y,\mathrm{DFT}}(t)$ and $s_{\epsilon,\mathrm{DFT}}(t)$, respectively.
-Our DFT analysis may fail due to four application limitations.
-The data sample is “too short” [22]
+Our DFT analyses of simulated time series may fail due to AL6-AL10.
+The time series is “too short” [25, AL6]
 $P>\Delta T.$ (35)
-The data contain “a trend” [26]
+The time series is non-stationary because it contains “a trend” [33, AL7]
 $p_{\mathrm{DFT}}(t)\neq 0.$ (36)
-The signal frequencies are “too close” [22].
+Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [25, AL9].
 $|f_{1}-f_{2}|<f_{0}={\Delta T}^{-1}.$ (37)
-The signals are not “pure sines” [6].
+The signals are not “pure sines” [5, AL10]
 $K_{2}\neq 1.$ (38)
-The $p_{i}=p_{\mathrm{DFT}}(t_{i})$ trend is first removed from the data $y_{i}=y(t_{i})$ [26].
-Then, the technique described by Ghaderpour et al.
-[15] is used to detect the pure sine signals.
-We apply the DFT to the detrended $y_{i}-p_{i}$ data.
-When searching for the second signal, we apply the DTF to the original model residuals $\epsilon_{i}=y_{i}-g_{\mathrm{y,DFT}}(t_{i})$.
+Many parametric frequency-domain time series analysis methods, like the DFT, can be applied only to stationary data.
+Trends changing the time series mean or variance must be removed before applying these methods [33].
+The DFT can detect only one period at the time.
+Such frequency-domain parametric time series analysis methods are hereafter called “one-dimensional”.
+Our DFT analysis of simulated data proceeds through two stages.
+First, the $p_{i}=p_{\mathrm{DFT}}(t_{i})$ trend is removed from the simulated data $y_{i}=y(t_{i})$.
+Then, the iterative pre-whitening technique described by Ghaderpour et al.
+[17] is applied to search for the pure sine signals.
+We search for the first signal by applying the DFT to the detrended $y_{i}-p_{i}$ simulated data.
+In the second signal search, the DTF is applied to the original model residuals $\epsilon_{i}=(y_{i}-p_{i})-g_{\mathrm{y,DFT}}(t_{i})$.
+At any stage, this combination of detrending and iterative pre-whitening analysis may fail and corrupt the results obtained at the next stages.
+For example, all DFT analyses of simulated time series will fail already at the detrending stage (Sections 3.1-3.7).
+Our DCM has no corrupting separate stages because the trend and the signal(-s) are detected (modelled) simultaneously.
 3 Results
-We apply DCM and DFT to seven different simulated data sets (Sections 3.1 - 3.7).
-As we proceed, these data sets become increasingly complex.
-DCM analysis succeeds for all data sets.
-DFT analysis fails for every data set.
+We apply DCM and DFT to seven different simulated time series (Sections 3.1 - 3.7).
+As we proceed, these time series become increasingly complex.
+DCM analysis succeeds for all time series.
+DFT analysis fails for every time series.
 3.1 Model 1
-Our first simulation model is the one signal model
+Our first time series simulation model is the one signal model
 $g(t)=(A_{1}/2)\cos{\left[{{2\pi(t-t_{\mathrm{1,max,1}})}\over{P_{1}}}\right]}+M_{0}.$ (39)
@@
 [^3] We publish all our data files and all DCM control files in https://zenodo.org/uploads/17018676 The detected $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$ and $M_{0}$ values are correct and accurate even for the combination $n=50$ and ${\mathrm{SN}}=50$.
-These values become more accurate when $n$ and $\mathrm{SN}$} increase.
-A graphical presentation of DCM analysis results is shown for Model 1 simulated data sample, where $n=50$ and ${\mathrm{SN}}=50$ (Figures 1a-d).
+Regardless of $\Delta T<P_{1}$, these model parameter values become more accurate and converge to the correct simulated values when $n$ and $\mathrm{SN}$} increase.
+This confirms the WD-effect.
+A graphical presentation of DCM analysis results is shown for Model 1 simulated time series, where $n=50$ and ${\mathrm{SN}}=50$ (Figures 1a-d).
 The DCM long search $z_{1}(f_{1})$ periodogram minimum is at $P_{1}=1.843$ (Figure 1a: diamond).
@@
 The DFT residuals (blue dots) offset to the level of -1.5 (blue dotted line) display obvious trends, especially at the end of analysed sample.
-For the simulated data of Model 1, the DCM analysis succeeds, but the DFT analysis fails.
+For the simulated time series of Model 1, the DCM analysis succeeds, but the DFT analysis fails.
 Table 2: Model 2. DCM analysis between $P_{\mathrm{min}}=0.63$ and $P_{\mathrm{max}}=4.70$. Notations as in Table 1.
@@
 3.2 Model 2
-Our next one signal data simulation model is
+Our next one signal time series simulation model is
 $g(t)$ $=$ $(A_{1}/2)\cos{\left[{{2\pi(t-t_{\mathrm{1,max,1}})}\over{P_{1}}}\right]}$
@@
 As a DCM model, the orders of Model 2 are $K_{1}=1$, $K_{2}=1$ and $K_{3}=2$.
-The simulated data sample is “too short” because the period $P_{1}$ is $1.9\times\Delta T$ (Equation 27).
+The simulated time series is “too short” because the period $P_{1}$ is $1.9\times\Delta T$ (Equation 35).
 The parabolic trend $p(t)$ is unknown (Equation 36).
@@
 These results are not very accurate for the $n=100$ and ${\mathrm{SN}}=100$ combination, but they definitely improve for larger $n$ and $\mathrm{SN}$} values.
+The WD-effect ensures that the detected values converge to the correct simulated model parameter values.
+The short sample window, $\Delta T<P_{1}$, does not mislead the DCM analysis.
 For Model 2, the DCM analysis results are illustrated for the n=10000n=10~000 and ${\mathrm{SN}}=100$ combination (Figures 2a-d).
@@
 The trends of these residuals confirm that the DFT analysis fails.
-Only DCM (not DFT) succeeds in the analysis of Model 2 simulated data.
+Only DCM (not DFT) succeeds in the analysis of Model 2 simulated time series.
 3.3 Model 3
-Our third data simulation model is the two signal model
+Our third time series simulation model is the two signal model
 $g(t)$ $=$ $(A_{1}/2)\cos{\left[{{2\pi(t-t_{\mathrm{1,max,1}})}\over{P_{1}}}\right]}$
@@
 The DCM orders of Model 3 are $K_{1}=2$, $K_{2}=1$ and $K_{3}=0$.
-The simulated data sample is not “too short” (Equation 27) because $P_{1}<P_{2}<\Delta T$.
+The simulated time series is not “too short” (Equation 35) because $P_{1}<P_{2}<\Delta T$.
 The $p(t)$ trend mean level $M_{0}$ is unknown (Equation 36).
@@
 The results are surprisingly accurate even for the small $n=50$ sample having a low ${\mathrm{SN}}=10$.
+Due to the WD-effect, the detected model parameter values converge to the correct simulated values when $n$ and $\mathrm{SN}$} increase.
 The DCM results for Model 3 are illustrated in Figures 1a-d ($n=50$ and ${\mathrm{SN}}=50$ combination).
@@
 The DFT residuals (blue dots) offset to the level of -3.0 (blue dotted line) are not stable.
-The DCM analysis of Model 3 simulated data succeeds, but the DFT analysis does not.
+The DCM analysis of Model 3 simulated time series succeeds, but the DFT analysis does not.
 3.4 Model 4
-The next data simulation model is
+The next time series simulation model is
 $g(t)$ $=$ $(A_{1}/2)\cos{\left[{{2\pi(t-t_{\mathrm{1,max,1}})}\over{P_{1}}}\right]}$
@@
 The DCM detects the correct $P_{1}$, $P_{2}$, $A_{1}$, $A_{2}$, $t_{\mathrm{max,1}}$, $t_{\mathrm{max,2}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values even for the lowest $n=50$ and ${\mathrm{SN}}=10$ combination.
-Our Figures 4a-d illustrate the DCM analysis results for one sample of simulated data (Model 4: $n=50$ and ${\mathrm{SN}}=50$).
+As the simulated data $n$ and $\mathrm{SN}$} values increase, the WD-effect ensures that all detected values converge to the correct simulated model parameter values.
+Our Figures 4a-d illustrate the DCM analysis results for one sample of simulated time series (Model 4: $n=50$ and ${\mathrm{SN}}=50$).
 The DCM long search best periods are at $P_{1}=0.159$ and $P_{2}=0.168$ (Figure 4a: diamonds).
@@
 The blue dots representing the DFT model residuals are offset to the level of -3.0 and show clear trends.
-Our Model 4 simulated data analysis succeeds for the DCM and fails for the DFT.
+Our Model 4 simulated time series analysis succeeds for the DCM and fails for the DFT.
 Table 5: Model 5. DCM analysis between $P_{\mathrm{min}}=0.47$ and $P_{\mathrm{max}}=4.20$. Notations as in Table 1.
@@
 3.5 Model 5
-The mathematical equation
+The mathematical time series equation
 $g(t)$ $=$ $(A_{1}/2)\cos{\left[{{2\pi(t-t_{\mathrm{1,max,1}})}\over{P_{1}}}\right]}$
@@
 $+$ $M_{0}+M_{1}T+M_{2}T^{2}$
-for our fifth model for simulated data is the same for Model 4 (Equation 3.4).
+for our fifth model is the same as for Model 4 (Equation 3.4).
 However, this new Model 5 differs from the earlier Model 4 because we use totally different $P_{1}$, $P_{2}$, $A_{1}$$A_{2}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{2,max,1}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values (Table 5, Column 1).
-The simulated data sample is “too short” for both $P_{1}$ and $P_{2}$ periods (Equation 27).
+The simulated time series is “too short” for both $P_{1}$ and $P_{2}$ periods (Equation 35).
 There is “a trend” $p(t)$ (Equation 36).
@@
 This shows that the DCM can fail, just like any other time series analysis method, if the quality of data is too low.
-The DCM results for higher $n$ and $\mathrm{SN}$} combinations are correct (Table 5, Columns 3-4).
-We show the DCM analysis results for one sample of Model 5 simulated data (Figures 5a-d: n=10000n=10~000 and ${\mathrm{SN}}=10~000$).
+Due to the WD-effect, the DCM results for higher $n$ and $\mathrm{SN}$} combinations are correct (Table 5, Columns 3-4).
+We show the DCM analysis results for one sample of Model 5 simulated time series (Figures 5a-d: n=10000n=10~000 and ${\mathrm{SN}}=10~000$).
 The long and short DCM searches give $P_{1}=1.407$ and $P_{2}=1.917$ (Figure 5a: diamonds), and $P_{1}=1.393$ and $P_{2}=2.024$ (Figure 5b: Diamonds).
@@
 These residuals show strong trends.
-The DCM analysis succeeds for Model 5 simulated data, but the DFT analysis fails.
+The DCM analysis succeeds for Model 5 simulated time series, but the DFT analysis fails.
 3.6 Model 6
-Our sixth data simulation model is
+Our sixth time series simulation model is
 $g(t)$ $=$ $c_{1}\cos{\left[{{2\pi t}\over{P_{1}}}\right]}$
@@
 The coefficients $c_{1}=0.3655$, $c_{2}=0.7310$ and $c_{3}=0.3000$ determine the $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,min,2}}$, $t_{\mathrm{1,max,1}}$ and $t_{\mathrm{1,max,2}}$ values given in Table 6 (Column 1).
-This simulated data sample is not “too short” because $P_{1}<\Delta T$ (Equation 27).
+This simulated time series is not “too short” because $P_{1}<\Delta T$ (Equation 35).
 There is no trend because $p(t)=M_{0}=0$ (Equation 36).
-These simulated data contain only one signal (Equation 29).
+This simulated time series contains only one signal (Equation 37).
 However, this simulation Model 6 is not a $K_{2}=1\texteuro$ pure sine model (Equation 38).
@@
 The DCM detects the correct $P_{1}$, $A_{1}$, $t_{\mathrm{1,min,1}}$ $t_{\mathrm{1,min,2}}$ $t_{\mathrm{1,max,1}}$ $t_{\mathrm{1,max,2}}$ and $M_{0}$ values even for the lowest $n=50$ and ${\mathrm{SN}}=10$ combination.
-We demonstrate the DCM analysis results for simulated data having $n=50$ and ${\mathrm{SN}}=50$ (Figures 6a-d).
+For increasing $n$ and $\mathrm{SN}$}, the results for the model parameters converge to correct values (WD-effect).
+We demonstrate the DCM analysis results for simulated time series having $n=50$ and ${\mathrm{SN}}=50$ (Figures 6a-d).
 The long and short searches give $P_{1}=0.159$ (Figure 6a: diamond) and $P_{1}=0.160$ (Figure 6b: diamond).
@@
 We conclude that the DFT “detects” the $P_{1}/2$ and $P_{1}$ periods, where $P_{1}$ is the correct simulated period value.
-However, the DFT two signal model is not the correct model for these Model 6 simulated data, which contain only one signal.
+However, the DFT two signal model is not the correct model for these Model 6 simulated time series, which contains only one signal.
 If the correct period is $P$ and the correct model is a double wave $(K_{2}=2)$, the DCM pure sine model $(K_{1}=1)$ analysis can also give the values $P/2$ and $P$.
-The DCM analysis succeeds for Model 6 simulated data. The DFT analysis fails.
+The DCM analysis succeeds for Model 6 simulated time series. The DFT analysis fails.
 3.7 Model 7
-In this section, we analyse our most complex data set. The seventh simulation model is
+In this section, we analyse our most complex time series. The seventh simulation model is
 $g(t)$ $=$ $c_{1}\cos{\left[{{2\pi t}\over{P_{1}}}\right]}+c_{2}\cos{\left[{{4\pi(t-c_{3})}\over{P_{1}}}\right]}$
@@
 The coefficients $c_{1}=0.3687$, $c_{2}=0.7374$, $c_{3}=0.4000$, $c_{4}=0.3708$, $c_{5}=0.7416$ and $c_{6}=0.6000$ determine the $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,min,2}}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{1,max,2}}$, $A_{2}$, $t_{\mathrm{2,min,1}}$, $t_{\mathrm{2,min,2}}$, $t_{\mathrm{2,max,1}}$ and $t_{\mathrm{2,max,2}}$ values given in Table 7 (Column 1).
-This simulated data sample is “too short” because $\Delta T<P_{1}<P_{2}$ (Equation 27).
+This simulated time series is “too short” because $\Delta T<P_{1}<P_{2}$ (Equation 35).
 The parabolic $p(t)$ represents “a trend” (Equation 36).
@@
 These signals are double waves having two unequal minima and maxima.
-All four main reasons that can cause the failure of DFT analysis are present (Equations 27-30).
-Therefore, this simulated data set is the most complex one of all analysed seven data sets.
-Our DCM and DFT time series analysis of Model 7 simulated data is performed between $P_{\mathrm{min}}=P_{1}/3=0.4$ and $P_{\mathrm{max}}=3P_{1}=3.6$.
+All four main reasons that can cause the failure of DFT analysis are present (Equations 35-38).
+Therefore, this simulated time series is the most complex one of all analysed seven time series.
+Our DCM and DFT time series analysis of Model 7 simulated time series is performed between $P_{\mathrm{min}}=P_{1}/3=0.4$ and $P_{\mathrm{max}}=3P_{1}=3.6$.
 The DCM orders of Model 7 are $K_{1}=2$, $K_{2}=2$ and $K_{3}=2$.
@@
 We give the DCM analysis results for different $n$ and $\mathrm{SN}$} combinations in Table 7 (Columns 2-4).
-The DCM analysis results are displayed for one sample of Model 7 simulated data (Figures 7a-d: $n=100$ and ${\mathrm{SN}}=5~000~000$).
-Since there are $\eta=13$ free model parameters, this $n=100$ sample is quite small for time series analysis.
+The DCM analysis results are displayed for one sample of Model 7 simulated time series (Figures 7a-d: $n=100$ and ${\mathrm{SN}}=5~000~000$).
+Since there are $\eta=13$ free model parameters, this $n=100$ time series is quite small for time series analysis.
 The long and short DCM searches give $P_{1}=1.200$ and $P_{2}=1.385$ (Figure 7a: diamonds), and $P_{1}=1.199$ and $P_{2}=1.405$ (Figure 7b: Diamonds).
@@
 These DCM model residuals are small and their level is stable.
-These results confirm that the DCM time series analysis method can detect complex non-linear models $(\eta=13)$ from very small data sets $(n=100)$, if these data are extremely accurate $({\mathrm{SN}}=5~000~000).$
+These results confirm that the DCM time series analysis method can detect complex non-linear models $(\eta=13)$ from very small time series $(n=100)$, if the data are extremely accurate $({\mathrm{SN}}=5~000~000).$ These results demonstrate the power of WD-effect because both periods $P_{1}$ and $P_{2}$ are shorter than the sample window $\Delta T$, the trend is a parabola and the signals are not pure sines.
 The DFT time series analysis gives the wrong periods for the original data (Figure 7e: diamond at $P_{1}=0.542$ and for the residuals (Figure 7f: diamond at $P_{2}=1.059$.
@@
 The DFT model $g_{\mathrm{DFT}}(t)$ (black continuous line) deviates from the data $y_{i}$ (black dots).
-This deviation is largest at the beginning and the end of the simulated data sample.
+This deviation is largest at the beginning and the end of the simulated time series.
 In our Figure 7h, the black dots denote the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$, the continuous black line denotes the detrended DFT model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ and the continuous thick red line denotes pure sine signal $s_{\mathrm{y,DFT}}(t)$ detected from the original data.
@@
 The correct simulated peak to peak amplitude for this second signal is much larger, $A_{2}=2.$ The DFT model residuals (blue dots) are offset to -1.5 level and show strong trends.
-The DCM time series analysis succeeds for Model 7 simulated data. The DFT time series analysis fails, as predicted by Equations 27-30.
+The DCM analysis succeeds for Model 7 simulated time series. The DFT analysis fails, as predicted by Equations 35-38.
+3.8 Summary of simulations
+For all seven simulated time series, all DCM analyses succeed and all DFT analyses fail (Sections 3.1 - 3.7).
+[…]
+The analyses of seven simulated time series expose the DFT weaknesses.
+Every analysis fails!
+The AL6-AL10 are not the only causes for these failures.
+In general, the solutions for one-dimensional time series analysis models are ill-posed, if the data contain an unknown trend and/or more than one signal.
+For example, the detrending of simulated data misleads DFT even when the correct $K_{3}$ value is known (Sections 3.1-3.7).
+This correct $K_{3}$ value can be unknown for real data.
+The combination of detrending and iterative pre-whitening can fail at any stage.
+Even if the detrending succeeds, the leakage can cause the detection of wrong first signal frequency.
+This corrupts the model residuals and the whole analysis.
+For any one-dimensional time series analysis method, like the DFT, the search for many signals superimposed on an unknown trend can fail when the combination of detrending and pre-whitening is applied.
+Our DCM outperforms the esteemed DTF.
+The DCM detects the correct signal(s) and trend, but DFT does not.
+For the sunspot data, the DCM detects signals superimposed on a constant trend (Jetsu 2025).
+The simulations presented here support the conclusion that those solar signals and the trend are correct.
+This would explain why the DCM is the first time series analysis method that detects Jupiter’s exact $11.^{\mathrm{y}}86$ period in the sunspot record.
+The one-dimensional time series analysis methods, like DFF, have failed to find exactly correct periods.
 Table 8: Fisher tests.
-DCM analysis of Model 1 simulated data ($n=100,{\mathrm{SN}}=100$, data file Model1n100SN100).
+DCM analysis of Model 1 simulated time series ($n=100,{\mathrm{SN}}=100$, data file Model1n100SN100).
 (1) Simple model: Model number $\mathcal{M}$, model $g_{K_{1},K_{2},K_{3}}$, number of free parameters $\eta$, chi-square $\chi^{2}$ and control file name.
-(2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.22footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
+(2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.33footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
 […]
 3.9 Best model
-The $K_{1}$, $K_{2}$ and $K_{3}$ orders of the best model are not necessarily known when some time series analysis method is applied to the real data.
-We know a priori the best DCM and DFT model orders for the simulated data of Models 1-7 (Sections 3.1-3.7).
+The $K_{1}$, $K_{2}$ and $K_{3}$ orders of the best model are not necessarily known when some time series analysis method is applied to the real time series.
+We know a priori the best DCM and DFT model orders for the simulated time series of Models 1-7 (Sections 3.1-3.7).
 It could be argued that our DCM analysis succeeds only for this reason.
@@
 If the number of complex model free parameters $\eta_{2}$ increases, the $R_{2}$ or $\chi^{2}_{2}$ values decrease.
-This increases the $F_{R}$ or $F_{\chi}$ values because the terms $(R_{1}/R_{2}-1)$ or $(\chi^{2}_{1}/\chi^{2}_{2}-1)$ increase (Equations 38 and 39: first terms).
-At the same time, the $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty term decreases (Equations 38 and 39: second terms), and this decreases the $F_{R}$ or $F_{\chi}$ values.
+This increases the $F_{R}$ or $F_{\chi}$ values because the terms $(R_{1}/R_{2}-1)$ or $(\chi^{2}_{1}/\chi^{2}_{2}-1)$ increase (Equations 46 and 47: first terms).
+At the same time, the $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty term decreases (Equations 46 and 47: second terms), and this decreases the $F_{R}$ or $F_{\chi}$ values.
 This penalty term prevents the use of too high $\eta_{2}$ values (too complex models).
@@
 […]
-Table 9: Fisher tests for Model 1 $n=100,{\mathrm{SN}}=100$ simulated data (electronic data file Model1n100SN100).
-(1) Polynomial model.
+In the above example, the Fisher-test finds the correct number of signals $(K_{1})$ and the correct trend $(K_{3})$ for the pure sine signal alternative $(K_{2}=1)$.
+We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large.
+One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu (2025, Table S7).
+We conclude that the best model for the real time series can be found by applying the Fisher-test to any arbitrary number of different nested DCM or DFT models.
+3.10 Significance estimates
+Jetsu [23] or Jetsu (2025) gave no signal significance estimate for the first detected period.
+Here, we use the Fisher-test for this purpose.
+The one signal model is the complex model.
+The logical simple model alternative is $g(t)=m\equiv$ white noise having standard deviation $s$.
+However, white noise is not the only possible alternative simple model.
+The other possible nested no signal DCM polynomial models are
+$g(t)=p(t,K_{3})$ (49)
+where $h(t)=0$, $K_{3}=0,1,2..$.
+The $K_{3}=0$ polynomial $g(t)=M_{0}$ represents white noise.
+The Fisher-test gives the critical level $Q_{\mathrm{F}}$ for rejecting the $H_{0}$ hypothesis when this hypothesis is in fact true (Equation 48).
+Therefore, this $Q_{\mathrm{F}}$ value represents the probability of false signal detection.
+In cases $Q_{\mathrm{F}}<\!10^{-16}$, the signal detection is absolutely certain.
+We use the Model 1 simulated time series combination $n=100$ and ${\mathrm{SN}}=100$ to demonstrate the significance estimation for the first detected period.
+The correct model for the first detected signal is $g_{1,1,0}(t)$.
+The Fisher-test is used to compare this correct $g_{1,1,0}(t)$ model to different polynomial models $g(t)=p(t,K_{3})$ having $K_{3}=0,1,..,7$ (Table 9: Equation 49).
+The $\mathcal{M}$=1, 2 and 3 models have less free parameters than the correct $g_{1,1,0}(t)$ model, but the $\chi^{2}$ values of these three polynomial models are so large that the Fisher-test is merely a formality.
+The $\mathcal{M}$=4 model has the same number of free parameters as the correct $g_{1,1,0}(t)$ model, but the comparison of $\chi^{2}$ values reveals that this third order $p(t)$ polynomial is not the correct model for the time series.
+The next model $\mathcal{M}$=5 has more free parameters than the correct $g_{1,1,0}(t)$ model.
+This fourth order $p(t)$ polynomial $\mathcal{M}$=5 model must be rejected because it has a larger $\chi^{2}$ value than the correct $g_{1,1,0}(t)$ model.
+The critical levels $Q_{\mathrm{F}}\gg\gamma=0.001$ for the remaining $\mathcal{M}$=6, 7 and 8 polynomial models are so large that these models must also be rejected.
+The results in Table 9 confirm that the one signal model is better than any polynomial model in Equation 49.
+Hence, the analysed Model 1 simulated time series must contain at least the $h_{1}(t)$ pure sine signal.
+For the constant, linear or parabolic $p(t)$ trend alternatives, the significance for this $h_{1}(t)$ signal is $Q_{\mathrm{F}}<\!10^{-16}$.
+After the detection of the first strongest signal, the Fisher-test critical level $Q_{\mathrm{F}}$ values increase for the next detected weaker signals.
+In other words, the signal significances decrease.
+Typical examples can be found in Jetsu 2025 (Tables S5-S16).
+No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected.
+For the Model 1 simulated $n=100$ and ${\mathrm{SN}}=100$ combination, Table 8 shows that this time series contains only one signal.
+For this Model 1 time series, there is no need to discuss the DFT significance estimates [21, their Equation 22] because this method fails to detect the correct period.
+We conclude that the Fisher-test identifies the correct DCM model (Section 3.9), as well as gives the signal significance $Q_{\mathrm{F}}$ estimates (Section 3.10).
+Table 9: Fisher tests for Model 1 $n=100,{\mathrm{SN}}=100$ simulated time series (electronic data file Model1n100SN100).
+(1) Polynomial model (Equation 49).
 (2) DCM model $g_{1,1,0}$.
@@
 Otherwise as in Table 8.
-33footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.
+44footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.
 (1) (2)
@@
 […]
-In the above example, the Fisher-test finds the correct number of signals $(K_{1})$ and the correct trend $(K_{3})$ for the pure sine signal alternative $(K_{2}=1)$.
-We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large.
-One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu (2025, Table S7).
-We conclude that the best model for the real data can be found by applying the Fisher-test to any arbitrary number of different nested DCM or DFT models.
-3.9 Significance estimates
-The Fisher-test gives the critical level $Q_{\mathrm{F}}$ for rejecting the $H_{0}$ hypothesis (Equation 40).
-After the detection of the first strongest signal, the values $Q_{\mathrm{F}}$ of these critical levels increase for the next detected weaker signals.
-Typical examples can be found in Jetsu 2025 (Tables S5-S16).
-No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected.
-The Fisher test critical level $Q_{\mathrm{F}}$ is the significance estimate for each signal detected after the first signal.
-However, this Fisher-test process gives no significance estimate for the first detected signal.
-We estimate the critical level $Q_{\mathrm{F}}$ for the first detected signal $h_{1}(t)$ by testing the alternative that there are no signals in the data. The “no signal hypothesis” is
-$h(t)=0\Rightarrow h_{1}(t)=0\Rightarrow g(t)=p(t).$ (41)
-In Model 1 simulated data combination $n=100$ and ${\mathrm{SN}}=100$, the correct model is $g_{1,1,0}(t)$.
-We use the Fisher-test to compare this correct $g_{1,1,0}(t)$ model to different polynomial models $g(t)=p(t)$ having $K_{3}=0,1,..,7$ (Table 9).
-The $\mathcal{M}$=1, 2 and 3 models have less free parameters than the correct $g_{1,1,0}(t)$ model, but the $\chi^{2}$ values of these three polynomial models are so large that the Fisher-test is merely a formality.
-The $\mathcal{M}$=4 model has the same number of free parameters as the correct $g_{1,1,0}(t)$ model, but the comparison of $\chi^{2}$ values reveals that this third order $p(t)$ polynomial is not the correct model for the data.
-The next model $\mathcal{M}$=5 has more free parameters than the correct $g_{1,1,0}(t)$ model.
-This fourth order $p(t)$ polynomial $\mathcal{M}$=5 model must be rejected because it has a larger $\chi^{2}$ value than the correct $g_{1,1,0}(t)$ model.
-The critical levels $Q_{\mathrm{F}}\gg\gamma=0.001$ for the remaining $\mathcal{M}$=6, 7 and 8 polynomial models are so large that these models must also be rejected.
-The results in Table 9 confirm the “no signal hypothesis” (Equation 41) rejection.
-The analysed Model 1 simulated data must contain at least the $h_{1}(t)$ pure sine signal.
-For the constant, linear or parabolic $p(t)$ trend alternatives, the critical level for this $h_{1}(t)$ signal is $Q_{\mathrm{F}}<\!10^{-16}$.
-There is no need to discuss the DFT significance estimates [19, their Equation 22] because this method fails to detect the correct Model 1 period.
-Table 10: DCM time series analysis between $P_{\mathrm{min}}=0.53$ and $P_{\mathrm{max}}=4.80$ for predictive data (Model = 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data: first forty observations).
+Table 10: DCM time series analysis between $P_{\mathrm{min}}=0.53$ and $P_{\mathrm{max}}=4.80$ for forecasting data (Model = 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data: first forty observations).
 (1) Model.
-(2) Predictive data test statistic $z$ (Equation 16).
-(3) Predicted data test statistic $z_{\mathrm{pred}}$ (Equation 3.10).
+(2) Forecasting data test statistic $z$ (Equation 17).
+(3) Forecasted data test statistic $z_{\mathrm{Fore}}$ (Equation 3.12).
 (4) Data file.
 (5) Control file.
-(6) Figure where prediction is shown.
+(6) Figure where forecast is shown.
 (1) (2) (3) (4) (5) (6)
-Model $z$ $z_{\mathrm{pred}}$ Data file Control file Figure
+Model $z$ $z_{\mathrm{Fore}}$ Data file Control file Figure
 $g_{1,1,0}$ 2.27 8.61 Model3n40SN10.dat dcmModel3K110.dat 8a
@@
 […]
-Figure 8: DCM model prediction (Model 3 simulated data for combination $n=50$ and ${\mathrm{SN}}=10$: first forty observations are predictive data).
+Figure 8: DCM model forecast (Model 3 simulated time series for combination $n=50$ and ${\mathrm{SN}}=10$: first forty observations are forecasting data).
 (a) One signal model $g_{1,1,0}$ results.
-Predictive data are $t_{i}$, $y_{i}$ and $\sigma_{i}$ ($n=40$ black dots) and predicted data are ten last observations $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$ ($n=10$ open black dots).
-Continuous black line of predictive data model $g(t,\bm{\beta}_{\mathrm{pred}})$ ends to vertical dotted blue line, where dotted black line of predicted model $g^{\prime}(t,\bm{\beta}_{\mathrm{pred}})$ begins.
+Forecasting data are $t_{i}$, $y_{i}$ and $\sigma_{i}$ ($n=40$ black dots) and forecasted data are ten last observations $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$ ($n=10$ open black dots).
+Continuous black line of forecasting data model $g(t,\bm{\beta}_{\mathrm{Fore}})$ ends to vertical dotted blue line, where dotted black line of forecasted model $g^{\prime}(t,\bm{\beta}_{\mathrm{Fore}})$ begins.
 Dotted red line denotes $\pm 3\sigma_{g(t)}$ errors of both models (Equation 30).
-Residuals of predictive data model $\epsilon_{i}=y_{i}-g_{i}$ (blue dots), residuals of predicted data model $\epsilon_{i}=^{\prime}y_{i}^{\prime}-g_{i}^{\prime}$ (open blue dots) and $\pm 3\sigma$ errors of both models (red dotted line) are offset to level -1 (blue dotted line).
+Residuals of forecasting data model $\epsilon_{i}=y_{i}-g_{i}$ (blue dots), residuals of forecasted data model $\epsilon_{i}=^{\prime}y_{i}^{\prime}-g_{i}^{\prime}$ (open blue dots) and $\pm 3\sigma$ errors of both models (red dotted line) are offset to level -1 (blue dotted line).
 (b) Two signal model $g_{2,1,0}$ results.
@@
 Otherwise as in “a”.
-3.10 Prediction
-There are numerous techniques for predicting (forecasting) a time series [e.g., 18, 20, 23].
-The DCM model $g(t_{i},\bm{\beta})$ can be used to predict.
-We divide all data into the predictive data and the predicted data.
-The time points, the observations and the errors of these samples are
-  • $n$ predictive data values $t_{i}$, $y_{i}$ and $\sigma_{i}$
-  • $n^{\prime}$ predicted data values $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$
-The DCM gives the best predictive data model
-$g_{i}=g(t_{i},\bm{\beta}_{\mathrm{pred}}),$ (42)
-where $\bm{\beta}_{\mathrm{pred}}$ are the free parameter values.
-The $t_{\mathrm{mid,pred}}\!=\!(t_{n}\!+\!t_{1})/2$ and $\Delta T_{\mathrm{pred}}\!=\!t_{n}\!-\!t_{1}$ values are computed from the predictive data time points $t_{i}$.
-The predicted data model values are
-$g^{\prime}_{i}$ $=$ $g(t^{\prime}_{i},\bm{\beta}_{\mathrm{pred}}),$ (43)
-where $t_{\mathrm{mid}}=t_{\mathrm{mid,pred}}$ and $\Delta T=\Delta T_{\mathrm{pred}}$.
-We do not compute “new” $t_{\mathrm{mid}}$ and $\Delta T$ values from the predicted data time points $t^{\prime}_{i}$ because the correct $g^{\prime}_{i}$ values are obtained only from the $\bm{\beta}_{\mathrm{pred}}$, $t_{\mathrm{mid,pred}}$ and $\Delta T_{\mathrm{pred}}$ combination of Equation 42.
-The $n^{\prime}$ predicted data model residuals
-$\epsilon^{\prime}_{i}=y^{\prime}_{i}-g^{\prime}_{i}$
-give the predicted data test statistic
-$z_{\mathrm{pred}}$ $=$ z test statistic for predicted
-data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ (Equation 15 or 16).
-This parameter $z_{\mathrm{pred}}$ measures how well the prediction (Equation 43) obtained from the predictive data works for the predicted data.
-If the DCM detects a new signal from the predictive data, there are two alternatives:
-  • If this new signal is real, the $z_{\mathrm{pred}}$ value of predicted data decreases.
-  • If this new signal is unreal, the $z_{\mathrm{pred}}$ value of predicted data increases.
-This “Prediction-test” technique revealed at least five real signals in the sunspot record (Jetsu 2025).
-We compute the $z_{\mathrm{pred}}$ parameter value from the known predicted data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ values (Equation 3.10) .
-Predictions are possible even if all $t^{\prime}_{i}$, $y^{\prime}_{i}$ or $\sigma^{\prime}_{i}$ predicted data values are unknown.
-In this case, the $t_{i}$, $y_{i}$ and $\sigma_{i}$ values of all data can be used as predictive data.
-The best DCM model $g(t_{i},\bm{\beta}_{\mathrm{pred}})$ for all data determines the correct $\bm{\beta}_{\mathrm{pred}}$, $t_{\mathrm{mid,pred}}=(t_{n}+t_{1})/2$ and $\Delta T_{\mathrm{pred}}=t_{n}-t_{1}$ combination.
-The $n^{\prime}$ predicted data time points $t_{i}^{\prime}$ can be created for any arbitrary chosen time interval $\Delta T^{\prime}=t_{n}^{\prime}-t_{i}^{\prime}$.
-The $n^{\prime}$ predicted model $g_{i}^{\prime}=g^{\prime}_{i}(t_{i}^{\prime},\bm{\beta}_{\mathrm{pred}})$ values are obtained from Equation 43.
-These $g^{\prime}_{i}$ values can be used, for example, to compute the predicted mean level
-$m_{\mathrm{pred}}={{1}\over{n^{\prime}}}\sum_{i=1}^{n^{\prime}}g^{\prime}_{i}$ (45)
-during the chosen time interval $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$.
-Jetsu (2025) used this $m_{\mathrm{pred}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720.
-Since the known mean level of all sunspot data was $m$, Jetsu (2025) used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
-  • If DCM detects a real new signal in all data, the $m_{\mathrm{pred}}$ value decreases.
-  • If DCM detects many real signals in all data, the $m_{\mathrm{pred}}$ value falls below $m$.
-  • If DCM detects an unreal new signal in all data, the $m_{\mathrm{pred}}$ value increases.
-We use Model 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data (Table 3, column 2) to illustrate the DCM prediction technique.
-The black dots in Figures 8a-c are the first $n=40$ predictive data values $t_{i}$, $y_{i}$ and $\sigma_{i}$.
-The open black dots denote the $n^{\prime}=10$ predicted data values $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$.
-The continuous black line is the predictive model $g(t)$ and the dotted black line is the predicted model $g^{\prime}(t)$.
-The red dotted line shows the $\pm 3\sigma$ errors of both models.
-The blue dots are the predictive data $\epsilon_{i}=y_{i}-g_{i}$ residuals and the open blue dots are the predicted data $\epsilon_{i}^{\prime}=y_{i}^{\prime}-g_{i}^{\prime}$ residuals.
-Model 3 is the sum of two $P_{1}=0.16$ and $P_{2}=0.17$ pure sine signals $(K_{1}=2,K_{2}=1)$ superimposed on the constant mean level $M_{0}=1$ $(K_{3}=0)$.
-We compute the $z$ and $z_{\mathrm{pred}}$ values for the $g_{1,1,0}$, $g_{2,1,0}$ and $g_{3,1,0}$ models (Table 10).
-These three models have the same $K_{2}=1$ and $K_{3}=0$ orders as Model 3, but their signal numbers $K_{1}=1,2$ or 3 are different, the $g_{2,1,0}$ model being the correct simulation Model 3.
-The correct $g_{2,1,0}$ model gives the smallest $z_{\mathrm{pred}}=0.88$ value (Table 10).
-Therefore, it is a better predictive model than the $g_{1,1,0}$ and $g_{3,1,0}$ models.
-The one signal $g_{1,1,0}$ model prediction fails because the blue open circles denoting the predicted data residuals $\epsilon_{i}^{\prime}=y_{i}^{\prime}-g_{i}^{\prime}$ show large deviations from to the blue dotted line offset level of $\epsilon_{i}^{\prime}=-1$ (Figure 8a).
-The two signal $g_{2,1,0}$ model prediction succeeds because all open blue dots denoting the predicted data residuals stay close to the blue dotted line offset level $\epsilon_{i}^{\prime}=-1$, as well as inside the red dotted $\pm 3\sigma$ model error limits (Figure 8b).
-The three signal model $g_{3,1,0}$ periods are $P_{1}=0.0643\pm 0.0018$, $P_{2}=0.1592\pm 0.0012$ and $P_{3}=0.1716\pm 0.0012$.
-The amplitudes of these pure sine signals are $A_{1}=0.092\pm 0.025$, $A_{2}=1.879\pm 0.16$ and $A_{3}=1.82\pm 0.15$.
-The periods and the amplitudes of the two strongest $P_{2}$ and $P_{3}$ signals are correct because they are the same as in Model 3 simulation (Table 3, column 1).
-Therefore, the $g_{3,1,0}$ prediction appears nearly as good as the $g_{2,1,0}$ prediction because the residuals $\epsilon_{i}^{\prime}$ are close to the offset level of $\epsilon_{i}^{\prime}=-1$ (blue open dots), and these residuals also stay inside the red dotted $\pm 3\sigma$ model error limits (Figure 8c).
-The amplitude $A_{1}=0.092$ of the third $P_{1}=0.0643$ signal is very low.
-Due this weak “unreal” $P_{1}$ signal, the $g_{3,1,0}$ model has a larger $z_{\mathrm{pred}}=0.91$ value than the correct $g_{2,1,0}$ model (Table 10).
-The $n=40$ predictive data parameters $\eta_{1}=7$, $\eta_{2}=10$, $\chi_{1}=22.946$ and $\chi_{2}=19.790$ for the simple $g_{2,1,0}$ model and the complex $g_{3,1,0}$ model give Fisher-test values $F_{\chi}=1.32$ and $Q_{\mathrm{F}}=0.28>\gamma_{\mathrm{F}}=0.001$.
-The $H_{0}$ hypothesis is not rejected.
-For the predictive data, the Fisher-test also confirms that the $g_{2,1,0}$ model is better than the $g_{3,1,0}$ model.
-Finally, we note that the three signal model $g_{3,1,0}$ is not unstable ($\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”) although the simulated data contains only two signals.
-For the $g_{2,1,0}$ model, the DCM detects the correct Model 3 signal frequencies (Table 3), but these frequencies are “too close” for the correct DFT detection (Equation 29).
-Therefore, the DFT prediction can not succeed.
-The analyses of all seven complex data sets indicate that the relative accuracy of amplitude and period estimates are lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 1 - 7).
-This statistical effect is the same when DCM is applied to the data of any arbitrary phenomenon.
-This effect would, for example, explain why our solar cycle amplitude predictions are less accurate than our solar cycle minimum and maximum epoch predictions (Jetsu 2025).
-DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct.
-The leakage of DFT spectral power [22, 15] would explain why correct signal periods have not been detected earlier from the sunspot data.
-4 Discussion
+3.11 Ill-posed problem
 The solution for the non-linear DCM model is an ill-posed problem (Equations 1-5).
@@
 Our technique can be applied to solve other non-linear models: Use the free parameters that make the model non-linear (Equation 9: $\bm{\beta}_{I}$) for solving the remaining other free parameters (Equation 2.1: $\bm{\beta}_{II}$).
-4.1 Existence ( $\mathrm{C}_{1}$)
+3.11.1 Existence ( $\mathrm{C}_{1}$)
 [13] transformed the original function into the frequency domain.
-The modern DFT time series analysis method transforms the original data into the frequency domain and gives the best frequency for a pure sine model.
-[14] presented the least squares method, which minimises the differences between the data and the linear model.
+The modern DFT time series analysis method transforms the original time series into the frequency domain and gives the best frequency for a pure sine model.
+[15] presented the LS method, which minimises the differences between the data and the linear model.
 Our DCM does the same by testing a large number of linear models.
 The data spacing, even or uneven, is irrelevant for these models.
-For every chosen DCM model, the number of tested linear models is
-$\binom{n_{\mathrm{L}}}{K_{1}}+(1+n_{\mathrm{B}})\times\binom{n_{\mathrm{S}}}{K_{1}}$, (46)
+For every chosen DCM model, the total number of tested linear models is
+$n_{\mathrm{Lin}}=\binom{n_{\mathrm{L}}}{K_{1}}+(1+n_{\mathrm{B}})\times\binom{n_{\mathrm{S}}}{K_{1}}$, (50)
 where the number of tested long and short search frequencies is $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$, respectively.
@@
 In other words, we solve this ill-posed problem by using brute computational force.
-If the data contain only zero mean white noise, the Gauß-Markov theorem ensures that a linear least squares fit solution exists for every tested frequency combination.
-4.2 Uniqueness ( $\mathrm{C}_{2}$)
+If the time series contains only zero mean white noise, the Gauß-Markov theorem ensures that a LS fit solution exists for every tested frequency combination.
+3.11.2 Uniqueness ( $\mathrm{C}_{2}$)
 We make the most of the Gauß-Markov theorem.
@@
 All possible $\bm{\beta}_{I}$ frequency combinations are tested (Equation 12).
-For every tested frequency combination $\bm{\beta}_{I}$, the linear model gives a unique value for the test statistic $z$ (Equations  15 and 16).
+For every tested frequency combination $\bm{\beta}_{I}$, the linear model gives a unique value for the test statistic $z$ (Equations  16 and 17).
 From all tested frequency combinations, we select the best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ which minimises $z$.
@@
 The best DCM model is not necessarily the correct model, if this correct model is not among the compared models.
-The correct model must be able the predict the future and past data.
-We formulate the Predictability-test for alternative DCM models (Equation 3.10).
-The order of Fisher- and Predictability-tests can be reversed.
-However, the former uses all data, while the latter uses a subset of all data, the predictive data.
+The correct model must be able to forecast the future and past data.
+We formulate the Forecast-test for alternative DCM models (Equation 3.12).
+The order of Fisher- and Forecast-tests can be reversed.
+However, the former uses all data, while the latter uses a subset of all data, the forecasting data.
 In ideal cases, both tests identify the same best and correct model, like in Section 3.12.
-4.3 Stability ( $\mathrm{C}_{3}$)
+3.11.3 Stability ( $\mathrm{C}_{3}$)
 The artificial bootstrap data sets (Equation 31) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation).
@@
 1a-b), the DCM model solution is stable because it does not change by increasing the number of tested frequencies $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$.
-Furthermore, the solutions for all seven simulated data sets converge when the $n$ and $\mathrm{SN}$} values increase (Tables 1 - 7).
+Furthermore, the solutions for all seven simulated time series converge when the $n$ and $\mathrm{SN}$} values increase (Tables 1 - 7).
 The different $n$ and $\mathrm{SN}$} combinations give the same stable DCM model solution.
-We conclude that our computational statistical DCM model solution fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
-4.4 Performance
-DCM does not suffer from the DFT limitations (Equations 27-30).
-The simulated complex data sets of Models 1-7 expose these DFT weaknesses, but in all these cases, DCM detects all periodic signals superimposed on an unknown trend (Sections 3.1 - 3.7).
-The correct signal(-s) and trend are detected.
-If there is enough $(n)$ very accurate $(\sigma_{i})$ data $(y_{i})$, DCM can predict the future or postdict the past by finding the correct DCM model for these data.
-We do admit that the $n$ and $\mathrm{SN}$} values of Models 2, 5 and 7 are extreme and unrealistic for most cases of real data.
-However, those model simulations are necessary for demonstrating that DCM can detect signals from short $\Delta T$ samples of high quality data.
-Those DCM signal detections depend only on the number of observations and their accuracy.
-Our DCM computes the $R$ (Equation 13) or $\chi^{2}$ (Equation 14) values for a massive number of linear models (Equation 46).
-The Gauß-Markov theorem ensures that the model having the lowest $R$ or $\chi^{2}$ is eventually always found, and the result for the non-linear iteration is unique (Equation 21).
-We could have simulated more complex data sets by increasing the values of $K_{1}$ (number of signals), $K_{2}$ (signal shape) and $K_{3}$ (polynomial order of trend).
-DCM would eventually detect those signals and the trend when the sampled data sets get larger $(n)$ and/or the signal to noise ratio becomes higher $({\mathrm{SN}})$.
-In this sense, DCM has no limit for the complexity of the data sets.
-In the real world, it may be impossible to eliminate noise $({\mathrm{SN}})$, but it is possible increase the sample size $(n)$.
-Since the time span $(\Delta T)$ is irrelevant, the best solution is fast sampling of accurate data sets.
-For example, many observers could measure the same phenomenon simultaneously.
-We recommend such fast sampling combined to DCM parallel computation Python code analysis.
-We want to draw attention to three aspects of DCM computational performance.
+We conclude that our computational statistical DCM model solution fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 3.11.1 - 3.11.3).
+3.12 Forecast
+There are numerous techniques for forecasting a time series [e.g., 20, 22, 26].
+The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast.
+We divide all data into the forecasting data and the forecasted data.
+The time points, the observations and the errors of these samples are
+  • $n$ forecasting data values $t_{i}$, $y_{i}$ and $\sigma_{i}$
+  • $n^{\prime}$ forecasted data values $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$
+The DCM gives the best forecasting data model
+$g_{i}=g(t_{i},\bm{\beta}_{\mathrm{Fore}}),$ (51)
+where $\bm{\beta}_{\mathrm{Fore}}$ are the free parameter values.
+The $t_{\mathrm{mid,Fore}}\!=\!(t_{n}\!+\!t_{1})/2$ and $\Delta T_{\mathrm{Fore}}\!=\!t_{n}\!-\!t_{1}$ values are computed from the forecasting data time points $t_{i}$.
+The forecasting model values at any arbitrary time $t$ can be obtained from the $\bm{\beta}_{\mathrm{Fore}}$ and $t_{\mathrm{mid,Fore}}$ $\Delta T_{\mathrm{Fore}}$ values.
+The forecasted data model values are
+$g^{\prime}_{i}$ $=$ $g(t^{\prime}_{i},\bm{\beta}_{\mathrm{Fore}}),$ (52)
+where $t_{\mathrm{mid}}=t_{\mathrm{mid,Fore}}$ and $\Delta T=\Delta T_{\mathrm{Fore}}$.
+We do not compute “new” $t_{\mathrm{mid}}$ and $\Delta T$ values from the forecasted data time points $t^{\prime}_{i}$ because the correct $g^{\prime}_{i}$ values are obtained only from the $\bm{\beta}_{\mathrm{Fore}}$, $t_{\mathrm{mid,Fore}}$ and $\Delta T_{\mathrm{Fore}}$ combination of the forecasting model (Equation 52).
+The $n^{\prime}$ forecasted data model residuals
+$\epsilon^{\prime}_{i}=y^{\prime}_{i}-g^{\prime}_{i}$
+give the forecasted data test statistic
+$z_{\mathrm{Fore}}$ $=$ z test statistic for forecasted
+data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ (Equation 16 or 17).
+This parameter $z_{\mathrm{Fore}}$ measures how well the forecast (Equation 52) obtained from the forecasting data works for the forecasted data.
+If the DCM detects a new signal from the forecasting data, there are two alternatives:
+  • If this new signal is real, the $z_{\mathrm{Fore}}$ value of forecasted data decreases.
+  • If this new signal is unreal, the $z_{\mathrm{Fore}}$ value of forecasted data increases.
+This “Forecast-test” technique revealed at least five real signals in the sunspot record (Jetsu 2025).
+We compute the $z_{\mathrm{Fore}}$ parameter value from the known forecasted data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ values (Equation 3.12).
+Forecasts are possible even if all $t^{\prime}_{i}$, $y^{\prime}_{i}$ or $\sigma^{\prime}_{i}$ forecasted data values are unknown.
+In this case, the $t_{i}$, $y_{i}$ and $\sigma_{i}$ values of all data can be used as forecasting data.
+The best DCM model $g(t_{i},\bm{\beta}_{\mathrm{Fore}})$ for all data determines the correct $\bm{\beta}_{\mathrm{Fore}}$, $t_{\mathrm{mid,Fore}}=(t_{n}+t_{1})/2$ and $\Delta T_{\mathrm{Fore}}=t_{n}-t_{1}$ combination of the forecasting model.
+The $n^{\prime}$ forecasted data time points $t_{i}^{\prime}$ can be created for any arbitrary chosen sample window $\Delta T^{\prime}=t_{n}^{\prime}-t_{i}^{\prime}$.
+The $n^{\prime}$ forecasted $g_{i}^{\prime}=g^{\prime}_{i}(t_{i}^{\prime},\bm{\beta}_{\mathrm{Fore}})$ values are obtained from Equation 52.
+These $g^{\prime}_{i}$ values can be used, for example, to compute the forecasted mean level
+$m_{\mathrm{Fore}}={{1}\over{n^{\prime}}}\sum_{i=1}^{n^{\prime}}g^{\prime}_{i}$ (54)
+during the chosen sample window $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$.
+Jetsu (2025) used this $m_{\mathrm{Fore}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720.
+Since the known mean level of all sunspot data was $m$, Jetsu (2025) used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
+  • If DCM detects a real new signal in all data, the $m_{\mathrm{Fore}}$ value decreases.
+  • If DCM detects many real signals in all data, the $m_{\mathrm{Fore}}$ value falls below $m$.
+  • If DCM detects an unreal new signal in all data, the $m_{\mathrm{Fore}}$ value increases.
+We use Model 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data (Table 3, column 2) to illustrate the DCM forecasting technique.
+The black dots in Figures 8a-c are the first $n=40$ forecasting data values $t_{i}$, $y_{i}$ and $\sigma_{i}$.
+The open black dots denote the $n^{\prime}=10$ forecasted data values $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$.
+The continuous black line is the forecasting model $g(t)$ and the dotted black line is the forecasted model $g^{\prime}(t)$.
+The red dotted line shows the $\pm 3\sigma$ errors of both models.
+The blue dots are the forecasting data $\epsilon_{i}=y_{i}-g_{i}$ residuals and the open blue dots are the forecasted data $\epsilon_{i}^{\prime}=y_{i}^{\prime}-g_{i}^{\prime}$ residuals.
+Model 3 is the sum of two $P_{1}=0.16$ and $P_{2}=0.17$ pure sine signals $(K_{1}=2,K_{2}=1)$ superimposed on the constant mean level $M_{0}=1$ $(K_{3}=0)$.
+We compute the $z$ and $z_{\mathrm{Fore}}$ values for the $g_{1,1,0}$, $g_{2,1,0}$ and $g_{3,1,0}$ models (Table 10).
+These three models have the same $K_{2}=1$ and $K_{3}=0$ orders as Model 3, but their signal numbers $K_{1}=1,2$ or 3 are different, the $g_{2,1,0}$ model being the correct simulation Model 3.
+The correct $g_{2,1,0}$ model gives the smallest $z_{\mathrm{Fore}}=0.88$ value (Table 10).
+Therefore, it is a better forecasting model than the $g_{1,1,0}$ and $g_{3,1,0}$ models.
+The one signal $g_{1,1,0}$ model forecast fails because the blue open circles denoting the forecasted data residuals $\epsilon_{i}^{\prime}=y_{i}^{\prime}-g_{i}^{\prime}$ show large deviations from to the blue dotted line offset level of $\epsilon_{i}^{\prime}=-1$ (Figure 8a).
+The two signal $g_{2,1,0}$ model forecast succeeds because all open blue dots denoting the forecasted data residuals stay close to the blue dotted line offset level $\epsilon_{i}^{\prime}=-1$, as well as inside the red dotted $\pm 3\sigma$ model error limits (Figure 8b).
+The three signal model $g_{3,1,0}$ periods are $P_{1}=0.0643\pm 0.0018$, $P_{2}=0.1592\pm 0.0012$ and $P_{3}=0.1716\pm 0.0012$.
+The amplitudes of these pure sine signals are $A_{1}=0.092\pm 0.025$, $A_{2}=1.879\pm 0.16$ and $A_{3}=1.82\pm 0.15$.
+The periods and the amplitudes of the two strongest $P_{2}$ and $P_{3}$ signals are correct because they are the same as in Model 3 simulation (Table 3, column 1).
+Therefore, the $g_{3,1,0}$ forecast appears nearly as good as the $g_{2,1,0}$ forecast because the residuals $\epsilon_{i}^{\prime}$ are close to the offset level of $\epsilon_{i}^{\prime}=-1$ (blue open dots), and these residuals also stay inside the red dotted $\pm 3\sigma$ model error limits (Figure 8c).
+The amplitude $A_{1}=0.092$ of the third $P_{1}=0.0643$ signal is very low.
+Due this weak “unreal” $P_{1}$ signal, the $g_{3,1,0}$ model has a larger $z_{\mathrm{Fore}}=0.91$ value than the correct $g_{2,1,0}$ model (Table 10).
+Finally, we note that the three signal model $g_{3,1,0}$ is not unstable ($\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”) although the simulated time series contains only two signals.
+For the forecasting data, the extreme Fisher-test critical levels $Q_{\mathrm{F}}<10^{-13}$ confirm that the two signal $g_{2,1,0}$ and the three signal $g_{3,1,0}$ models are certainly better than the one signal $g_{1,1,0}$ model.
+The forecasting data parameters $n=40$, $\eta_{1}=7$, $\eta_{2}=10$, $\chi_{1}=22.946$ and $\chi_{2}=19.790$ for the simple $g_{2,1,0}$ model and the complex $g_{3,1,0}$ model give the Fisher-test critical level $Q_{\mathrm{F}}=0.28>\gamma_{\mathrm{F}}=0.001$.
+The $g_{2,1,0}$ model beats the $g_{3,1,0}$ model because the $H_{0}$ hypothesis is not rejected.
+Hence, the Fisher-test confirms that the $g_{2,1,0}$ model is the best model for the forecasting data.
+The best DCM model (Fisher-test, $g_{2,1,0}$ ) is also the correct DCM model (Forecast-test, $g_{2,1,0}$).
+The double-check works for this particular Model 3 simulated time series!
+The DFT can not detect the correct frequencies for the forecasting data because the simulated frequencies are “too close” (Equation 37). Therefore, the DFT forecast can not succeed.
+The analyses of all seven complex time series indicate that the relative accuracy of amplitude and period estimates are lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 1 - 7).
+This statistical effect is the same when DCM is applied to the time series of any arbitrary phenomenon.
+This effect would, for example, explain why our solar cycle amplitude forecasts are less accurate than our solar cycle minimum and maximum epoch forecasts (Jetsu 2025).
+DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct.
+The leakage of DFT spectral power [25, 17] would explain why correct signal periods have not been detected earlier from the sunspot data.
+The DCM analysis results can be double-checked.
+We use the Fisher-test to identify the best DCM model from all tested DCM models (Equation 48).
+The correct model may not be among the tested models.
+Hence, the best model is not necessarily the correct one.
+This best model must be correct if it passes the Forecast-test (Equation 3.12).
+This Fisher- and Forecast-test combination double-checks the DCM analysis results.
+The perfect result is that the best model is the correct model.
+Therefore, the Forecast-test also prevents overfitting.
+4 Discussion
+DCM proceeds through two stages.
+It computes the $R$ (Equation 14) or $\chi^{2}$ (Equation 15) values for a massive number of LS fits (Equation 50).
+The Gauß-Markov theorem ensures that the DCM model having the lowest $R$ or $\chi^{2}$ is inevitably always found, and the result for the non-linear iteration is unique (Equation 22).
+This first stage devours CPU.
+The second fast stage, the Fisher-test, then reveals the best DCM model alternative (Equation 48).
+Practical time series analysis applications, like digital signal processing, require fast computational algorithms [25].
+Long computation time is the main AL constraining DCM.
+For example, Jetsu (2025) computed about 4.5 million LS fits to search for four signals in the monthly sunspot data $(n=3287,K_{1}=4,K_{2}=1,K_{3}=0,n_{\mathrm{L}}=100,n_{\mathrm{S}}=30,n_{\mathrm{B}}=20)$.
+This required several months of CPU.
+The parallel Python code computations took a few days.
+Nevertheless, the CPU used is beside the point if a well-posed computational solution can be found to any challenging scientific ill-posed problem.
+The DCM computation time AL constraint is amply compensated in real observations because there is no need to wait for the repetition of the signal(-s).
+The WD-effect is the spearhead of DCM because the sample window $(\Delta T)$ can be infinitesimally short.
+Fast accurate observations are the best approach.
+If the noise $(\sigma\equiv{\mathrm{SN}})$ can not be eliminated it is always possible to increase the sample size $(n)$.
+It is time to summarise why the ALs of other frequency-domain parametric time series analysis methods (Section 1) do not constrain DCM.
   1. 1.
-    Practical time series analysis applications, like digital signal processing, require fast computational algorithms [22].
-Our computational statistical solution for the non-linear DCM model requires a lot of CPU.
-Yet, the CPU used is beside the point if a well-posed computational solution can be found to any challenging scientific non-linear problem.
-DCM proceeds through two stages.
-For every chosen alternative model, DCM performs a massive number of linear least squares fits (Equation 18) to find the unique initial free parameter values for the non-linear iteration (Equation 21).
-This first stage devours CPU.
-The second fast stage, the Fisher-test, then reveals the best model alternative (Equation 40).
+    Data errors (level of noise) are unknown.
+    • DCM can solve both alternatives: errors $\sigma_{i}$ known or unknown (Equations 16, 17, 46 and 47).
+All DCM model parameter solutions in Tables 1-7 converge to the correct values when the simulated data $n$ and/or $\mathrm{SN}$} increase ($\sigma_{i}$ decrease).
+Due to this WD-effect, sufficient noise reduction always leads to the correct model detection.
   2. 2.
-    Our frequency space symmetry idea dramatically reduces CPU (Equation 11).
-For example, were this symmetry not exploited, the four signal $K_{1}=4$ search would give $4!=24$ exactly the same solutions from four-dimensional tested grid in frequency space.
-This would cause unnecessary use of CPU in the preliminary testing all possible frequency combinations inside a tesseract (a four-dimensional cube).
-After one necessary iteration of Equation 21, and 23 unnecessary ones, the identification of the best frequency combination would be a daunting exercise.
+    Data error information is not utilised.
+    • DCM utilises this information.
+If the errors $\sigma_{i}$ are known, DCM performs weighted LS fits which minimise the model $\chi^{2}$ for every tested frequency combination (Equation 15).
+This gives the DCM test statistic $z$ (Equation 17).
+The Fisher-test utilises the $\chi^{2}$ values of all alternative tested DCM models to identify the best DCM model (Equation 47).
   3. 3.
-    The DCM signal detections depend only on the sample size $n$ and the signal to noise ratio $\mathrm{SN}$}, but not on the time span $\Delta T$.
-This dimensional effect makes sense.
-If the $t_{i}$ and $y_{i}$ data are inside a rectangle $\Delta T\times\Delta y=(t_{n}-t_{1})\times[{\mathrm{max}}(y_{i})-{\mathrm{min}}(y_{i})]$, then the $y_{i}/\sigma_{i}$ ratio reveals or hides the $g_{i}$ model details.
-Since the time span of data $\Delta T$ is irrelevant, DCM is an ideal forecasting method.
-We have been deluded into unnecessary waiting for the repetition of the signal(-s).
-DCM can perceive the future and the past much “earlier” (from shorter $\Delta T$) than previously thought, like in Equations 27 and 29.
-This should motivate fast sampling of large high quality data sets and their analysis using the DCM parallel Python code.
-The invention of computers has made it possible to use this computational statistical method, which is probably the simplest possible one.
+    Data must be evenly spaced.
+    • DCM performance is independent of data spacing.
+Even or uneven spacing is irrelevant for all $n_{\mathrm{Lin}}$ LS fits (Equation 50).
+However, we do admit that long gaps can mislead even these LS fits.
+  4. 4.
+    Model parameter errors are unknown.
+    • DCM gives error estimates for the model parameters of Equations 23-29.
+For non-linear models, such as the DCM model, the analytical solution for the model parameter errors is a highly complex effort [12].
+DCM solves these error estimates using the computational statistical bootstrap technique (Equation 31).
+  5. 5.
+    Model and forecast errors are unknown.
+    • The computational bootstrap technique (Equation 31) gives $n_{B}$ estimates for $\bm{\beta}_{J}$, $t_{\mathrm{mid,J}}$ and $\Delta T_{J}$, where $J=1,2,...n_{B}$.
+These estimates give $n_{B}$ values for $g(t)$ at any time $t$ inside and outside the sample window $\Delta T$.
+The standard deviation of these $n_{B}$ values gives the error limits $\sigma_{g(t)}$ in Equation 30.
+The $n_{B}$ estimates for $\bm{\beta}_{J}$, $t_{\mathrm{mid,J}}$ and $\Delta T_{J}$ also give the error limits for the functions $h(t),h_{i}(t)$ and $p(t)$ in the DCM model $g(t)$.
+  6. 6.
+    Sample window is shorter than signal period(s).
+    • DCM solves this problem.
+The sample window is shorter than the period(s) in Model 1, 2, 5 and 7 time series simulations.
+Regardless of this, all model parameter estimates converge to correct values when $n$ and/or $\mathrm{SN}$} increase (Tables 1, 2, 5 and 7).
+Due to this WD-effect, the sample window $\Delta T$ value is irrelevant for DCM.
+It can perceive the future and the past much “earlier” (from shorter $\Delta T$) than has been previously thought, like in Equations 35 and 37.
+  7. 7.
+    Presence and shape of trend are unknown.
+    • DCM can test any $K_{1}$, $K_{2}$ and $K_{3}$ model combination.
+The Fisher-test reveals the combination of the best DCM model.
+The Forecast-test can double-check this result.
+These tests give the correct $K_{3}$ trend order.
+Due to the WD-effect, the polynomial trend coefficients $M_{k}$ converge to correct values in all simulated time series (Sections 3.1-3.7).
+The trend is absent (stationary time series) if $K_{3}=-1\equiv$ $p(t)=0$ or $K_{3}=0\equiv$ $p(t)=M_{0}=\text{constant}$ (Equations 4 and 5).
+  8. 8.
+    Sample window causes leakage.
+    • Leakage does not constrain DCM because the sample window $\Delta T$ has no effect on the LS fits (Equation 50: $n_{\mathrm{Lin}}$).
+All frequencies converge to exactly correct values when the $n$ and/or $\mathrm{SN}$} values of simulated time series increase (Tables 1-7).
+This confirms the absence of leakage.
+Due to the WD-effect, the performance of DCM does not depend on the sample window $\Delta T$.
+  9. 9.
+    Leakage weakens frequency resolution.
+    • There is no leakage because the sample window is irrelevant (AL8).
+The DCM frequency resolution is limited only by the sample size $n$ and the data accuracy $\sigma$.
+The WD-effect ensures that all frequency detections for the simulated time series converge to exactly correct values (Tables 1-7).
+  10. 10.
+    Signal shapes are not pure sines.
+    • The DCM always finds the correct $K_{1}$, $K_{2}$ and $K_{3}$ values (AL7).
+The $K_{2}$ value determines the signal shape.
+The $K_{2}=1$ pure sine shape is the simplest.
+Higher $K_{2}$ values allow the DCM modelling of more complex shapes.
+We demonstrate the $K_{2}=2$ double wave signal detections in Sections 3.6 and 3.7.
+  11. 11.
+    Number of signals is unknown.
+[…]
+  12. 12.
+    Correct model alternative is unknown.
+    • The values of $K_{1}$, $K_{2}$ and $K_{3}$ determine the correct model (AL7).
+The $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty terms prevent overfitting when the Fisher-test is used to identify the best DCM model from all tested DCM models (Equations 46 and 47).
+This best model is also the correct model if it passes the Forecast-test (Equation 3.12).
+The WD-effect ensures that better data can inevitably reveal the correct model.
+  13. 13.
+    Signal significances are unknown.
+    • The Fisher-test gives the $Q_{\mathrm{F}}$ critical levels (signal significances) for all detected signals (Sections 3.9 and 3.10).
+This $Q_{\mathrm{F}}$ is the probability of falsely rejecting the $H_{0}$ hypothesis when it is in fact true.
+Thus, the $Q_{\mathrm{F}}$ value represents the probability of false signal detection.
+This detection is absolutely certain in the $Q_{\mathrm{F}}<\!10^{-16}$ cases.
+Our pre-assigned significance level for signal detection is $\gamma=0.001$ (Equation 48).
+  14. 14.
+    Model solution is ill-posed.
     • The analytical solution for the non-linear DCM model is ill-posed (Equation 1).
-Regardless of the data set complexity, DCM can find a well-posed computational statistical solution for this problem (Equation 21).
-So far, we have compared DCM only to DFT.
-Both methods are parametric and operate in the frequency-domain.
-There are innumerable similar parametric time series analysis methods [e.g.
-22, 15].
-Each one of these other time series analysis methods suffers from their own particular application limitations.
-We list some of those application limitations below.
-  1. 1.
-    The required data spacing is even or uneven.
-  2. 2.
-    The number of signals is unknown.
-  3. 3.
-    The presence and shape of trend are unknown.
-  4. 4.
-    The degree of noise is unknown.
-  5. 5.
-    The sample window causes leakage.
-  6. 6.
-    The leakage weakens frequency resolution.
-  7. 7.
-    The signal frequencies are too close.
-  8. 8.
-    The signal periods exceed the data time span.
-  9. 9.
-    The signal shapes are not pure sinusoids.
-  10. 10.
-    The forecasts of non-linear models fail.
-DCM does not suffer from any of these application limitations. This simple computational statistical method can fail if, and only if, the Gauß-Markov theorem is not valid. That is impossible.
+We show that there exists a well-posed computational solution (Section 3.11).
+The WD-effect ensures that this solution can always be found by increasing the sample size $(n)$ and/or the signal to noise ratio $({\mathrm{SN}})$, regardless of the time series complexity ($K_{1}$, $K_{2}$ and $K_{3}$ combination).
+The main DCM constraint is that the solutions for more complex models require more computation time when the number of signals increases (Equation 50).
+  15. 15.
+    Complex non-linear model forecasts fail.
+    • Our computational solution for the non-linear DCM model is well-posed (Section 3.11).
+The forecast of this well-posed solution can be double-checked (Section 3.12).
+First, the Fisher-test identifies the best DCM model from all tested DCM models (Section 3.9).
+Then, the Forecast-test reveals if this best DCM model is also the correct one (Equation 3.12).
+Again, the WD-effect ensures that the correct model and forecast can be found for any complex time series ($K_{1}$, $K_{2}$ and $K_{3}$ combination) when $n$ and/or $\mathrm{SN}$} increase.
+We conclude that the DCM rises to meet the challenge of “forecasting the evolution of complex systems” [7].
+The DCM turns things upside down.
+There would be no need for time series analysis, if the correct frequencies were already known.
+The tested frequencies are already known.
+DCM does not search for unknown frequencies, it just tests known frequencies.
+DCM tests all possible frequency combinations.
+The LS fit for every frequency combination gives unique $R$ and $\chi^{2}$ values.
+The best frequency combination minimises $R$ or $\chi^{2}$.
+This gives the unique initial free parameter $\bm{\beta}$ values for the non-linear iteration.
+The frequencies are never unknown in this process.
+The Gauss-Markov theorem, the well-posed computational model solution and the revolutionary WD-effect make absolutely sure that DCM succeeds for any number of signals $(K_{1})$, all signal shapes $(K_{2})$ and every polynomial trend $(K_{3})$.
+The Fisher-test and the Forecast-test can double-check that the DCM model solution is correct.
+The DCM analysis of intensive, large and accurate time series can see through time: a glimpse of the future and the past.
+Our DCM is a remarkable method because it outperforms the DFT.
+Gauß (1777–1855) developed an early form of the Fast Fourier Transform (FFT) algorithm for astronomical purposes, but did he not publish it.
+Perhaps our DCM should be re-named the Slow Gauß Transform (SGT).
 […]
@@
 […]
-Author contribution Lauri Jetsu created the simulated data sets, performed the DCM and DFT analysis and wrote the manuscript.
+Author contribution Lauri Jetsu performed this research and wrote the manuscript.
 […]
-Consent for publication The author of this manuscript grants permission for the publication in Computational Statistics.
+Consent for publication The author of this manuscript grants permission for the publication in Journal of Statistical Theory and Practice.
 […]

### v5

@@
 Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series.
+In our appendix, we show that DCM can model and forecast El Niño.
 keywords:
-Time series analysis, Non-linear and non-stationary processes, Ill-posed problems, Computational statistics, Forecasting, Discrete Chi-square Method, Discrete Fourier Transform
+Time series analysis, Non-linear and non-stationary processes, Ill-posed problems, Computational statistics, Forecasting, Discrete Chi-square Method, Discrete Fourier Transform, El Niño
 1 Introduction
@@
 (g) DFT model detrended $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (black continuous line), DFT pure sine $s_{\mathrm{DFT}}(t)$ (red thick continuous line), detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and DFT model residuals (blue dots) offset to -1.5 level (blue dotted line).
-The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [24, TSPA, Equation 1] and the Continuous Period Search [29, CPS, Equation 3].
+The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [29, TSPA, Equation 1] and the Continuous Period Search [36, CPS, Equation 3].
 TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$.
@@
 We search for signals in these simulated time series by using the frequently applied DFT version formulated by [23] [^2] This paper had over 2600 citations in December 2025.
-Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [21, Equation 1].
+Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [23, Equation 1].
 The notations for the DFT model are
@@
 Our DFT analyses of simulated time series may fail due to AL6-AL10.
-The time series is “too short” [25, AL6]
+The time series is “too short” [30, AL6]
 $P>\Delta T.$ (35)
-The time series is non-stationary because it contains “a trend” [33, AL7]
+The time series is non-stationary because it contains “a trend” [42, AL7]
 $p_{\mathrm{DFT}}(t)\neq 0.$ (36)
-Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [25, AL9].
+Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [30, AL9].
 $|f_{1}-f_{2}|<f_{0}={\Delta T}^{-1}.$ (37)
@@
 For the Model 1 simulated $n=100$ and ${\mathrm{SN}}=100$ combination, Table 8 shows that this time series contains only one signal.
-For this Model 1 time series, there is no need to discuss the DFT significance estimates [21, their Equation 22] because this method fails to detect the correct period.
+For this Model 1 time series, there is no need to discuss the DFT significance estimates [23, their Equation 22] because this method fails to detect the correct period.
 We conclude that the Fisher-test identifies the correct DCM model (Section 3.9), as well as gives the signal significance $Q_{\mathrm{F}}$ estimates (Section 3.10).
@@
 3.12 Forecast
-There are numerous techniques for forecasting a time series [e.g., 20, 22, 26].
+There are numerous techniques for forecasting a time series [e.g., 22, 26, 31].
 The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast.
@@
 Perhaps our DCM should be re-named the Slow Gauß Transform (SGT).
-[…]
-Acknowledgements This work has made use of NASA’s Astrophysics Data System (ADS).
-[…]
-Author contribution Lauri Jetsu performed this research and wrote the manuscript.
-[…]
-Consent for publication The author of this manuscript grants permission for the publication in Journal of Statistical Theory and Practice.
-[…]
-Data availability All data files and all DCM control files are stored to
-[…]
+5 Appendix
+In this appendix, we show that our DCM can model and forecast the El Niño phenomenon between the years 1870 and 2024.
+This result may be the “holy grail” of climatology.
+The solar forcing most probably causes the detected $5.850\pm 0.085$, $12.82\pm 0.40$ and $19.30\pm 0.83$ year signals.
+These strict periodicities indicate that the solar cycle is caused by the planets.
+Our results confirm that the DCM is an ideal time series analysis method for detecting an unknown number of signals superimposed on an unkown trend.
+The El Niño phenomenon is considered unpredictable [52, 39, 6, 25, 32, 35, 50, 40].
+Our analysed data are the yearly means in the The Seasonal Mean Nino 4 HadISST1.1 (NOAA PSL) sample.
+[^5] The data from https://psl.noaa.gov/data/timeseries/month/data/nino4.long.anom.data were retrieved on January 29th, 2026.
+There were no values for the last four months of 2025.
+The full data contains $n=155$ yearly means between 1870 and 2024.
+5.1 Trend
+We detect significant signals only from the weighted data.
+Therefore, the trend and the forecasts are solved from the weighted data.
+The weighted DCM analysis results for all data are given in Table S1.
+For one, two, three and four signals, the best trend alternative is the linear trend $K_{1}=1$.
+We use this $K_{1}=1$ trend in our DCM analysis of El Niño data.
+For the $\mathcal{M}$=3 model, the positive linear trend is $2M_{1}=0.56\pm 0.20$ $\mathrm{C}^{o}$during $\Delta T=154$ years (Equation 5), like in all $\mathcal{M}$=1-4 models.
+5.2 Forecasting data
+We use the first half of all data $(n=78)$ to forecast the second half $(n=77)$.
+The results of weighted DCM analysis are given in Table S2.
+The periods of the three strongest signals are $5.850\pm 0.085$, $12.82\pm 0.40$ and $19.30\pm 0.83$ years (Model $\mathcal{M}$=3).
+The forecasts of the two and three signals models, open circles, are shown in Figures S1 and S2.
+For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signals models are 421, 336, 370 and 428, respectively.
+Hence, the two signal $\mathcal{M}$=2 model gives the best forecast.
+5.3 All data
+The DCM analysis results for all weighted data are given in Table S3.
+The three signals $\mathcal{M}$=3 model is the best.
+The four signal model is unstable $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}})$.
+The three and four signal models are shown in Figures S3 and S4.
+For both alternatives, the predicted values, open circles, after 2024 are given in Table S4.
+Both alternatives give essentially the same forecast.
+For the sake of consistency, we also give the results of non-weighted analysis of all data (Table S5).
+We detect the same signals as from non-weighted data, but the signal significances $Q_{\mathrm{F}}$ are lower.
+This confirms that the data error information is crucial.
+5.4 Main result
+Our DCM can model and forecast the El Niño phenomenon.
+The solar forcing is the only possible explanation for these results.
+Within their error limits, the detected three El Niño signals could represent the Hale cycle [24], the solar cycle[46, 56] and half the solar cycle.
+These strict El Niño periodicities indicate that the planets could cause a stationary and deterministic solar cycle [28].
+A stochastic and non-stationary solar cycle [8, 54] could not cause the detected regular El Niño periodicities.
+Acknowledgements
+This work has made use of NASA’s Astrophysics Data System (ADS) services.
+Author contributions statement
+L.J. peformed this research and wrote the manuscript.
+Consent for publication
+The author grants permission for publication.
+Additional information
+Supplementary information The online version contains supplementary material available.
+Data availability
+Main manuscript analysis: All data files and all DCM control files are stored to
 https://zenodo.org/uploads/17018676.
-There is also a manual of how to use those codes.
+Appendix analysis: All files, DCM Python code, DCM control files and Python code are stored to the
+https://zenodo.org/records/18595333.
 Declarations
-[…]
-[…]
+Competing interests
+The author declares no competing interests.
+Correspondence and requests for materials should be addressed to L.J.
+Funding No funding was received.
+Ethical approval and consent to participate Not applicable.
+Supplementary material
+The $\mathcal{M}$= 1 analysis in Table S1 can be repeated using these two commands
+cp longC14K110.dat dcm.dat
+python dcm.py
+The whole El Niño analysis can be repeated in similar fashion.
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S1: First half forecast (Table S2: $\mathcal{M}$=2): Forecasting data (closed black circles) and forecasted data (open black circles).
+Green and red dottted lines denote one and three sigma model $g(t)$ errors.
+Residuals (closed and open blue circles) are offset to level of -3.
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S2: First half forecast (Table S2: $\mathcal{M}$=3). Otherwise as in Figure S1.
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S3: All data model $\mathcal{M}$=3 in Table S3. Otherwise as in Figure S1.
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S4: All data model $\mathcal{M}$=4 in Table S3. Otherwise as in Figure S1.
+Table S1: Trend in all data.
+Weighted DCM analysis performed between $P_{\mathrm{min}}=3$ and $P_{\mathrm{max}}=100$.
+(1) $\mathcal{M}$model, (2) $\eta$ free parameters and $\chi^{2}$ of model, (3) First signal: $P_{1}$ period, $A_{1}$ peak to peak amplitude and $t_{\mathrm{min},1}$ primary minimum epoch.
+(4-6) Next signals.
+(7-8) Fisher-test: $F_{\chi}$ test statistic and $Q_{\mathrm{F}}$ critical level.
+(9) Control file.
+Arrows point to best model in Fisher-test comparison.
+Period analysis Fisher-test
+Data: Original weighted data ($n=155,\Delta T=154$: Clong.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{\chi}$ (-) $F_{\chi}$ (-) Control file
+$\chi^{2}$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9)
+One signal
+$\mathcal{M}$=1 $12.67\pm 0.33$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$$_{1,1,0,\chi^{2}}$ 4 $0.62\pm 0.15$ - - - 26.9 14.0 longC14K111.dat
+516 $1871.2\pm 1.3$ - - - $6.7\times 10^{7}$ $2.7\times 10^{-6}$
+$\mathcal{M}$=2 $12.65\pm 0.20$ - - - - $\leftarrow$
+$\mathcal{M}$$_{1,1,1,\chi^{2}}$ 5 $0.51\pm 0.14$ - - - - 1.02 longC14K111.dat
+437 $1871.3\pm 1.3$ - - - - 0.31
+$\mathcal{M}$=3 $12.65\pm 0.43$ - - -
+$\mathcal{M}$$_{1,1,2,\chi^{2}}$ 6 $0.52\pm 0.12$ - - - longC41K112.dat
+434 $1871.4\pm 1.1$ - - -
+Two signals
+$\mathcal{M}$=5 $\mathcal{M}$=6
+$\mathcal{M}$=4 $12.70\pm 0.17$ $26.8\pm 2.0$ - - $\uparrow$ $\uparrow$
+$\mathcal{M}$$_{2,1,0,\chi^{2}}$ 7 $0.62\pm 0.12$ $0.39\pm 0.16$ - - 27.4 13.6 longC14K210.dat
+456 $1870.8\pm 1.0$ $1870.3\pm 3.4$ - - $5.6\times 10^{-7}$ $3.8\times 10^{-6}$
+$\mathcal{M}$=5 $5.677\pm 0.089$ $12.64\pm 0.31$ - - - $\leftarrow$
+$\mathcal{M}$$_{2,1,1,\chi^{2}}$ 8 $0.393\pm 0.086$ $0.47\pm 0.13$ - - - 0 longC14K211.dat
+384 $1875.65\pm.56$ $1871.3\pm 1.2$ - - - 1
+$\mathcal{M}$=6 $5.677\pm 0.077$ $12.64\pm 0.42$ - - -
+$\mathcal{M}$$_{2,1,2,\chi^{2}}$ 9 $0.388\pm 0.093$ $0.47\pm 0.12$ - - - longC41K212.dat
+384 $1875.66\pm 0.76$ $1871.4\pm 1.2$ - - -
+Three signals
+$\mathcal{M}$=8 $\mathcal{M}$=9
+$\mathcal{M}$=7 $5.619\pm 0.091$ $12.74\pm 0.48$ $21.83\pm 0.81$ - $\uparrow$ $\uparrow$
+$\mathcal{M}$$_{3,1,0,\chi^{2}}$ 10 $0.41\pm 0.13$ $0.52\pm 0.11$ $0.40\pm 0.12$ - 29.2 15.0 longC14K310.dat
+407 $1871.04\pm 0.77$ $1870.7\pm 1.7$ $1870.0\pm 3.2$ - $2.6\times 10^{-7}$ $1.2\times 10^{-6}$
+$\mathcal{M}$=8 $5.662\pm 0.077$ $12.77\pm 0.12$ $21.3\pm 1.5$ - - $\leftarrow$
+$\mathcal{M}$$_{3,1,1,\chi^{2}}$ 11 $0.40\pm 0.13$ $0.47\pm 0.10$ $0.37\pm 0.12$ - - 0.82 longC14K311.dat
+338 $1870.3\pm 0.62$ $1870.4\pm 0.87$ $1870.8\pm 3.2$ - - 0.36
+$\mathcal{M}$=9 $5.661\pm 0.073$ $76\pm 0.22 21.2±1.$$2\pm 1.2 - - -$
+$\mathcal{M}$$_{3,1,2,\chi^{2}}$ 12 $0.40\pm 0.11$ $0.49\pm 0.10$ $0.39\pm 0.12$ - - - longC41K312.dat
+336 $1870.3\pm 0.74$ $1870.6\pm 1.2$ $1871.1\pm 2.8$ - - -
+Four signals
+$\mathcal{M}$=11 $\mathcal{M}$=12
+$\mathcal{M}$=10 $5.597\pm 0.066$ $12.77\pm 0.19$ $19.4\pm 1.6~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $21.5\pm 2.1~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $\uparrow$ $\uparrow$
+$\mathcal{M}$$_{4,1,0,\chi^{2}}$ 13 $0.44\pm 0.10$ $0.57\pm 0.11$ $0.4\pm 1.3~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.4\pm 1.4~AD$ 28.1 14.2 longC14K410.dat
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 362 $1871.56\pm 0.57$ $1870.3\pm 1.2$ $1872.8\pm 3.7$ $1871.8\pm 3.3$ $4.5\times 10^{-7}$ $2.5\times 10^{-6}$
+$\mathcal{M}$=11 $5.496\pm 0.093~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $5.66\pm 0.13~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $12.76\pm 0.15$ $21.1\pm 1.5$ - $\leftarrow$
+$\mathcal{M}$$_{4,1,1,\chi^{2}}$ 14 $0.3\pm 1.8~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.4\pm 1.8~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.49\pm 0.13$ $0.39\pm 0.11$ - 0.45 longC14K411.dat
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 300 $1873.8\pm 1.2$ $1870.2\pm 1.2$ $1870.60\pm 0.97$ $1871.3\pm 2.1$ - 0.50
+$\mathcal{M}$=12 $5.497\pm 0.087~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $0.566\pm 0.062~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $12.75\pm 0.17$ $21.0\pm 1.0$ - -
+$\mathcal{M}$$_{4,1,2,\chi^{2}}$ 15 $0.34\pm 0.94~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.42\pm 0.90~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.50\pm 0.11$ $0.398\pm 0.084$ - - longC41K412.dat
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 299 $1873.8\pm 1.0$ $1870.18\pm 0.99$ $1870.6\pm 0.97$ $1871.4\pm 2.2$ - -
+Table S2: Periods in first half data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original weighted data ($n=78,\Delta T=77$: Chalf.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{\chi}$ (-) $F_{\chi}$ (-) $F_{\chi}$(-) Control file
+$\chi^{2}$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $19.1\pm 1.2$ - - - $\leftarrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$$_{1,1,1,\chi^{2}}$ 5 $0.44\pm 0.12$ - - - 5.62 5.95 5.78 halfC14K111.dat
+168 $1873.6\pm 2.0$ - - - 0.0016 $5.0\times 10^{-5}$ $7.5\times 10^{-6}$
+Two signals
+$\mathcal{M}$=2 $12.82\pm 0.54$ $18.78\pm 0.65$ - - - $\leftarrow$ $\uparrow$
+$\mathcal{M}$$_{2,1,1,\chi^{2}}$ 8 $0.44\pm 0.12$ $0.53\pm 0.13$ - - - 5.3 4.91 halfC14K211.dat
+135 $1870.4\pm 1.4$ $1874.3\pm 1.9$ - - - 0.0025 0.00034
+Three signals
+$\mathcal{M}$=3 $5.580\pm 0.085$ $12.82\pm 0.40$ $19.32\pm 0.83$ - - - $\leftarrow$
+$\mathcal{M}$$_{3,1,1,\chi^{2}}$ 11 $0.42\pm 0.14$ $0.488\pm 0.087$ $0.499\pm 0.096$ - - - 3.88 halfC14K311.dat
+109 $1875.01\pm 0.51$ $1870.2\pm 1.4$ $1873.5\pm 2.3$ - - - 0.012
+Four signals
+$\mathcal{M}$=4 $3.437\pm 0.052~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $5.70\pm 0.13{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $71\pm x0.17 21.2±1.$$2\pm 1.2 - - -$
+$\mathcal{M}$$_{4,1,1,\chi^{2}}$ 14 $0.30\pm 0.12~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.40\pm 0.31~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $6\pm 0.13 0.440$$40\pm 0.096$ - - - halfC14K411.dat
+92 $1872.51\pm 0.66$ $1875.6\pm 1.2$ $1870.8\pm 1.1$ $1870.9\pm 2.9$ - - -
+Table S3: Periods in all weighted data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original weighted data ($n=155,\Delta T=154$: Clong.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{\chi}$ (-) $F_{\chi}$ (-) $F_{\chi}$(-) Control file
+$\chi^{2}$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $12.65\pm 0.20$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$$_{1,1,1,\chi^{2}}$ 5 $0.51\pm 0.14$ - - - 12.4 6.98 7.10 longC14K111.dat
+437 $1871.3\pm 1.3$ - - - $2.7\times 10^{-7}$ $1.6\times 10^{-6}$ $1.9\times 10^{-8}$
+Two signals
+$\mathcal{M}$=2 $5.677\pm 0.089$ $12.64\pm 0.31$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$$_{2,1,1,\chi^{2}}$ 8 $0.393\pm 0.086$ $0.47\pm 0.13$ - - - 6.48 6,53 longC14K211.dat
+384 $1875.65\pm 0.56$ $1871.3\pm 1.2$ - - - 0.00038 $4.1\times 10^{-6}$
+Three signals
+$\mathcal{M}$=3 $5.662\pm 0.077$ $12.78\pm 0.12$ $21.3\pm 1.5$ - - - $\uparrow$
+$\mathcal{M}$$_{3,1,1,\chi^{2}}$ 11 $0.40\pm 0.13$ $0.47\pm 0.10$ $0.37\pm 0.12$ - - - 5.91 longC14K311.dat
+338 $1870.27\pm 0.62$ $1870.36\pm 0.88$ $1870.8\pm 3.3$ - - - 0.00079
+Four signals
+$\mathcal{M}$=4 $5.496\pm 0.093~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $5.66\pm 0.13~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $12.76\pm 0.15$ $21.09\pm 1.5$ - - -
+$\mathcal{M}$$_{4,1,1,\chi^{2}}$ 14 $0.3\pm 1.8~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.4\pm 1.8~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.491\pm 0.13$ $0.39\pm 0.11$ - - - longC14K411.dat
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 300 $1873.8\pm 1.2$ $1870.2\pm 1.1$ $1870.6\pm 0.97$ $1871.3\pm 2.$ - - -
+Table S4: Forecasts (Figures S3 and S4: open circles).
+Figure S3 Figure S4
+$t$ $y$ $\sigma$ $y$ $\sigma$
+(y) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$)
+2025 0.09 0.24 0.35 0.28
+2026 0.35 0.20 0.47 0.29
+2027 0.41 0.21 0.31 0.28
+2028 0.35 0.22 0.21 0.28
+2029 0.40 0.19 0.42 0.34
+2030 0.60 0.20 0.79 0.31
+2031 0.79 0.20 0.94 0.21
+2032 0.75 0.27 0.66 0.21
+2033 0.45 0.34 0.18 0.23
+2034 0.11 0.30 -0.10 0.31
+2035 -0.03 0.21 -0.03 0.38
+2036 0.06 0.22 0.16 0.34
+Table S5: Periods in all non-weighted data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Rlong.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ (-) $A_{2}$ (-) $A_{3}$ (-) $A_{4}$ (-) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $12.71\pm 0.12$ - - - $\leftarrow$ $\leftarrow$ $\leftarrow$
+$\mathcal{M}$1,1,1,R 5 $0.389\pm 0.074$ - - - 3.67 3.77 1.11 longR14K110.dat
+27.1 $1871.19\pm 0.86$ - - - 0.013 0.0016 0.36
+Two signals
+$\mathcal{M}$=2 $5.651\pm 0.030$ $12.69\pm 0.15$ - - - $\leftarrow$ $\leftarrow$
+$\mathcal{M}$2,1,1,R 8 $0.319\pm 0.064$ $0.384\pm 0.088$ - - - 3.77 -0.092 longR14K211.dat
+25.2 $1870.26\pm 0.41$ $1871.3\pm 1.1$ - - - 0.0016 1
+Three signals
+$\mathcal{M}$=3 $5.650\pm 0.032$ $9.12\pm 0.12$ $12.69\pm 0.27$ - - - $\leftarrow$
+$\mathcal{M}$3,1,1,R 11 $0.327\pm 0.084$ $0.299\pm 0.087$ $0.381\pm 0.079$ - - - -3.50 longR14K311.dat
+23.4 $1870.28\pm 0.46$ $1872.5\pm 1.2$ $1871.4\pm 1.3$ - - - 1
+Four signals
+$\mathcal{M}$=4 $5.496\pm 0.093~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $5.662\pm 1.3~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $12.76\pm 0.16$ $21.1\pm 1.5$ -
+$\mathcal{M}$4,1,1,R 14 $0.3\pm 1.9~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.4\pm 1.8~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.491\pm 0.49\pm 0.13$ $0.39\pm 0.11$ - longR14K411.dat
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 25.3 $1873.8\pm 1.2$ $1870.2\pm 1.2$ $1870.6\pm 0.97$ $1871.3\pm 2.1$ -

### v6

@@
-Discrete Chi-Square Method can model and forecast complex time series
+Discrete Chi-Square Method can model and forecast complex time series, like El Nino data between 1870 and 2024
 Abstract
-We show how intensive, large and accurate time series can allow us to see through time.
-Many phenomena have aperiodic and periodic components.
-An ideal time series analysis method would detect such trend and signal(-s) combinations.
-The widely-used Discrete Fourier Transform (DFT) and other frequency-domain parametric time series analysis methods have many application limitations constraining the trend and signal(-s) detection.
-We show that none of those limitations constrains our Discrete Chi-square Method (DCM) which can detect signal(-s) superimposed on an unknown trend.
+Forecasting El Niño is one of the greatest challenges of science.
+We show how intensive, large and accurate time series allow us to see through time.
+Our Discrete Chi-square Method (DCM) can detect arbitrary trend and signal(-s) combinations.
+It can forecast complex time series.
+The widely-used Discrete Fourier Transform (DFT) and other frequency-domain parametric time series analysis methods have many application limitations.
+None of those limitations constrains the DCM.
 Our simulated time series analyses ascertain the revolutionary Window Dimension Effect (WD-effect): “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.” The simulations also expose the DFT’s weaknesses and the DCM’s efficiency.
@@
 We present a computational well-posed solution.
-The DCM can forecast complex time series.
 The best DCM model must be correct if it passes our Forecast-test.
 Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series.
-In our appendix, we show that DCM can model and forecast El Niño.
+In our appendix, we show that the DCM can model and forecast El Nino data between 1870 and 2024.
+An immediate, independent and objective validity check of our analysis may save some money.
 keywords:
-Time series analysis, Non-linear and non-stationary processes, Ill-posed problems, Computational statistics, Forecasting, Discrete Chi-square Method, Discrete Fourier Transform, El Niño
+El Niño, Time series analysis, Non-linear and non-stationary processes, Ill-posed problems, Computational statistics, Forecasting, Discrete Chi-square Method, Discrete Fourier Transform
 1 Introduction
@@
 Finally, we discuss our results (Section 4).
+In our Appendix, we show that the DCM can model and forecast El Niño (Section A1).
 2 Methods
@@
 These functions are the periodic function
-$h(t)$ $=$ $h(t,K_{1},K_{2})=\sum_{i=1}^{K_{1}}h_{i}(t,f_{i})$ (2)
+$h(t)=h(t,K_{1},K_{2})$ $=$ $\begin{cases}0,\text{if }K_{1}=0\\ \sum_{i=1}^{K_{1}}h_{i}(t),\text{if }K_{1}\geq 1\end{cases}$ (2)
 $h_{i}(t,f_{i})$ $=$ $\sum_{j=1}^{K_{2}}B_{i,j}\cos{(2\pi jf_{i}t)}$
@@
 For example, DCM can be forced search for too few or too many $K_{1}$ signals, or the selected $p(t)$ trend order $K_{3}$ can be wrong, as shown in Figures 5-10 by [29].
-Such DCM models are unstable and we denote them with $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”, like in Jetsu [^1] Jetsu L.
-“Do planets cause the sunspot cycle?”, submitted to Scientific Reports.
-August 13th, 2025.
-(2025).
+Such DCM models are unstable and we denote them with $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”, like in [30].
 These unstable models have three signatures
@@
 (g) DFT model detrended $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (black continuous line), DFT pure sine $s_{\mathrm{DFT}}(t)$ (red thick continuous line), detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and DFT model residuals (blue dots) offset to -1.5 level (blue dotted line).
-The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [29, TSPA, Equation 1] and the Continuous Period Search [36, CPS, Equation 3].
+The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [31, TSPA, Equation 1] and the Continuous Period Search [39, CPS, Equation 3].
 TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$.
@@
 We search for signals in these simulated time series by using the frequently applied DFT version formulated by [24] [^1] This paper had over 2600 citations in December 2025.
-Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [23, Equation 1].
+Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [24, Equation 1].
 The notations for the DFT model are
@@
 Our DFT analyses of simulated time series may fail due to AL6-AL10.
-The time series is “too short” [30, AL6]
+The time series is “too short” [32, AL6]
 $P>\Delta T.$ (35)
-The time series is non-stationary because it contains “a trend” [42, AL7]
+The time series is non-stationary because it contains “a trend” [45, AL7]
 $p_{\mathrm{DFT}}(t)\neq 0.$ (36)
-Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [30, AL9].
+Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [32, AL9].
 $|f_{1}-f_{2}|<f_{0}={\Delta T}^{-1}.$ (37)
@@
 For each sample, this table specifies the electronic data file and the electronic DCM control file.
-[^3] We publish all our data files and all DCM control files in https://zenodo.org/uploads/17018676 The detected $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$ and $M_{0}$ values are correct and accurate even for the combination $n=50$ and ${\mathrm{SN}}=50$.
+[^2] We publish all our data files and all DCM control files in
+https://zenodo.org/records/18726181/
+The detected $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$ and $M_{0}$ values are correct and accurate even for the combination $n=50$ and ${\mathrm{SN}}=50$.
 Regardless of $\Delta T<P_{1}$, these model parameter values become more accurate and converge to the correct simulated values when $n$ and $\mathrm{SN}$} increase.
@@
 The DFT detects the wrong period $P_{1}=1.163$ for the original data (Figure 3e: diamond).
-This is an expected result because the detected period should be close to $(P_{1}+P_{2})/2$ when the peak to peak amplitudes of the simulated data, $A_{1}=A_{2}$, are equal (Jetsu 2025).
+This is an expected result because the detected period should be close to $(P_{1}+P_{2})/2$ when the peak to peak amplitudes of the simulated data, $A_{1}=A_{2}$, are equal [30].
 The two DFT periodogram $z_{\mathrm{DFT}}(f)$ peaks at frequencies $1/P_{1}$ and $1/P_{2}$ overlap and merge into one peak.
@@
 The DCM residuals (blue dots) offset to the level of -3.0 show no trends and are extremely stable.
-Since the peak to peak amplitudes of the simulated data signals are equal, $A_{1}=A_{2}$, the expected result for the DFT analysis of original data is $(P_{1}+P_{2})/2$, where $P_{1}$ and $P_{2}$ are the simulated signal periods (Jetsu 2025).
+Since the peak to peak amplitudes of the simulated data signals are equal, $A_{1}=A_{2}$, the expected result for the DFT analysis of original data is $(P_{1}+P_{2})/2$, where $P_{1}$ and $P_{2}$ are the simulated signal periods [30].
 The DFT detects this expected wrong period $P_{1}=1.165$ for the original data (Figure 4e: diamond).
@@
 The DCM detects the correct signal(s) and trend, but DFT does not.
-For the sunspot data, the DCM detects signals superimposed on a constant trend (Jetsu 2025).
+For the sunspot data, the DCM detects signals superimposed on a constant trend [30].
 The simulations presented here support the conclusion that those solar signals and the trend are correct.
@@
 (1) Simple model: Model number $\mathcal{M}$, model $g_{K_{1},K_{2},K_{3}}$, number of free parameters $\eta$, chi-square $\chi^{2}$ and control file name.
-(2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.33footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
+(2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.22footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
 […]
@@
 We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large.
-One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu (2025, Table S7).
+One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu [30, Table S7].
 We conclude that the best model for the real time series can be found by applying the Fisher-test to any arbitrary number of different nested DCM or DFT models.
 3.10 Significance estimates
-Jetsu [27] or Jetsu (2025) gave no signal significance estimate for the first detected period.
+Jetsu [29] or Jetsu [30] gave no signal significance estimate for the first detected period.
 Here, we use the Fisher-test for this purpose.
@@
 In other words, the signal significances decrease.
-Typical examples can be found in Jetsu 2025 (Tables S5-S16).
+Typical examples can be found in Jetsu [30, Tables S5-S1].
 No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected.
 For the Model 1 simulated $n=100$ and ${\mathrm{SN}}=100$ combination, Table 8 shows that this time series contains only one signal.
-For this Model 1 time series, there is no need to discuss the DFT significance estimates [23, their Equation 22] because this method fails to detect the correct period.
+For this Model 1 time series, there is no need to discuss the DFT significance estimates [24, their Equation 22] because this method fails to detect the correct period.
 We conclude that the Fisher-test identifies the correct DCM model (Section 3.9), as well as gives the signal significance $Q_{\mathrm{F}}$ estimates (Section 3.10).
@@
 Otherwise as in Table 8.
-44footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.
+33footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.
 (1) (2)
@@
 The artificial bootstrap data sets (Equation 31) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation).
-We routinely check the stability of these solutions (e.g.
-Jetsu 2025, Figure S5).
+We routinely check the stability of these solutions [30, Figure S5].
 The unstable models, where the model parameter changes are large, are rejected (Section 2.1: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}” models).
@@
 3.12 Forecast
-There are numerous techniques for forecasting a time series [e.g., 22, 26, 31].
+There are numerous techniques for forecasting a time series [e.g., 23, 28, 33].
 The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast.
@@
   • If this new signal is unreal, the $z_{\mathrm{Fore}}$ value of forecasted data increases.
-This “Forecast-test” technique revealed at least five real signals in the sunspot record (Jetsu 2025).
+This “Forecast-test” technique revealed at least five real signals in the sunspot record [30]
 We compute the $z_{\mathrm{Fore}}$ parameter value from the known forecasted data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ values (Equation 3.12).
@@
 during the chosen sample window $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$.
-Jetsu (2025) used this $m_{\mathrm{Fore}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720.
-Since the known mean level of all sunspot data was $m$, Jetsu (2025) used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
+Jetsu [30] used this $m_{\mathrm{Fore}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720.
+Since the known mean level of all sunspot data was $m$, Jetsu [30] used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
   • If DCM detects a real new signal in all data, the $m_{\mathrm{Fore}}$ value decreases.
@@
 This statistical effect is the same when DCM is applied to the time series of any arbitrary phenomenon.
-This effect would, for example, explain why our solar cycle amplitude forecasts are less accurate than our solar cycle minimum and maximum epoch forecasts (Jetsu 2025).
+This effect would, for example, explain why our solar cycle amplitude forecasts are less accurate than our solar cycle minimum and maximum epoch forecasts [30].
 DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct.
@@
 Long computation time is the main AL constraining DCM.
-For example, Jetsu (2025) computed about 4.5 million LS fits to search for four signals in the monthly sunspot data $(n=3287,K_{1}=4,K_{2}=1,K_{3}=0,n_{\mathrm{L}}=100,n_{\mathrm{S}}=30,n_{\mathrm{B}}=20)$.
+For example, Jetsu [30] computed about 4.5 million LS fits to search for four signals in the monthly sunspot data $(n=3287,K_{1}=4,K_{2}=1,K_{3}=0,n_{\mathrm{L}}=100,n_{\mathrm{S}}=30,n_{\mathrm{B}}=20)$.
 This required several months of CPU.
@@
 Perhaps our DCM should be re-named the Slow Gauß Transform (SGT).
-5 Appendix
-In this appendix, we show that our DCM can model and forecast the El Niño phenomenon between the years 1870 and 2024.
-This result may be the “holy grail” of climatology.
-The solar forcing most probably causes the detected $5.850\pm 0.085$, $12.82\pm 0.40$ and $19.30\pm 0.83$ year signals.
-These strict periodicities indicate that the solar cycle is caused by the planets.
-Our results confirm that the DCM is an ideal time series analysis method for detecting an unknown number of signals superimposed on an unkown trend.
-The El Niño phenomenon is considered unpredictable [52, 39, 6, 25, 32, 35, 50, 40].
-Our analysed data are the yearly means in the The Seasonal Mean Nino 4 HadISST1.1 (NOAA PSL) sample.
-[^5] The data from https://psl.noaa.gov/data/timeseries/month/data/nino4.long.anom.data were retrieved on January 29th, 2026.
-There were no values for the last four months of 2025.
-The full data contains $n=155$ yearly means between 1870 and 2024.
-5.1 Trend
+A1 Appendix
+Introduction
+The modelling and forecasting of El Niño is considered notoriously difficult [55, 42, 6, 26, 35, 38, 53, 43].
+We will show that our DCM can easily model and forecast the El Niño time series.
+Since the challenging analyses of all simulated extremely complex time series succeed (Sections 3.1-3.7), there are no rational reasons why our analysis of the simple El Niño time series would fail.
+Data
+Our El Niño data are from The Seasonal Mean Nino 4 HadISST1.1 (NOAA PSL) sample. These NOAA data were retrieved on January 29th, 2026. The twelve monthly sea surface temperature anomaly values give
+  • $t_{i}=$ year
+  • $y_{i}=$ mean of monthly values
+  • $\sigma_{i}=$ standard deviation of monthly values
+There were no values for the last four months of 2025. The full data contains $n=155$ yearly values between 1870 and 2024.
+Results
+The DCM can model and forecast the El Niño time series. This result may be the “holy grail” of climatology.
+Trend
 We detect significant signals only from the weighted data.
 Therefore, the trend and the forecasts are solved from the weighted data.
-The weighted DCM analysis results for all data are given in Table S1.
-For one, two, three and four signals, the best trend alternative is the linear trend $K_{1}=1$.
-We use this $K_{1}=1$ trend in our DCM analysis of El Niño data.
-For the $\mathcal{M}$=3 model, the positive linear trend is $2M_{1}=0.56\pm 0.20$ $\mathrm{C}^{o}$during $\Delta T=154$ years (Equation 5), like in all $\mathcal{M}$=1-4 models.
-5.2 Forecasting data
+The weighted DCM analysis results for all data are given in Table A1.
+For one, two, three, and four signals, the best trend alternative is the linear trend $K_{3}=1$.
+We use this $K_{3}=1$ trend in our DCM analysis of El Niño data.
+For the $\mathcal{M}$=3 model, the positive linear trend is $2M_{1}=0.56\pm 0.20$ $\mathrm{C}^{o}$ during $\Delta T=154$ years (Equation 5).
+This trend is approximately the same in all $\mathcal{M}$=1-4 models.
+Forecasting data
 We use the first half of all data $(n=78)$ to forecast the second half $(n=77)$.
-The results of weighted DCM analysis are given in Table S2.
-The periods of the three strongest signals are $5.850\pm 0.085$, $12.82\pm 0.40$ and $19.30\pm 0.83$ years (Model $\mathcal{M}$=3).
-The forecasts of the two and three signals models, open circles, are shown in Figures S1 and S2.
-For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signals models are 421, 336, 370 and 428, respectively.
+The results of weighted DCM analysis are given in Table A2.
+The periods of the three strongest signals are $P_{1}=5.580\pm 0.085$, $P_{2}=12.82\pm 0.40$ and $P_{3}=19.30\pm 0.83$ years (Model $\mathcal{M}$=3).
+The forecasts of the two and three signals models, open circles, are shown in Figures A1 and A2.
+For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signal models are 421, 336, 370 and 428, respectively.
 Hence, the two signal $\mathcal{M}$=2 model gives the best forecast.
-5.3 All data
-The DCM analysis results for all weighted data are given in Table S3.
-The three signals $\mathcal{M}$=3 model is the best.
-The four signal model is unstable $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}})$.
-The three and four signal models are shown in Figures S3 and S4.
-For both alternatives, the predicted values, open circles, after 2024 are given in Table S4.
-Both alternatives give essentially the same forecast.
-For the sake of consistency, we also give the results of non-weighted analysis of all data (Table S5).
-We detect the same signals as from non-weighted data, but the signal significances $Q_{\mathrm{F}}$ are lower.
-This confirms that the data error information is crucial.
-5.4 Main result
-Our DCM can model and forecast the El Niño phenomenon.
-The solar forcing is the only possible explanation for these results.
-Within their error limits, the detected three El Niño signals could represent the Hale cycle [24], the solar cycle[46, 56] and half the solar cycle.
-These strict El Niño periodicities indicate that the planets could cause a stationary and deterministic solar cycle [28].
-A stochastic and non-stationary solar cycle [8, 54] could not cause the detected regular El Niño periodicities.
+All data
+The DCM analysis results for all weighted data are given in Table A3.
+We detect the same periods that are already detected from the first half of all data.
+The three signal $\mathcal{M}$=3 model is the best.
+The El Niño signal periods are $P_{1}=5.662\pm 0.077$, $P_{2}=12.78\pm 0.12$ and $P_{3}=21.3\pm 1.5$ years.
+The DCM periodogram of the $\mathcal{M}$=3 model shows three clearly defined minima (Figure A3a).
+These three signals have the peak to peak amplitudes $A_{1}=0.40\pm 0.13$ $\mathrm{C}^{o}$, $A_{2}=0.47\pm 0.10$ $\mathrm{C}^{o}$ and $A_{3}=0.37\pm 0.12$ $\mathrm{C}^{o}$ (Figure A3b).
+The four signal $\mathcal{M}$=4 model is unstable $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}})$.
+The three and four signal models are shown in Figures A4 and A5.
+For both alternatives, the predicted values after 2024, open circles, are given in Table A4.
+Both model alternatives give essentially the same forecast.
+For the sake of consistency, we also give the results of non-weighted analysis of all data (Table A5).
+We detect the same signals from these non-weighted data, but the signal significances $Q_{\mathrm{F}}$ are lower.
+This confirms that the data error information of weighted data is crucial.
+For all weighted data, the detection of the first signal is absolutely certain.
+The Fisher-test comparison between the $K_{3}=0$ pure noise model $g_{0,1,0,\chi^{2}}$ and the one signal model $g_{1,1,1,\chi^{2}}$ gives signal significance $Q_{\mathrm{F}}=3.9\times 10^{-13}$ (Table A6, $\mathcal{M}$=1).
+All other polynomial model alternatives, $1\leq K_{3}\leq 7$, are also rejected.
+The second and third signal significances are $Q_{\mathrm{F}}=2.7\times 10^{-7}$ and $Q_{\mathrm{F}}=0.00038$ (Table A3: $\mathcal{M}$=2-3).
+Holy grail: El Niño, Sun and Planets
+The DCM can model and forecast the El Niño phenomenon.
+The first half data forecast fits to the second half data (Figures A1b-d and A2b-d: open circles).
+For the two times larger all data sample, time will tell if our next decade forecast is also correct (Figures A4d and A5d: open circles).
+For example, we forecast an extreme El Niño between 2030 and 2032.
+The solar forcing is the most probable explanation for these results.
+The three detected El Niño signal periods are close to the 22 years Hale cycle period [25], the 11 years solar cycle period [49, 59] and half the solar cycle period.
+A stochastic and non-stationary solar cycle [8, 57] can not cause these strict El Niño periodicities.
+Only very regular solar energy output changes can cause these El Niño signals because the noise added by the terrestrial climate response can not prevent our signal detection from the sea surface temperature data.
+These strictly periodic El Niño signals support the claim that the planets cause a stationary, multi-periodic and deterministic solar cycle(-s) [30].
+Our mathematical model can detect regularities that the current physical models fail to detect, let alone forecast.
+Mathematics detects signals that should not be there according to Physics.
+The counterargument is that simple mathematics can not explain a complex physical phenomenon.
+We argue that the El Niño physical model is simple.
+The solar energy output determines El Niño.
+It is no surprise that the Sun warms the largest battery on the Earth, the Pacific Ocean.
+And if many cycles moderate the solar energy output, it is no surprise either that we have failed to discover the interference of these cycles, El Niño, in our yearly climate temperatures.
+The current El Niño ”tool kit” is admittedly impressive [6].
+Generations of hard work has given us those tools.
+Masses of scientists develop and use them.
+Yet, even the best currently available physical El Niño forecasts fail in 1.5 years [62].
+Our mathematical DCM forecast can connect seven El Niño decades to the next seven El Niño decades (Figures A1 and A2).
+That is quite a coincidence if the chaos of physical models takes over in 1.5 years.
+The annual cost of El Niño damage is roughly one trillion dollars [9, 34].
+An accurate El Niño prediction for just one year ahead would save the global economy trillions by mitigating long-term productivity losses [27, 61].
+This should motivate other scientists to check our results immediately.
+[^4] All necessary file are in
+https://zenodo.org/records/18726181/
+No supercomputers are needed because an ordinary PC can detect the first two signals in about one minute. The third signal detection takes a few minutes.
+William of Ockham (1287–1347) stated that if two competing hypotheses make the same predictions, the simplest hypothesis is usually the best one (Occam’s razor).
+The connection between El Niño, the Sun and the Planets may seem like the holy grail.
+However, those three are just physical phenomena.
+The real scientific holy grail is the DCM.
 Acknowledgements
 This work has made use of NASA’s Astrophysics Data System (ADS) services.
-Author contributions statement
-L.J. peformed this research and wrote the manuscript.
+Author contributions
+L.J. did this research and wrote the manuscript.
 Consent for publication
-The author grants permission for publication.
+The author grants permission for publication in …
 Additional information
@@
 Data availability
-Main manuscript analysis: All data files and all DCM control files are stored to
-https://zenodo.org/uploads/17018676.
-Appendix analysis: All files, DCM Python code, DCM control files and Python code are stored to the
-https://zenodo.org/records/18595333.
+All data files and Python codes are stored to
+https://zenodo.org/records/18726181/
+The manual for repeating all analysis is also published there.
 Declarations
@@
 Ethical approval and consent to participate Not applicable.
-Supplementary material
-The $\mathcal{M}$= 1 analysis in Table S1 can be repeated using these two commands
-cp longC14K110.dat dcm.dat
-python dcm.py
-The whole El Niño analysis can be repeated in similar fashion.
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S1: First half forecast (Table S2: $\mathcal{M}$=2): Forecasting data (closed black circles) and forecasted data (open black circles).
+Figure A1: First half forecast (Table A2: $\mathcal{M}$=2): Forecasting data (closed black circles) and forecasted data (open black circles).
 Green and red dottted lines denote one and three sigma model $g(t)$ errors.
 Residuals (closed and open blue circles) are offset to level of -3.
+Units are x-axis: $t_{i}]$=y, y-axis: $y_{i}]=$ $\mathrm{C}^{o}$ and error bars = $1\times\sigma_{i}]=$ $\mathrm{C}^{o}$.
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S2: First half forecast (Table S2: $\mathcal{M}$=3). Otherwise as in Figure S1.
+Figure A2: First half forecast (Table A2: $\mathcal{M}$=3). Otherwise as in Figure A1.
+Figure A3: All data model $\mathcal{M}$=3 in Table A3.
+(a) Periodogram (Equation 21).
+Otherwise as in Figures 2ab.
+(b) Three $h_{1}(t)$, $h_{2}(t)$ and $h_{3}(t)$ signals as a function of time $t$ (left-hand panels) and phase $\phi$ (right-hand panels) [30, Equation 29].
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S3: All data model $\mathcal{M}$=3 in Table S3. Otherwise as in Figure S1.
+Figure A4: All data model $\mathcal{M}$=3 in Table A3. Otherwise as in Figure A1.
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S4: All data model $\mathcal{M}$=4 in Table S3. Otherwise as in Figure S1.
-Table S1: Trend in all data.
-Weighted DCM analysis performed between $P_{\mathrm{min}}=3$ and $P_{\mathrm{max}}=100$.
-(1) $\mathcal{M}$model, (2) $\eta$ free parameters and $\chi^{2}$ of model, (3) First signal: $P_{1}$ period, $A_{1}$ peak to peak amplitude and $t_{\mathrm{min},1}$ primary minimum epoch.
-(4-6) Next signals.
-(7-8) Fisher-test: $F_{\chi}$ test statistic and $Q_{\mathrm{F}}$ critical level.
-(9) Control file.
-Arrows point to best model in Fisher-test comparison.
+Figure A5: All data model $\mathcal{M}$=4 in Table A3. Otherwise as in Figure A1.
+Table A1: Trend in all weighted data. DCM analysis between $P_{\mathrm{min}}=3$ and $P_{\mathrm{max}}=100$ years. Otherwise as in Table 8
 Period analysis Fisher-test
@@
 $\mathcal{M}$=1 $12.67\pm 0.33$ - - - $\uparrow$ $\uparrow$
-$\mathcal{M}$$_{1,1,0,\chi^{2}}$ 4 $0.62\pm 0.15$ - - - 26.9 14.0 longC14K111.dat
+$\mathcal{M}$$_{1,1,0,\chi^{2}}$ 4 $0.62\pm 0.15$ - - - 26.9 14.0 longC14K110.dat
 516 $1871.2\pm 1.3$ - - - $6.7\times 10^{7}$ $2.7\times 10^{-6}$
@@
 $\mathcal{M}$=3 $12.65\pm 0.43$ - - -
-$\mathcal{M}$$_{1,1,2,\chi^{2}}$ 6 $0.52\pm 0.12$ - - - longC41K112.dat
+$\mathcal{M}$$_{1,1,2,\chi^{2}}$ 6 $0.52\pm 0.12$ - - - longC14K112.dat
 434 $1871.4\pm 1.1$ - - -
@@
 $\mathcal{M}$=6 $5.677\pm 0.077$ $12.64\pm 0.42$ - - -
-$\mathcal{M}$$_{2,1,2,\chi^{2}}$ 9 $0.388\pm 0.093$ $0.47\pm 0.12$ - - - longC41K212.dat
+$\mathcal{M}$$_{2,1,2,\chi^{2}}$ 9 $0.388\pm 0.093$ $0.47\pm 0.12$ - - - longC14K212.dat
 384 $1875.66\pm 0.76$ $1871.4\pm 1.2$ - - -
@@
 $\mathcal{M}$=9 $5.661\pm 0.073$ $76\pm 0.22 21.2±1.$$2\pm 1.2 - - -$
-$\mathcal{M}$$_{3,1,2,\chi^{2}}$ 12 $0.40\pm 0.11$ $0.49\pm 0.10$ $0.39\pm 0.12$ - - - longC41K312.dat
+$\mathcal{M}$$_{3,1,2,\chi^{2}}$ 12 $0.40\pm 0.11$ $0.49\pm 0.10$ $0.39\pm 0.12$ - - - longC1¤K312.dat
 336 $1870.3\pm 0.74$ $1870.6\pm 1.2$ $1871.1\pm 2.8$ - - -
@@
 $\mathcal{M}$=12 $5.497\pm 0.087~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $0.566\pm 0.062~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{IF}}}$ $12.75\pm 0.17$ $21.0\pm 1.0$ - -
-$\mathcal{M}$$_{4,1,2,\chi^{2}}$ 15 $0.34\pm 0.94~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.42\pm 0.90~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.50\pm 0.11$ $0.398\pm 0.084$ - - longC41K412.dat
+$\mathcal{M}$$_{4,1,2,\chi^{2}}$ 15 $0.34\pm 0.94~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.42\pm 0.90~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.50\pm 0.11$ $0.398\pm 0.084$ - - longC14K412.dat
 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 299 $1873.8\pm 1.0$ $1870.18\pm 0.99$ $1870.6\pm 0.97$ $1871.4\pm 2.2$ - -
-Table S2: Periods in first half data. Otherwise as in Table S1.
+Table A2: Periods in first half data. Otherwise as in Table A1.
 Period analysis Fisher-test
@@
 92 $1872.51\pm 0.66$ $1875.6\pm 1.2$ $1870.8\pm 1.1$ $1870.9\pm 2.9$ - - -
-Table S3: Periods in all weighted data. Otherwise as in Table S1.
+Table A3: Periods in all weighted data. Otherwise as in Table A1.
 Period analysis Fisher-test
@@
 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 300 $1873.8\pm 1.2$ $1870.2\pm 1.1$ $1870.6\pm 0.97$ $1871.3\pm 2.$ - - -
-Table S4: Forecasts (Figures S3 and S4: open circles).
-Figure S3 Figure S4
-$t$ $y$ $\sigma$ $y$ $\sigma$
+Table A4: Forecasts (Figures A4 and A5: open circles).
+Figure A4 Figure A5
+$t$ $y$ $\pm 1\sigma$ $y$ $\pm 1\sigma$
 (y) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$)
@@
 2036 0.06 0.22 0.16 0.34
-Table S5: Periods in all non-weighted data. Otherwise as in Table S1.
+Table A5: Periods in all non-weighted data. Otherwise as in Table A1.
 Period analysis Fisher-test
@@
 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 25.3 $1873.8\pm 1.2$ $1870.2\pm 1.2$ $1870.6\pm 0.97$ $1871.3\pm 2.1$ -
+Table A6: Significance one signal model $\mathcal{M}$=1 in Table A3. Otherwise as in Table 9. 44footnotetext: Note: No test for $\mathcal{M}$=5  $(\eta_{1}=\eta_{2}=5)$.
+(1) (2)
+$g_{1,1,1}$
+Polynomial $\eta=5,\chi^{2}=437$
+$\mathcal{M}$=1 $\uparrow$
+$g(t)=p(t,K_{3}=0)$ $F=19.9$
+$\eta=1$, $\chi^{2}=670$ $Q_{\mathrm{F}}=3.9\times 10^{-13}$
+$\mathcal{M}$=2 $\uparrow$
+$g(t)=p(t,K_{3}=1)$ $11.4$
+$\eta=2$, $\chi^{2}=537$ $Q_{\mathrm{F}}=9.2\times 10^{-7}$
+$\mathcal{M}$=3 $\uparrow$
+$g(t)=p(t,K_{3}=2)$ $F=17.0$
+$\eta=3$, $\chi^{2}=537$ $Q_{\mathrm{F}}=2.1\times 10^{-7}$
+$\mathcal{M}$=4 $\uparrow$
+$g(t)=p(t,K_{3}=3)$ $F=26.5$
+$\eta=4$, $\chi^{2}=515$ $Q_{\mathrm{F}}=7.8\times 10^{-7}$
+$\mathcal{M}$=5 $\uparrow$
+$g(t)=p(t,K_{3}=4)$ No test11footnotemark: 1
+$\eta=5$, $\chi^{2}=514$ $Q_{\mathrm{F}}=1$
+$\mathcal{M}$=6 $\uparrow$
+$g(t)=p(t,K_{3}=5)$ F=22F=-22
+$\eta=6$, $\chi^{2}=514$ $Q_{\mathrm{F}}=1$
+$\mathcal{M}$=7 $\uparrow$
+$g(t)=p(t,K_{3}=6)$ F=8.2F=-8.2
+$\eta=7$, $\chi^{2}=492$ $Q_{\mathrm{F}}=1$
+$\mathcal{M}$=8 $\uparrow$
+$g(t)=p(t,K_{3}=7)$ F=5.4F=-5.4
+$\eta=8$, $\chi^{2}=492$ $Q_{\mathrm{F}}=1$
+[…]

### v7

@@
 The modelling and forecasting of El Niño is considered notoriously difficult [55, 42, 6, 26, 35, 38, 53, 43].
-We will show that our DCM can easily model and forecast the El Niño time series.
+We will show that our DCM can model and forecast the “Big Wave” in El Niño time series.
 Since the challenging analyses of all simulated extremely complex time series succeed (Sections 3.1-3.7), there are no rational reasons why our analysis of the simple El Niño time series would fail.
 Data
-Our El Niño data are from The Seasonal Mean Nino 4 HadISST1.1 (NOAA PSL) sample. These NOAA data were retrieved on January 29th, 2026. The twelve monthly sea surface temperature anomaly values give
+Our El Niño data are from The Seasonal Mean Nino 4 HadISST1.1 (NOAA PSL) sample. These NOAA1444https://psl.noaa.gov/data/timeseries/month/
+data/nino4.long.anom.dat
+data were retrieved on January 29th, 2026. The twelve monthly sea surface temperature (SST) anomaly values give the yearly values
   • $t_{i}=$ year
@@
   • $\sigma_{i}=$ standard deviation of monthly values
-There were no values for the last four months of 2025. The full data contains $n=155$ yearly values between 1870 and 2024.
+Four monthly values for the year 2025 were missing. Therefore, we analyse $n=155$ yearly values between 1870 and 2024.
 Results
-The DCM can model and forecast the El Niño time series. This result may be the “holy grail” of climatology.
+The DCM can model and forecast the “Big Wave” in El Niño time series. This result may be the “holy grail” of climatology.
 Trend
-We detect significant signals only from the weighted data.
+We detect significant signals only from the weighted yearly data.
 Therefore, the trend and the forecasts are solved from the weighted data.
-The weighted DCM analysis results for all data are given in Table A1.
-For one, two, three, and four signals, the best trend alternative is the linear trend $K_{3}=1$.
+The weighted DCM analysis results for all yearly data are given in Table S1.
+For one, two, three, and four signals, the best “Big Wave” alternative is the linear trend $K_{3}=1$.
 We use this $K_{3}=1$ trend in our DCM analysis of El Niño data.
@@
 This trend is approximately the same in all $\mathcal{M}$=1-4 models.
+This pattern supports climate change-induced global warming.
 Forecasting data
-We use the first half of all data $(n=78)$ to forecast the second half $(n=77)$.
-The results of weighted DCM analysis are given in Table A2.
+We use the first half of all yearly data $(n=78)$ to forecast the second half $(n=77)$.
+The results of weighted DCM analysis are given in Table S2.
 The periods of the three strongest signals are $P_{1}=5.580\pm 0.085$, $P_{2}=12.82\pm 0.40$ and $P_{3}=19.30\pm 0.83$ years (Model $\mathcal{M}$=3).
-The forecasts of the two and three signals models, open circles, are shown in Figures A1 and A2.
+The forecasts of the two and three signals models, open circles, are shown in Figures S1 and S2.
 For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signal models are 421, 336, 370 and 428, respectively.
-Hence, the two signal $\mathcal{M}$=2 model gives the best forecast.
+Hence, the two signal $\mathcal{M}$=2 model gives the best “Big Wave” forecast.
 All data
-The DCM analysis results for all weighted data are given in Table A3.
-We detect the same periods that are already detected from the first half of all data.
+The DCM analysis results for all weighted yearly data are given in Table S3.
+We detect the same “Big Wave” periods that are already detected from the first half of all data.
 The three signal $\mathcal{M}$=3 model is the best.
-The El Niño signal periods are $P_{1}=5.662\pm 0.077$, $P_{2}=12.78\pm 0.12$ and $P_{3}=21.3\pm 1.5$ years.
-The DCM periodogram of the $\mathcal{M}$=3 model shows three clearly defined minima (Figure A3a).
-These three signals have the peak to peak amplitudes $A_{1}=0.40\pm 0.13$ $\mathrm{C}^{o}$, $A_{2}=0.47\pm 0.10$ $\mathrm{C}^{o}$ and $A_{3}=0.37\pm 0.12$ $\mathrm{C}^{o}$ (Figure A3b).
+The El Niño  “Big Wave” signal periods are $P_{1}=5.662\pm 0.077$, $P_{2}=12.78\pm 0.12$ and $P_{3}=21.3\pm 1.5$ years.
+The DCM periodogram of the $\mathcal{M}$=3 model shows three clearly defined minima (Figure S3a).
+These three signals have the peak to peak amplitudes $A_{1}=0.40\pm 0.13$ $\mathrm{C}^{o}$, $A_{2}=0.47\pm 0.10$ $\mathrm{C}^{o}$ and $A_{3}=0.37\pm 0.12$ $\mathrm{C}^{o}$.
 The four signal $\mathcal{M}$=4 model is unstable $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}})$.
-The three and four signal models are shown in Figures A4 and A5.
-For both alternatives, the predicted values after 2024, open circles, are given in Table A4.
-Both model alternatives give essentially the same forecast.
-For the sake of consistency, we also give the results of non-weighted analysis of all data (Table A5).
+The three and four signal models are shown in Figures S3 and S4.
+For both alternatives, the predicted values after 2024, open circles, are given in Table S4.
+The $\mathcal{M}$=3 and $\mathcal{M}$=4 model alternatives give essentially the same “Big Wave” forecast between 2025 and 2036.
+We might as well have forecasted the “Big Wave” for the next century because our first half data forecasts for the second half data are quite accurate (Figures S1 and S2).
+For the sake of consistency, we also give the “Big Wave” results of non-weighted analysis of all data (Table S5).
 We detect the same signals from these non-weighted data, but the signal significances $Q_{\mathrm{F}}$ are lower.
 This confirms that the data error information of weighted data is crucial.
-For all weighted data, the detection of the first signal is absolutely certain.
-The Fisher-test comparison between the $K_{3}=0$ pure noise model $g_{0,1,0,\chi^{2}}$ and the one signal model $g_{1,1,1,\chi^{2}}$ gives signal significance $Q_{\mathrm{F}}=3.9\times 10^{-13}$ (Table A6, $\mathcal{M}$=1).
+For all weighted data, the detection of the first “Big Wave” signal is absolutely certain.
+The Fisher-test comparison between the $K_{3}=0$ pure noise model $g_{0,1,0,\chi^{2}}$ and the one signal model $g_{1,1,1,\chi^{2}}$ gives signal significance $Q_{\mathrm{F}}=3.9\times 10^{-13}$ (Table S6, $\mathcal{M}$=1).
 All other polynomial model alternatives, $1\leq K_{3}\leq 7$, are also rejected.
-The second and third signal significances are $Q_{\mathrm{F}}=2.7\times 10^{-7}$ and $Q_{\mathrm{F}}=0.00038$ (Table A3: $\mathcal{M}$=2-3).
+The second and third “Big Wave” signal significances are $Q_{\mathrm{F}}=2.7\times 10^{-7}$ and $Q_{\mathrm{F}}=0.00038$ (Table S3: $\mathcal{M}$=2-3).
 Holy grail: El Niño, Sun and Planets
-The DCM can model and forecast the El Niño phenomenon.
-The first half data forecast fits to the second half data (Figures A1b-d and A2b-d: open circles).
-For the two times larger all data sample, time will tell if our next decade forecast is also correct (Figures A4d and A5d: open circles).
+The DCM can model and forecast the “Big Wave” of El Niño.
+The first half data forecast fits to the second half data (Figures S1b-d and S2b-d: open circles).
+For the two times larger all data sample, time will tell if our next decade “Big Wave” forecast is also correct (Figures S3d and S4d: open circles).
 For example, we forecast an extreme El Niño between 2030 and 2032.
-The solar forcing is the most probable explanation for these results.
+The solar forcing is the most probable explanation for this El Niño  “Big Wave” result.
 The three detected El Niño signal periods are close to the 22 years Hale cycle period [25], the 11 years solar cycle period [49, 59] and half the solar cycle period.
-A stochastic and non-stationary solar cycle [8, 57] can not cause these strict El Niño periodicities.
-Only very regular solar energy output changes can cause these El Niño signals because the noise added by the terrestrial climate response can not prevent our signal detection from the sea surface temperature data.
-These strictly periodic El Niño signals support the claim that the planets cause a stationary, multi-periodic and deterministic solar cycle(-s) [30].
+A stochastic and non-stationary solar cycle [8, 57] can not cause these strict “Big Wave” periodicities.
+Only very regular solar energy output changes can cause these signals because the noise added by the terrestrial climate response can not prevent our signal detection from the sea surface temperature data.
+These strictly periodic El Niño  “Big Wave” signals support the claim that the planets cause a stationary, multi-periodic and deterministic solar cycle(-s) [30].
 Our mathematical model can detect regularities that the current physical models fail to detect, let alone forecast.
@@
 The counterargument is that simple mathematics can not explain a complex physical phenomenon.
-We argue that the El Niño physical model is simple.
-The solar energy output determines El Niño.
+We argue that the El Niño physical model for the “Big Wave” is simple.
+The solar energy output changes could determine the “Big Wave”.
 It is no surprise that the Sun warms the largest battery on the Earth, the Pacific Ocean.
-And if many cycles moderate the solar energy output, it is no surprise either that we have failed to discover the interference of these cycles, El Niño, in our yearly climate temperatures.
+And if many cycles moderate the solar energy output, it is no surprise either that we have failed to discover the interference of these cycles, the “Big Wave”, in our yearly climate temperatures.
+The immense Pacific Ocean surface acts like a giant thermometer that measures solar energy output in real time.
 The current El Niño ”tool kit” is admittedly impressive [6].
@@
 Yet, even the best currently available physical El Niño forecasts fail in 1.5 years [62].
-Our mathematical DCM forecast can connect seven El Niño decades to the next seven El Niño decades (Figures A1 and A2).
+Our mathematical DCM forecast can connect seven “Big Wave” decades to the next seven “Big Wave” decades (Figures S1 and S2).
 That is quite a coincidence if the chaos of physical models takes over in 1.5 years.
@@
 https://zenodo.org/records/18726181/
-No supercomputers are needed because an ordinary PC can detect the first two signals in about one minute. The third signal detection takes a few minutes.
+No supercomputers are needed because an ordinary PC can detect the first two “Big Wave” signals in about one minute. The third signal detection takes a few minutes.
 William of Ockham (1287–1347) stated that if two competing hypotheses make the same predictions, the simplest hypothesis is usually the best one (Occam’s razor).
-The connection between El Niño, the Sun and the Planets may seem like the holy grail.
+The simple connection between El Niño  “Big Wave”, the Sun and the Planets may seem like the holy grail.
 However, those three are just physical phenomena.
 The real scientific holy grail is the DCM.
+Note added to the proof
+We arrive at a "Mexican standoff" because only future data can confirm that our “Big Wave” forecasts are correct. Fortunately, one minor detail saves us one year of waiting.
+On January 29th, 2026, we computed the yearly values from the NOAA1 data.
+Four monthly values for the year 2025 were missing and our analysed data ended in the year 2024.
+Today February 25th 2026, NOAA1 still has these same four missing monthly values for the year 2025 (“-99.99”).
+To get these missing values, we are advised to use the Climate Prediction Center (CPC) data in NOAA2.666https://www.cpc-ncep.noaa.gov/products/
+analysis-monitoring/ensostuff/ON-v5.php
+All twelve monthly values for the year 2025 are given in Table S7 (column 2).
+These values give $y_{i}\pm 1\sigma_{i}=-0.26\pm 0.21$ $\mathrm{C}^{o}$.
+A red circle now highlights this observed new 2025 year value which is added into Figures S3e and Figures S4e.
+Our “Big Wave” forecast for the year 2025 succeeds.
+The year 2025 forecast for the best $\mathcal{M}$=3 model for all data (Table S3) is more accurate than the forecast for the unstable $\mathcal{M}$=4 model.
+Our forecast for this 2025 observation can, of course, be dismissed just as one lucky guess.
+Who wants to take the risk that our severe “Big Wave” forecasts for the years 2030 and 2031 are also “lucky”?
+We also analyse the monthly SST values (Tables S8-S19, Figures S5-S16).
+The red dots in our figures highlight the available new monthly values between January 2025 and February 2026.
+Most of those red dots hit their mark.
+We consistently detect the “new” 3.6 and 4.6 year periods from these monthly values.
+The monthly mean computation eliminates the detection of these periods from yearly data.
+We conclude that the chaotic $\pm 1$$\mathrm{C}^{o}$ weather changes can prevent accurate DCM monthly forecasts.
+However, we can compute the “Big Wave” monthly values (Table S20).
+We admit that this does not fully solve the El Niño riddle but it may help.
+If no internal mechanism can cause the strictly multi-periodic “Big Wave” in the Pacific Ocean then the Sun is the only possible external cause.
+The planets may indeed cause regular solar cycles [30].
 Acknowledgements
-This work has made use of NASA’s Astrophysics Data System (ADS) services.
+This work has made use of NASA’s Astrophysics Data System (ADS) services. We thank Dr. Jouni Räisänen for his comments of the manuscript.
 Author contributions
@@
 Ethical approval and consent to participate Not applicable.
+Supplementary material
 Refer to caption Refer to caption
-            (a)             (b)
+            (b)             (c)
 Refer to caption Refer to caption
-            (c)             (d)
-Figure A1: First half forecast (Table A2: $\mathcal{M}$=2): Forecasting data (closed black circles) and forecasted data (open black circles).
+            (d)             (e)
+Figure S1: First half forecast (Table S2: $\mathcal{M}$=2): (a) Periodogram (Equation 21).
+(b-e) Forecasting data (closed black circles) and forecasted data (open black circles).
 Green and red dottted lines denote one and three sigma model $g(t)$ errors.
@@
 Refer to caption Refer to caption
-            (a)             (b)
+            (b)             (c)
 Refer to caption Refer to caption
-            (c)             (d)
-Figure A2: First half forecast (Table A2: $\mathcal{M}$=3). Otherwise as in Figure A1.
-Figure A3: All data model $\mathcal{M}$=3 in Table A3.
-(a) Periodogram (Equation 21).
-Otherwise as in Figures 2ab.
-(b) Three $h_{1}(t)$, $h_{2}(t)$ and $h_{3}(t)$ signals as a function of time $t$ (left-hand panels) and phase $\phi$ (right-hand panels) [30, Equation 29].
+            (d)             (e)
+Figure S2: First half forecast (Table S2: $\mathcal{M}$=3). Otherwise as in Figure S1.
 Refer to caption Refer to caption
-            (a)             (b)
+            (b)             (c)
 Refer to caption Refer to caption
-            (c)             (d)
-Figure A4: All data model $\mathcal{M}$=3 in Table A3. Otherwise as in Figure A1.
+            (d)             (e)
+Figure S3: “Big Wave” for all yearly mean data model $\mathcal{M}$=3 in Table S3. Red circle is added observed 2025 year value. Otherwise as in Figure S1.
 Refer to caption Refer to caption
-            (a)             (b)
+            (b)             (c)
 Refer to caption Refer to caption
-            (c)             (d)
-Figure A5: All data model $\mathcal{M}$=4 in Table A3. Otherwise as in Figure A1.
-Table A1: Trend in all weighted data. DCM analysis between $P_{\mathrm{min}}=3$ and $P_{\mathrm{max}}=100$ years. Otherwise as in Table 8
+            (d)             (e)
+Figure S4: “Big Wave” for all yearly mean data model $\mathcal{M}$=4 in Table S3. Otherwise as in Figure S2.
+Table S1: Trend in all yearly mean weighted data.
+DCM analysis between $P_{\mathrm{min}}=3$ and $P_{\mathrm{max}}=100$ years.
+We use $\gamma=0.001$ (Equation 48) for yearly data.
+Otherwise as in Table 8
 Period analysis Fisher-test
@@
 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 299 $1873.8\pm 1.0$ $1870.18\pm 0.99$ $1870.6\pm 0.97$ $1871.4\pm 2.2$ - -
-Table A2: Periods in first half data. Otherwise as in Table A1.
+Table S2: Periods in first half yaerly mean data. Otherwise as in Table S1.
 Period analysis Fisher-test
@@
 92 $1872.51\pm 0.66$ $1875.6\pm 1.2$ $1870.8\pm 1.1$ $1870.9\pm 2.9$ - - -
-Table A3: Periods in all weighted data. Otherwise as in Table A1.
+Table S3: Periods in all yearly mean weighted data. Otherwise as in Table S1.
 Period analysis Fisher-test
@@
 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 300 $1873.8\pm 1.2$ $1870.2\pm 1.1$ $1870.6\pm 0.97$ $1871.3\pm 2.$ - - -
-Table A4: Forecasts (Figures A4 and A5: open circles).
-Figure A4 Figure A5
+Table S4: “Big Wave”yearly forecasts (Figures S3d and S4d: open circles).
+Figure S3 Figure S4
 $t$ $y$ $\pm 1\sigma$ $y$ $\pm 1\sigma$
 (y) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$)
-2025 0.09 0.24 0.35 0.28
-2026 0.35 0.20 0.47 0.29
-2027 0.41 0.21 0.31 0.28
-2028 0.35 0.22 0.21 0.28
-2029 0.40 0.19 0.42 0.34
-2030 0.60 0.20 0.79 0.31
-2031 0.79 0.20 0.94 0.21
-2032 0.75 0.27 0.66 0.21
-2033 0.45 0.34 0.18 0.23
-2034 0.11 0.30 -0.10 0.31
-2035 -0.03 0.21 -0.03 0.38
-2036 0.06 0.22 0.16 0.34
-Table A5: Periods in all non-weighted data. Otherwise as in Table A1.
+2025 0.3 0.2 0.2 0.2
+2026 -0.2 0.2 0.2 0.2
+2027 -0.4 0.3 -0.0 0.2
+2028 0.2 0.3 -0.1 0.2
+2029 0.5 0.3 0.2 0.2
+2030 0.3 0.3 0.4 0.2
+2031 0.1 0.2 0.4 0.2
+2032 0.2 0.3 0.2 0.2
+2033 0.2 0.3 0.1 0.3
+2034 0.0 0.2 0.0 0.2
+2035 0.1 0.2 -0.1 0.3
+2036 -0.0 0.3 -0.2 0.3
+Table S5: Periods in all non-weighted yearly mean data. Otherwise as in Table S1.
 Period analysis Fisher-test
@@
 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 25.3 $1873.8\pm 1.2$ $1870.2\pm 1.2$ $1870.6\pm 0.97$ $1871.3\pm 2.1$ -
-Table A6: Significance one signal model $\mathcal{M}$=1 in Table A3. Otherwise as in Table 9. 44footnotetext: Note: No test for $\mathcal{M}$=5  $(\eta_{1}=\eta_{2}=5)$.
+Table S6: Significance one signal model $\mathcal{M}$=1 in Table S3. Otherwise as in Table 9. 66footnotetext: Note: No test for $\mathcal{M}$=5  $(\eta_{1}=\eta_{2}=5)$.
 (1) (2)
@@
 […]
+Table S7: Year 2025.
+Temperature anomalies $(y)$ of 12 months.
+(1) NOAA1
+(2) NOAA2
+$y$ $y$
+$\mathrm{C}^{o}$ $\mathrm{C}^{o}$
+(1) (2)
+-0.50 -0.4
+-0.36 -0.2
+-0.13 -0.1
+-0.09 0.0
+0.02 0.0
+0.14 0.0
+0.05 -0.1
+0.06 -0.3
+-99.99 -0.4
+-99.99 -0.5
+-99.99 -0.6
+-99.99 -0.5
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S5: January data model $\mathcal{M}$=4 in Table S8. Red dots are -0.76 and -0.54 for 2025 and 2026. Otherwise as in Figure S2.
+Table S8: Periods in January data. Note that we the pre-assigned sigficance level $\gamma=0.05$ (Equation 48) because monthly SST data are so noisy. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Jan.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $3.766\pm 0.030$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.46\pm 0.12$ - - - 3.33 3.40 3.26 JanR14K111.dat
+64.0 $1872.14\pm 0.45$ - - - 0.021 0.0035 0.0012
+Two signals
+$\mathcal{M}$=2 $3.767\pm 0.040$ $4.734\pm 0.046$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.45\pm 0.11$ $0.46\pm 0.13$ - - - 3.14 3.09 JanR14K211.dat
+59.9 $1872.14\pm 0.41$ $1870.98\pm 0.59$ - - - 0.027 0.0071
+Three signals
+$\mathcal{M}$=3 $3.766\pm 0.045$ $4.734\pm 0.026$ $12.72\pm 0.18$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.45\pm 0.10$ $6\pm 0.11 0.44±0$$4\pm 0.11$ - - - 2.91 JanR14K311.dat
+56.2 $1872.15\pm 0.43$ $1870.98\pm 0.43$ $1871.6\pm 1.2$ - - - 0.037
+Four signals
+$\mathcal{M}$=4 $3.765\pm 0.018$ $4.737\pm 0.048$ $5.643\pm 0.15$ $12.68\pm 0.30$ - - -
+$\mathcal{M}$4,1,1,R 14 $5\pm 0.12 0.44±0$$4\pm 0.12$ $0.41\pm 0.10$ $0.43\pm 0.11$ - - - JanR14K411.dat
+52.9 $1872.17\pm 0.45$ $1870.90\pm 0.67$ $1870.80\pm 0.66$ $1871.9\pm 1.4.$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S6: February data model $\mathcal{M}$=4 in Table S9. Red dots are -0.35 and -0.40 for 2025 and 2026. Otherwise as in Figure S3
+Table S9: Periods in February data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Feb.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $5.652\pm 0.088$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.48\pm 0.13$ - - - 3.55 3.36 3.23 FebR14K111.dat
+55.8 $1870.60\pm 0.76$ - - - 0.016 0.0039 0.0013
+Two signals
+$\mathcal{M}$=2 $5.656\pm 0.025$ $12.74\pm 0.22$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.47\pm 0.12$ $0.442\pm 0.086$ - - - 3.02 2.92 FebR14K211.dat
+52.0 $1870.55\pm 0.40$ $1871.2\pm 1.2$ - - - 0.032 0.010
+Three signals
+$\mathcal{M}$=3 $4.7314PM0.050$ $5.672\pm 0.032$ $12.68\pm 0.44$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.40\pm 0.11$ $0.450\pm 0.098$ $0.42\pm 0.10$ - - - 2.72 FebR14K311.dat
+48.9 $1870.840\pm 0.69$ $1870.37\pm 0.54$ $1871.6\pm 1.6$ - - - 0.046
+Four signals
+$\mathcal{M}$=4 $3.771\pm 0.035$ $4.741\pm 0.058$ $5.647\pm 0.031$ $12.73\pm 0.22$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.3\pm 1.8$ $0.40\pm 0.10$ $0.45\pm 0.13$ $0.436\pm 0.094$ - - - FebR14K411.dat
+46.2 $1873.8\pm 1.2$ $1870.8\pm 0.53$ $1870.68\pm 0.040$ $1871.32\pm 1.7$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S7: March data model $\mathcal{M}$=3 in Table S10. Red dot is -0.01 for 2025. Otherwise as in Figure S3
+Table S10: Periods in March data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Mar.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $5.665\pm 0.029$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.451\pm 0.086$ - - - 3.94 4.27 3.71 MarR14K111.dat
+44.1 $1870.46\pm 0.45$ - - - 0.0097 0.00054 0.00033
+Two signals
+$\mathcal{M}$=2 $5.669\pm 0.042$ $12.68\pm 0.56$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.44\pm 0.11$ $0.412\pm 0.085$ - - - 4.33 3.41 MarR14K211.dat
+40.8 $1870.41\pm 0.72$ $1871.7\pm 1.6$ - - - 0.0059 0.0036
+Three signals
+$\mathcal{M}$=3 $3.633\pm 0.019$ $5.644\pm 0.052$ $12.73\pm 0.44$ - - - $\leftarrow$
+$\mathcal{M}$3,1,1,R 11 $0.418\pm 0.090$ $0.45\pm 0.10$ $0.43\pm 0.13$ - - - 2.36 MarR14K311.dat
+37.4 $1872.96\pm 0.37$ $1870.74\pm 0.81$ $1871.2\pm 1.2$ - - - 0.074
+Four signals
+$\mathcal{M}$=4 $3.633\pm 0.027$ $4.585\pm 0.079$ $5.671\pm 0.034$ $12.68\pm 0.15$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.419\pm 0.093$ $0.311\pm 0.084$ $0.45\pm 0.10$ $0.418\pm 0.080$ - - - MarR14K411.dat
+35.6 $1872.94\pm 0.37$ $1870.30\pm 0.92$ $1870.42\pm 0.51$ $1871.7\pm 1.1$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S8: April data model $\mathcal{M}$=4 in Table S11. Red dot is -0.20 for 2025. Otherwise as in Figure S3
+Table S11: Periods in April data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Apr.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $12.68\pm 0.13$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.42\pm 0.10$ - - - 5.06 4.29 3.92 AprR14K111.dat
+30.8 $1871.6\pm 1.0$ - - - 0.0023 0.00052 0.00018
+Two signals
+$\mathcal{M}$=2 $5.663\pm 0.052$ $12.66\pm 0.14$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.388\pm 0.073$ $0.416\pm 0.068$ - - - 3.29 3.13 AprR14K211.dat
+27.9 $1870.44\pm 0.62$ $1871.8\pm 1.0$ - - - 0.022 0.0065
+Three signals
+$\mathcal{M}$=3 $5.665\pm 0.017$ $12.73\pm 0.13$ $18.84\pm 0.84$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.382\pm 0.097$ $0.416\pm 0.068$ $0.308\pm 0.065$ - - - 2.82 AprR14K311.dat
+26.1 $1870.42\pm 0.29$ $1871.36\pm 0.85$ $1873.7\pm 2.0$ - - - 0.040
+Four signals
+$\mathcal{M}$=4 $3.760\pm 0.039$ $5.669\pm 0.069$ $12.74\pm 0.12$ $18.87\pm 0.37$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.276\pm 0.074$ $0.382\pm 0.096$ $0.42\pm 0.13$ $0.310\pm 0.090$ - - - AprR14K411.dat
+24.6 $1871.98\pm 0.44$ $1870.36\pm 0.47$ $1871.30\pm 0.84$ $1873.6\pm 2.2$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S9: May data model $\mathcal{M}$=4 in Table S12. Red dot is -0.04 for 2025. Otherwise as in Figure S3
+Table S12: Periods in May data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: May.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $12.67\pm 0.24$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.36\pm 0.13$ - - - 5.67 3.39 3.38 MayR14K111.dat
+29.7 $1872.0\pm 1.7$ - - - 0.0010 0.0037 0.00088
+Two signals
+$\mathcal{M}$=2 $5.664\pm 0.045$ $12.65\pm 0.15$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.329\pm 0.073$ $0.353\pm 0.075$ - - - 2.93 3.37 MayR14K211.dat
+27.6 $1870.38\pm 0.54$ $1872.1\pm 1.2$ - - - 0.036 0.0060
+Three signals
+$\mathcal{M}$=3 $3.388\pm 0.025$ $5.666\pm 0.062$ $12.65\pm 0.30$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.288\pm 0.090$ $0.330\pm 0.057$ $0.356\pm 0.080$ - - - 3.26 MayR14K311.dat
+26.0 $1873.05\pm 0.51$ $1870.38\pm 0.41$ $1872.18\pm 0.92$ - - - 0.023
+Four signals
+$\mathcal{M}$=4 $3.383\pm 0.024$ $3.551\pm 0.042$ $5.666\pm 0027$ $12.65\pm 0.12$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.304\pm 0.098$ $0.30\pm 0.10$ $0.332\pm 0.067$ $0.356\pm 0.081$ - - - MayR14K411.dat
+24.3 $1873.15\pm 0.50$ $1871.66\pm 0.54$ $1870.37\pm 0.40$ $1872.14\pm 0.90$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S10: June data model $\mathcal{M}$=4 in Table S13. Red dot is -0.02 for 2025. Otherwise as in Figure S3
+Table S13: Periods in June data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Jun.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $12.72\pm 0.15$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.42\pm 0.10$ - - - 4.67 4.07 4.00 JunR14K111.dat
+27.4 $1871.5\pm 1.1$ - - - 0.0038 0.00084 0.00014
+Two signals
+$\mathcal{M}$=2 $4.740\pm 0.022$ $12.72\pm 0.14$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.344\pm 0.076$ $0.424\pm 0.082$ - - - 3.26 3.42 JunR14K211.dat
+25.0 $1870.28\pm 0.52$ $1871.60\pm 0.97$ - - - 0.023 0.0035
+Three signals
+$\mathcal{M}$=3 $4.740\pm 0.018$ $12.75\pm 0.15$ $22.46\pm 0.55$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.342\pm 0.072$ $0.411\pm 0.072$ $0.297\pm 0.068$ - - - 3.42 JunR14K311.dat
+23.4 $1870.27\pm 0.30$ $1871.40\pm 0.92$ $1891.0\pm 2.1$ - - - 0.019
+Four signals
+$\mathcal{M}$=4 $3.647\pm 0.012$ $4.741PM0.037$ $12.76\pm 0.12$ $22.49\pm 1.7$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.289\pm 0.069$ $0.338\pm 0.094$ $0.415\pm 0.077$ $0.302\pm 0.086$ - - - JunR14K411.dat
+21.8 $1872.07\pm 0.35$ $1870.23\pm 0.65$ $1871.35\pm 0.85$ $1891.0\pm 1.8$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S11: July data model $\mathcal{M}$=4 in Table S14. Red dot is -0.10 for 2025. Otherwise as in Figure S3
+Table S14: Periods in July data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Jul.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $12.76\pm 0.17$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.372\pm 0.049$ - - - 4.07 4.23 4.24 JulR14K111.dat
+28.5 $1870.8\pm 1.3$ - - - 0.0082 0.00059 $7.1\times 10^{-5}$
+Two signals
+$\mathcal{M}$=2 $12.77\pm 0.18$ $20.0\pm 1.0$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $71\pm 0.078 0.330$$30\pm 0.081$ - - - 4.14 4.06 JulR14K211.dat
+26.3 $1870.6\pm 1.5$ $1872.50\pm 2.2$ - - - 0.0076 0.00087
+Three signals
+$\mathcal{M}$=3 $4.738\pm 0.048$ $12.77\pm 0.16$ $19.90\pm 0.71$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.331\pm 0.085$ $0.369\pm 0.078$ $0.327\pm 0.069$ - - - 3.75 JulR14K311.dat
+24.2 $1870.23\pm 0.44$ $1870.7\pm 1.0$ $1872.7\pm 1.8$ - - - 0.012
+Four signals
+$\mathcal{M}$=4 $3.545\pm 0.010$ $4.740\pm 0.043$ $12.77\pm 0.12$ $19.97\pm 0.98$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.302\pm 0.070$ $0.333\pm 0.08$ $0.370\pm 0.081$ $0.329\pm 0.047$ - - - JulR14K411.dat
+22.4 $1871.68\pm 0.34$ $1870.19\pm 0.42$ $1870.62\pm 0.80$ $1872.5\pm 1.7$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S12: August data model $\mathcal{M}$=4 in Table S15. Red dot is -0.30 for 2025. Otherwise as in Figure S3
+Table S15: Periods in August data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Aug.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $3.643\pm 0.036$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.38\pm 0.14$ - - - 3.86 37.4 3.56 AugR14K111.dat
+35.4 $1871.87\pm 0.44$ - - - 0.011 0.0017 0.00051
+Two signals
+$\mathcal{M}$=2 $3.647\pm 0.011$ $20.58\pm 0.89$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.38\pm 0.10$ $0.368\pm 0.071$ - - - 3.43 3.26 Aug14K211.dat
+32.8 $1871.88\pm 0.26$ $1871.1\pm 1.7$ - - - 0.019 0.0050
+Three signals
+$\mathcal{M}$=3 $3.647\pm 0.017$ $12.56\pm 0.42$ $20.58\pm 0.44$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.38\pm 0.10$ $0.331\pm 0.081$ $0.367\pm 0.071$ - - - 2.92 AugR14K311.dat
+30.6 $1871.88\pm 0.41$ $1872.4\pm 1.7$ $1871.2\pm 1.9$ - - - 0.036
+Four signals
+$\mathcal{M}$=4 $3.648\pm 0.010$ $9.13\pm 0.19$ $12.51\pm 0.31$ $20.6\pm 1.2$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.382\pm 0.084$ $0.311\pm 0.064$ $0.334\pm 0.084$ $0.353\pm 0.092$ - - - AugR14K411.dat
+28.8 $1871.88\pm 0.26$ $1872.31\pm 0.99$ $1872.9\pm 1.2$ $1871.3\pm 2.4$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S13: September data model $\mathcal{M}$=3 in Table S16. Red dot is -0.40 for 2025. Otherwise as in Figure S3
+Table S16: Periods in September data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Sep.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $9.12\pm 0.11$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.350\pm 0.078$ - - - 2.84 2.81 2.78 SepR14K111.dat
+43.5 $1872.5\pm 1.3$ - - - 0.040 0.013 0.0050
+Two signals
+$\mathcal{M}$=2 $3.546\pm 0.022$ $20.10\pm 0.68$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.349\pm 0.087$ $0.3550\pm 0.076$ - - - 2.70 2.66 Sep14K211.dat
+41.1 $1871.60\pm 0.43$ $1871.6\pm 2.6$ - - - 0.048 0.018
+Three signals
+$\mathcal{M}$=3 $3.546\pm 0.018$ $9.13\pm 0.21$ $20.0\pm 1.4$ - - - $\leftarrow$
+$\mathcal{M}$3,1,1,R 11 $0.35\pm 0.11$ $0.340\pm 0.094$ $0.341\pm 0.083$ - - - 2.53 SepR14K311.dat
+38.9 $1871.61\pm 0.37$ $1872.47\pm 0.96$ $1871.9\pm 3.5$ - - - 0.060
+Four signals
+$\mathcal{M}$=4 $3.545\pm 0.025$ $4.749\pm 0.071$ $9-12\pm 0.11$ $19.92\pm 0.93$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.35\pm 0.12$ $0.32\pm 0.11$ $0.343\pm 0.079$ $0.340\pm 0.082$ - - - SepR14K411.dat
+36.9 $1871.64\pm 0.30$ $1874.44\pm 0.63$ $1872.5\pm 1.1$ $1872.0\pm 3.1$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S14: October data model $\mathcal{M}$=4 in Table S17. Red dot is -0.48 for 2025. Otherwise as in Figure S3
+Table S17: Periods in October data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Oct.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $3.354\pm 0.049$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.38\pm 0.13$ - - - 2.51 2.70 2.78 OctR14K111.dat
+56.0 $1871.59\pm 0.53$ - - - 0.039 0.016 0.0049
+Two signals
+$\mathcal{M}$=2 $3.542\pm 0.022$ $4.576\pm 0.045$ - - - $\leftarrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.38\pm 0.11$ $0.40\pm 0.12$ - - - 2.46 2.65 Oct14K211.dat
+52.9 $1871.62\pm 0.41$ $1874.13\pm 0.64$ - - - 0.065 0.018
+Three signals
+$\mathcal{M}$=3 $3.542\pm 0.022$ $4.578\pm 0.047$ $12.82\pm 0.47$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.381\pm 0.093$ $0.419\pm 0.11$ $0.367\pm 0.093$ - - - 2.75 OctR14K311.dat
+50.3 $1871.63\pm 0.52$ $1874.09\pm 0.44$ $1882.6PM1.3$ - - - 0.045
+Four signals
+$\mathcal{M}$=4 $3.542\pm 0.035$ $5.579\pm 0.058$ $12.84\pm 0.39$ $20.4\pm 1.0$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.38\pm 0.10$ $0.40\pm 0.15$ $0.38\pm 0.13$ $0.38\pm 0.11$ - - - OctR14K411.dat
+47.5 $1871.64\pm 0.50$ $1874.07\pm 0.52$ $1882.2\pm 1.1$ $1871.3\pm 3.0$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S15: Novermber data model $\mathcal{M}$=3 in Table S18. Red dot is -0.80 for 2025. Otherwise as in Figure S3
+Table S18: Periods in November data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Nov.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $5.665\pm 0.052$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.47\pm 0.10$ - - - 2.86 3.02 3.08 NovR14K111.dat
+64.8 $1875.17\pm 0.75$ - - - 0.039 0.0081 0.0021
+Two signals
+$\mathcal{M}$=2 $3.647\pm 0.019$ $5.668\pm 0.078$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.43\pm 0.12$ $0.47\pm 0.14$ - - - 3.07 3.06 Nov14K211.dat
+61.2 $1871.67\pm 0.42$ $1875.13\pm 0.59$ - - - 0.039 0.0076
+Three signals
+$\mathcal{M}$=3 $3.647\pm 0.041$ $5.670\pm 0.049$ $12.86\pm 0.63$ - - - $\leftarrow$
+$\mathcal{M}$3,1,1,R 11 $0.44\pm 0.14$ $0.46\pm 0.12$ $0.43\pm 0.13$ - - - 2.67 NovR14K311.dat
+57.5 $1871.68\pm 0.42$ $1875.09\pm 0.57$ $1882.2\pm 2.2$ - - - 0.050
+Four signals
+$\mathcal{M}$=4 $3.647\pm 0.029$ $5.67\pm 0.11$ $9.10\pm 0.20$ $12.90\pm 0.31$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.43\pm 0.13$ $0.47\pm 0.15$ $0.42\pm 0.12$ $0.432\pm 0.096$ - - - NovR14K411.dat
+54.1 $1871.67\pm 0.53$ $1875.14\pm 0.67$ $72.67\pm 0.77 1881$$81.9\pm 1.2$ - - -
+Refer to caption Refer to caption
+            (a)             (b)
+Refer to caption Refer to caption
+            (c)             (d)
+Figure S16: December data model $\mathcal{M}$=4 in Table S19. Red dot is -0.92 for 2025. Otherwise as in Figure S3
+Table S19: Periods in December data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=155,\Delta T=154$: Dec.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $5.645\pm 0.033$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.48\pm 0.13$ - - - 4.97 4.20 3.94 DecR14K111.dat
+64.8 $1871.70\pm 0.44$ - - - 0.0026 0.00064 0.00017
+Two signals
+$\mathcal{M}$=2 $3.645\pm 0.016$ $12.75\pm 0.38$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.49\pm 0.15$ $0.50\pm 0.16$ - - - 3.20 3.20 Dec14K211.dat
+58.8 $1871.71\pm 0.36$ $1870.3\pm 1.5$ - - - 0.025 0.0056
+Three signals
+$\mathcal{M}$=3 $3.643\pm 0.029$ $5.664\pm 0.090$ $12.74\pm 0.18$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.49\pm 0.10$ $0.43\pm 0.10$ $0.50\pm 0.11$ - - - 3.06 DecR14K311.dat
+55.1 $1871.73\pm 0.31$ $1875.23\pm 0.67$ $1870.4\pm 1.1$ - - - 0.030
+Four signals
+$\mathcal{M}$=4 $3.644\pm 0.040$ $5.566\pm 0.020$ $5.664\pm 0.058$ $12.74\pm 0.32$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.49\pm 0.10$ $0.42\pm 0.10$ $0.428\pm 0.092$ $0.50\pm 0.12$ - - - DecR14K411.dat
+51.7 $1871.76\pm 0.38$ $1874.40\pm 0.38$ $1875.26\pm 0.50$ $1870.4\pm 2.0$ - - -
+Table S20: “Big Wave” monthly forecast 2025-2036.
+Year Jan Feb Mar Apr May Jun Jul Aug Sep Oct Nov Dec
+($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$)
+2025 $0.3\pm 0.2$ $0.4\pm 0.2$ $0.2\pm 0.2$ $0.1\pm 0.2$ $0.2\pm 0.2$ $0.3\pm 0.2$ $2\pm 0.2 0.0$$0\pm 0.2$ $2\pm 0.2 0.0$$\pm 0.3 0.0$$0\pm 0.2$ $0.3\pm 0.2$
+2026 $0.0\pm 0.3$ $1\pm 0.2 0.0$$0\pm 0.2$ $0.1\pm 0.1$ $0.3\pm 0.1$ $0.3\pm 0.2$ $0.1\pm 0.2$ $\pm 0.2 0.2±0.2$$\pm 0.2$ $0.2\pm 0.2$ $0.3\pm 0.3$ $0.2\pm 0.3$
+2027 $0.3\pm 0.2$ $0.2\pm 0.3$ $0.4\pm 0.2$ $0.1\pm 0.2$ $0.4\pm 0.1$ $0.1\pm 0.1$ $0.1\pm 0.2$ $0.4\pm 0.1$ $0.2\pm 0.2$ $0.3\pm 0.2$ $0.3\pm 0.2$ $0.5\pm 0.2$
+2028 $0\pm 0.2 0.0$$0\pm 0.2$ $\pm 0.2 0.2±0.2$$\pm 0.2$ $0.0\pm 0.2$ $0.1\pm 0.2$ $0.1\pm 0.2$ $0.2\pm 0.2$ $\pm 0.2 0.3±0.3$$\pm 0.3$ $1\pm 0.2 0.0$$0\pm 0.3$
+2029 $0.3\pm 0.3$ $0.3\pm 0.2$ $0.2\pm 0.2$ $1\pm 0.1 0.0$$0\pm 0.2$ $0.2\pm 0.2$ $0.5\pm 0.2$ $1\pm 0.2 0.0$$0\pm 0.2$ $3\pm 0.2 0.0$$0\pm 0.3$ $0.1\pm 0.2$
+2030 $0.2\pm 0.2$ $0.3\pm 0.2$ $\pm 0.2 0.1±0.1$$\pm 0.1$ $0.5\pm 0.2$ $0.4\pm 0.2$ $0.3\pm 0.2$ $4\pm 0.2 0.0$$0\pm 0.2$ $0.1\pm 0.2$ $0.6\pm 0.2$ $0.4\pm 0.2$
+2031 $0.3\pm 0.3$ $0.3\pm 0.2$ $0.6\pm 0.2$ $0.6\pm 0.2$ $0.4\pm 0.2$ $0.3\pm 0.1$ $0.1\pm 0.1$ $0.5\pm 0.1$ $1\pm 0.2 0.0$$0\pm 0.2$ $0.7\pm 0.2$ $0.6\pm 0.2$
+2032 $0.4\pm 0.2$ $0.4\pm 0.2$ $0.4\pm 0.2$ $0.7\pm 0.2$ $0.2\pm 0.1$ $1\pm 0.1 0.0$$0\pm 0.2$ $1\pm 0.1 0.0$$0\pm 0.3$ $0.3\pm 0.3$ $0.2\pm 0.3$ $0.3\pm 0.2$
+2033 $0.1\pm 0.3$ $2\pm 0.2 0.0$$0\pm 0.2$ $0.4\pm 0.2$ $0.4\pm 0.2$ $0.3\pm 0.2$ $3\pm 0.1 0.0$$0\pm 0.2$ $0.2\pm 0.2$ $2\pm 0.3 0.0$$\pm 0.3 0.0$$0\pm 0.2$
+2034 $0.2\pm 0.3$ $2\pm 0.3 0.0$$\pm 0.2 0.0$$0\pm 0.1$ $0.2\pm 0.2$ $0.5\pm 0.2$ $0.1\pm 0.2$ $1\pm 0.2 0.0$$0\pm 0.2$ $0.4\pm 0.3$ $0.1\pm 0.3$ $0.2\pm 0.3$
+2035 $0.1\pm 0.3$ $0.1\pm 0.3$ $\pm 0.3 0.1±0.1$$\pm 0.1$ $0.4\pm 0.2$ $0.3\pm 0.2$ $\pm 0.1 0.2±0.2$-$\pm 0.2$ $0.2\pm 0.3$ $0.5\pm 0.3$ $0.1\pm 0.2$ $0.4\pm 0.3$
+2036 $0\pm 0.3 0.0$$0\pm 0.2$ $0.3\pm 0.3$ $0.1\pm 0.2$ $0.1\pm 0.2$ $0.2\pm 0.1$ $0.2\pm 0.1$ $5\pm 0.1 0.0$$\pm 0.2 0.0$$0\pm 0.2$ $0.1\pm 0.2$ $0.1\pm 0.2$

### v8

@@
-Discrete Chi-Square Method can model and forecast complex time series, like El Nino data between 1870 and 2024
+Discrete Chi-Square Method can model and forecast complex time series, like El Nino data between 1870 and 2026
 Abstract
@@
 Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series.
-In our appendix, we show that the DCM can model and forecast El Nino data between 1870 and 2024.
+In our appendix, we show that the DCM can model and forecast El Nino data between 1870 and 2026.
 An immediate, independent and objective validity check of our analysis may save some money.
@@
 Finally, we discuss our results (Section 4).
-In our Appendix, we show that the DCM can model and forecast El Niño (Section A1).
+In our Appendix, we show that the DCM can model and forecast El Niño (Section Appendix 1).
 2 Methods
@@
 (g) DFT model detrended $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (black continuous line), DFT pure sine $s_{\mathrm{DFT}}(t)$ (red thick continuous line), detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and DFT model residuals (blue dots) offset to -1.5 level (blue dotted line).
-The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [31, TSPA, Equation 1] and the Continuous Period Search [39, CPS, Equation 3].
+The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [32, TSPA, Equation 1] and the Continuous Period Search [40, CPS, Equation 3].
 TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$.
@@
 We search for signals in these simulated time series by using the frequently applied DFT version formulated by [25] [^1] This paper had over 2600 citations in December 2025.
-Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [24, Equation 1].
+Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [25, Equation 1].
 The notations for the DFT model are
@@
 Our DFT analyses of simulated time series may fail due to AL6-AL10.
-The time series is “too short” [32, AL6]
+The time series is “too short” [33, AL6]
 $P>\Delta T.$ (35)
-The time series is non-stationary because it contains “a trend” [45, AL7]
+The time series is non-stationary because it contains “a trend” [46, AL7]
 $p_{\mathrm{DFT}}(t)\neq 0.$ (36)
-Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [32, AL9].
+Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [33, AL9].
 $|f_{1}-f_{2}|<f_{0}={\Delta T}^{-1}.$ (37)
@@
 We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large.
-One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu [30, Table S7].
+One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu [31, Table S7].
 We conclude that the best model for the real time series can be found by applying the Fisher-test to any arbitrary number of different nested DCM or DFT models.
@@
 In other words, the signal significances decrease.
-Typical examples can be found in Jetsu [30, Tables S5-S1].
+Typical examples can be found in Jetsu [31, Tables S5-S1].
 No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected.
 For the Model 1 simulated $n=100$ and ${\mathrm{SN}}=100$ combination, Table 8 shows that this time series contains only one signal.
-For this Model 1 time series, there is no need to discuss the DFT significance estimates [24, their Equation 22] because this method fails to detect the correct period.
+For this Model 1 time series, there is no need to discuss the DFT significance estimates [25, their Equation 22] because this method fails to detect the correct period.
 We conclude that the Fisher-test identifies the correct DCM model (Section 3.9), as well as gives the signal significance $Q_{\mathrm{F}}$ estimates (Section 3.10).
@@
 The artificial bootstrap data sets (Equation 31) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation).
-We routinely check the stability of these solutions [30, Figure S5].
+We routinely check the stability of these solutions [31, Figure S5].
 The unstable models, where the model parameter changes are large, are rejected (Section 2.1: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}” models).
@@
 3.12 Forecast
-There are numerous techniques for forecasting a time series [e.g., 23, 28, 33].
+There are numerous techniques for forecasting a time series [e.g., 24, 29, 34].
 The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast.
@@
 Perhaps our DCM should be re-named the Slow Gauß Transform (SGT).
-A1 Appendix
+Appendix 1
 Introduction
@@
 The forecasts of the two and three signals models, open circles, are shown in Figures S1 and S2.
+Both forecasts reproduce the overall yearly mean changes during the second half.
 For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signal models are 421, 336, 370 and 428, respectively.
@@
 The Fisher-test comparison between the $K_{3}=0$ pure noise model $g_{0,1,0,\chi^{2}}$ and the one signal model $g_{1,1,1,\chi^{2}}$ gives signal significance $Q_{\mathrm{F}}=3.9\times 10^{-13}$ (Table S6, $\mathcal{M}$=1).
+This is the probability that the first detected signal is not real.
 All other polynomial model alternatives, $1\leq K_{3}\leq 7$, are also rejected.
@@
 And if many cycles moderate the solar energy output, it is no surprise either that we have failed to discover the interference of these cycles, the “Big Wave”, in our yearly climate temperatures.
-The immense Pacific Ocean surface acts like a giant thermometer that measures solar energy output in real time.
+The immense Pacific Ocean surface may act like a giant thermometer that measures solar energy output in real time.
 The current El Niño ”tool kit” is admittedly impressive [6].
@@
 The real scientific holy grail is the DCM.
-Note added to the proof
-We arrive at a "Mexican standoff" because only future data can confirm that our “Big Wave” forecasts are correct. Fortunately, one minor detail saves us one year of waiting.
+Appendix 2
+The analysed data in Appendix 1 were the yearly values before December 2024.
+Future data
+We write this Appendix 2 because our research has arrived at a "Mexican standoff, where only future data can confirm our “Big Wave” forecasts.
+Fortunately, one minor detail saves us one year of waiting.
+Year 2025
 On January 29th, 2026, we computed the yearly values from the NOAA1 data.
-Four monthly values for the year 2025 were missing and our analysed data ended in the year 2024.
-Today February 25th 2026, NOAA1 still has these same four missing monthly values for the year 2025 (“-99.99”).
-To get these missing values, we are advised to use the Climate Prediction Center (CPC) data in NOAA2.666https://www.cpc-ncep.noaa.gov/products/
+These yearly data ended in December 2024 because the four last monthly values for the year 2025 were missing.
+On February 25th, 2026, NOAA1 still had these same four missing monthly values for the year 2025 (“-99.99”).
+To get these missing monthly values, we were advised to use the Climate Prediction Center (CPC) data in NOAA2.666https://www.cpc-ncep.noaa.gov/products/
 analysis-monitoring/ensostuff/ON-v5.php
-All twelve monthly values for the year 2025 are given in Table S7 (column 2).
-These values give $y_{i}\pm 1\sigma_{i}=-0.26\pm 0.21$ $\mathrm{C}^{o}$.
-A red circle now highlights this observed new 2025 year value which is added into Figures S3e and Figures S4e.
+All twelve monthly values for the year 2025 are given in Table S7 (column 3).
+These monthly values give the 2025 yearly value $y_{i}\pm 1\sigma_{i}=-0.26\pm 0.21$ $\mathrm{C}^{o}$.
+A red circle now highlights this observed new 2025 yearly value in Figures S3e and S4e.
 Our “Big Wave” forecast for the year 2025 succeeds.
-The year 2025 forecast for the best $\mathcal{M}$=3 model for all data (Table S3) is more accurate than the forecast for the unstable $\mathcal{M}$=4 model.
-Our forecast for this 2025 observation can, of course, be dismissed just as one lucky guess.
+This 2025 forecast for the best all yearly data $\mathcal{M}$=3 model (Table S3) is more accurate than the forecast for the unstable $\mathcal{M}$=4 model (Table S3).
+Our forecast for this new 2025 observation can, of course, be dismissed just as one lucky guess.
 Who wants to take the risk that our severe “Big Wave” forecasts for the years 2030 and 2031 are also “lucky”?
-We also analyse the monthly SST values (Tables S8-S19, Figures S5-S16).
-The red dots in our figures highlight the available new monthly values between January 2025 and February 2026.
-Most of those red dots hit their mark.
-We consistently detect the “new” 3.6 and 4.6 year periods from these monthly values.
-The monthly mean computation eliminates the detection of these periods from yearly data.
-We conclude that the chaotic $\pm 1$$\mathrm{C}^{o}$ weather changes can prevent accurate DCM monthly forecasts.
-However, we can compute the “Big Wave” monthly values (Table S20).
-We admit that this does not fully solve the El Niño riddle but it may help.
+All months
+We did not analyse the monthly data in Appendix 1.
+The analysed yearly data ended in December 2024.
+Here, we analyse the NOAA1 monthly data ending in December 2024.
+This ensures that the same numbers are used to compute the yearly and the monthly data values.
+The computation of monthly means eliminated the detection of shorter periods from yearly data.
+Therefore we lower the $P_{\mathrm{min}}$ value from 3 to 2 years when analysing these monthly data.
+The value $P_{\mathrm{min}}=100$ years remains the same.
+The DCM analysis results for all monthly values between January 1870 and December 2024 are given in Table S8.
+The four signal $\mathcal{M}$=4 model is the best.
+The signal periods are $P_{1}=3.643\pm 0.034$ years, $P_{2}=4.7475\pm 0.0082$ years $P_{3}=5.649\pm 0.012$ years and $P_{4}=12.706\pm 0.050$ years.
+The 12.7 year signal has the highest amplitude.
+The periodogram shows four clear minima (Figure S5a).
+The 1st signal significance is $Q_{\mathrm{F}}<\!10^{-16}$ (Table S9).
+The 2nd, 3rd, and 4th signal significances are also $Q_{\mathrm{F}}<\!10^{-16}$ (Table S8).
+These signal detections are absolutely certain.
+The DCM search for the fifth signal from the $\mathcal{M}$=4 model residuals gives $R=528.7$.
+Hence, these monthly data contain only four signals because the $\mathcal{M}$=4 model has $R=475.6$.
+The best DCM model for all monthly data is shown in Figures S5b-e.
+The residuals show strong correlation, which is typical for climatological time series [64, 17].
+Due to this correlation, the bootstrap procedure gives too small model error $\sigma_{g(t)}$ estimates (Equation 30).
+Therefore, only the next decade monthly forecast values, but not their error estimates, are given in Table S10.
+Separate months
+We also analyse all months separately because that may eliminate the residual correlation problem.
+The results are given in Tables S11-S22.
+The forecasted monthly values are compared to the NOAA2 values (Table S7, column 3), which are highlighted as red dots in Figures S6-S17.
+Except for November 2025, all other red dots are within the forecast error $\pm 3\sigma_{g(t)}$ red dotted lines.
+For separate months analysis, all next decade forecasted monthly values are given in Table S23.
+Note that this table also gives the forecast error estimates because residual correlation for separate months is weaker, or sometimes even totally absent.
+Connection to solar cycle periods
+[…]
+$P_{\mathrm{Syn}}=[P_{1}^{-1}\pm P_{2}^{-1}]^{-1}$ (A1)
+predicts the synodic periods of these signals given in Table S24.
+For example, the strongest 12.82 years period in El Niño data could arise from combination $P_{1}=11.86/2$ years and $P_{2}=11$ years given in line 15 at column 4.
+We think that it is not necessary to discuss all possible synodic period combinations in this short appendix.
+Physics
+We use the relation
+$F_{\mathrm{Sun}}(t)\longrightarrow T_{\mathrm{Sea}}(t)$ (A2)
+to describe how the solar output $F_{\mathrm{Sun}}]={\mathrm{W}}{\mathrm{m}}^{-2}$ triggers the Pacific Ocean sea surface temperature $T_{\mathrm{Sea}}]={\mathrm{C}}^{\mathrm{o}}$ changes.
+Our symbol $\longrightarrow$” contains all unknown climatological processes.
+The monthly data residuals consistently show how rapidly and strongly the $T_{\mathrm{Sea}}$ values react to the $F_{\mathrm{Sun}}$ peaks and valleys (Figures S5b-e, blue dots).
+The $F_{\mathrm{Sun}}$ and $T_{\mathrm{Sea}}$ extrema coincide.
+The DCM detects the $F_{\mathrm{Sun}}$ signals.
+The solar maximum effect is $A_{1}+A_{2}+A_{3}+A_{4}=1.4$$\mathrm{C}^{o}$ (Table S8).
+However, the chaotic $\longrightarrow$” trigger effect can reach $\pm 1$$\mathrm{C}^{o}$.
+We can not forecast the exact $T_{\mathrm{Sea}}$ values, but we can show that the general weather pattern follows the solar energy output “Big Wave”.
+This does not fully solve the El Niño riddle, but it gives a good starting point for modelling the climatological $\longrightarrow$” effects.
 If no internal mechanism can cause the strictly multi-periodic “Big Wave” in the Pacific Ocean then the Sun is the only possible external cause.
 The planets may indeed cause regular solar cycles [31].
+NOAA and other official agencies report El Niño probabilities like 60%, …, 90%.
+The DCM probability that the four strongest “Big wave” signals are real is larger than $1-10^{-16}$ (Table S8).
 Acknowledgements
@@
 Data availability
-All data files and Python codes are stored to
+All manuscript data files and Python codes are stored to
 https://zenodo.org/records/18726181/
 The manual for repeating all analysis is also published there.
+The newest additional Appendix data files and control files are stored to
+https://zenodo.org/records/19183791/
 Declarations
@@
             (d)             (e)
-Figure S3: “Big Wave” for all yearly mean data model $\mathcal{M}$=3 in Table S3. Red circle is added observed 2025 year value. Otherwise as in Figure S1.
+Figure S3: “Big Wave” for all yearly mean data model $\mathcal{M}$=3 in Table S3. Red circle is added new observed 2025 year value. Otherwise as in Figure S1.
 Refer to caption Refer to caption
@@
 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 299 $1873.8\pm 1.0$ $1870.18\pm 0.99$ $1870.6\pm 0.97$ $1871.4\pm 2.2$ - -
-Table S2: Periods in first half yaerly mean data. Otherwise as in Table S1.
+Table S2: Periods in first half yearly mean data. Otherwise as in Table S1.
 Period analysis Fisher-test
@@
 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$} 300 $1873.8\pm 1.2$ $1870.2\pm 1.1$ $1870.6\pm 0.97$ $1871.3\pm 2.$ - - -
-Table S4: “Big Wave”yearly forecasts (Figures S3d and S4d: open circles).
+Table S4: “Big Wave”yearly forecasts (Figures S3e and S4e: open circles).
 Figure S3 Figure S4
@@
 […]
-Table S7: Year 2025.
-Temperature anomalies $(y)$ of 12 months.
-(1) NOAA1
-(2) NOAA2
-$y$ $y$
-$\mathrm{C}^{o}$ $\mathrm{C}^{o}$
+Table S7: Data.
+(1) Month
+(2) NOAA1 on 28.01.2026. (3) NOAA2 on 13.03.2026.
+Month $y$ $y$
+(-) $\mathrm{C}^{o}$ $\mathrm{C}^{o}$
+(1) (2) (3)
+Jan-25 -0.50 -0.4
+Feb-25 -0.36 -0.2
+Mar-25 -0.13 -0.1
+Apr-25 -0.09 0.0
+May-25 0.02 0.0
+Jun-25 0.14 0.0
+Jul-25 0.05 -0.1
+Aug-25 0.06 -0.3
+Sep-25 -99.99 -0.4
+Oct-25 -99.99 -0.5
+Nov-25 -99.99 -0.6
+Dec-25 -99.99 -0.5
+Jan-26 - -0.4
+Feb-26 - -
+Refer to caption Refer to caption
+            (b)             (c)
+Refer to caption Refer to caption
+            (d)             (e)
+Figure S5: All monthly data model $\mathcal{M}$=4 in Table S8. Red dots are monthly values from Table S7. Otherwise as in Figure S2.
+Table S8: Periods in non-weighted all months data. Otherwise as in Table S1.
+Period analysis Fisher-test
+Data: Original non-weighted data ($n=1860,\Delta T=154.9$: mAll.dat)
+$\mathcal{M}$ $P_{1}$ (y) $P_{2}$ (y) $P_{3}$ (y) $P_{4}$ (y) $\mathcal{M}$=2 $\mathcal{M}$=3 $\mathcal{M}$=4
+$\eta$ (-) $A_{1}$ ($\mathrm{C}^{o}$) $A_{2}$ ($\mathrm{C}^{o}$) $A_{3}$ ($\mathrm{C}^{o}$) $A_{4}$ ($\mathrm{C}^{o}$) $F_{R}$ (-) $F_{R}$ (-) $F_{R}$(-) Control file
+$R$ (-) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $t_{\mathrm{min,1}}$ (y) $Q_{F}$ (-) $Q_{F}$ (-) $Q_{F}$ (-)
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+One signal
+$\mathcal{M}$=1 $12.716\pm 0.047$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.396\pm 0.029$ - - - 32.9 33.2 35.3 mAllR14K111.dat
+556.8 $1871.62\pm 0.37$ - - - $Q_{\mathrm{F}}<\!10^{-16}$ $Q_{\mathrm{F}}<\!10^{-16}$ $Q_{\mathrm{F}}<\!10^{-16}$
+Two signals
+$\mathcal{M}$=2 $5.651\pm 0.012$ $12.700\pm 0.042$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.348\pm 0.029$ $0.389\pm 0.028$ - - - 31.8 34.7 mAllR14K211.dat
+528.6 $1870.75\pm 0.19$ $1871.76\pm 0.32$ - - - $Q_{\mathrm{F}}<\!10^{-16}$ $Q_{\mathrm{F}}<\!10^{-16}$
+Three signals
+$\mathcal{M}$=3 $4.7459\pm 0.0082$ $5.640\pm 0.012$ $12.694\pm 0.043$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.335\pm 0.023$ $0.340\pm 0.034$ $0.388\pm 0.030$ - - - 35.7 mAllR14K311.dat
+502.6 $1870.60\pm 0.13$ $1870.93\pm 0.15$ $1871.83\pm 0.29$ - - - $Q_{\mathrm{F}}<\!10^{-16}$
+Four signals
+$\mathcal{M}$=4 $3.6436\pm 0.0046$ $4.7477\pm 0.0080$ $5.647\pm 0.011$ $12.704\pm 0.063$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.345\pm 0.030$ $0.332\pm 0.031$ $0.342\pm 0.025$ $0.392\pm 0.038$ - - - mAllR14K411.dat
+475.0 $1872.68\pm 0.11$ $1870.55\pm 0.14$ $1870.82\pm 0.19$ $1871.74\pm 0.39.$ - - -
+Table S9: Significance one signal model $\mathcal{M}$=1 in Table S8. Otherwise as in Table 9. 66footnotetext: Note: No test for $\mathcal{M}$=5  $(\eta_{1}=\eta_{2}=5)$.
 (1) (2)
--0.50 -0.4
--0.36 -0.2
--0.13 -0.1
--0.09 0.0
-0.02 0.0
-0.14 0.0
-0.05 -0.1
-0.06 -0.3
--99.99 -0.4
--99.99 -0.5
--99.99 -0.6
--99.99 -0.5
+$g_{1,1,1}$
+Polynomial $\eta=5,R=557.1$
+$\mathcal{M}$=1 $\uparrow$
+$g(t)=p(t,K_{3}=0)$ $F=63.9$
+$\eta=1$, $R=634$ $Q_{\mathrm{F}}<\!10^{-16}$
+$\mathcal{M}$=2 $\uparrow$
+$g(t)=p(t,K_{3}=1)$ $F=39.8$
+$\eta=2$, $R=593$ $Q_{\mathrm{F}}<\!10^{-16}$
+$\mathcal{M}$=3 $\uparrow$
+$g(t)=p(t,K_{3}=2)$ $F=59.7$
+$\eta=3$, $R=593$ $Q_{\mathrm{F}}<\!10^{-16}$
+$\mathcal{M}$=4 $\uparrow$
+$g(t)=p(t,K_{3}=3)$ $F=92.8$
+$\eta=4$, $R=585$ $Q_{\mathrm{F}}<\!10^{-16}$
+$\mathcal{M}$=5 $\uparrow$
+$g(t)=p(t,K_{3}=4)$ No test11footnotemark: 1
+$\eta=5$, $R=584$ $Q_{\mathrm{F}}=1$
+$\mathcal{M}$=6 $\uparrow$
+$g(t)=p(t,K_{3}=5)$ F=34F=-34
+$\eta=6$, $R=584$ $Q_{\mathrm{F}}=1$
+$\mathcal{M}$=7 $\uparrow$
+$g(t)=p(t,K_{3}=6)$ F=33.5F=-33.5
+$\eta=7$, $R=578$ $Q_{\mathrm{F}}=1$
+$\mathcal{M}$=8 $\uparrow$
+$g(t)=p(t,K_{3}=7)$ F=22.3F=-22.3
+$\eta=8$, $R=578$ $Q_{\mathrm{F}}=1$
+[…]
+Table S10: All months. “Big Wave” forecast (Figure S5). Symbols $\uparrow$ and $\downarrow$ tell that preceding value is smaller or bigger.
+Year Jan Feb Mar Apr May Jun Jul Aug Sep Oct Nov Dec
+2025 0.06 $\uparrow$ 0.05 $\downarrow$ 0.04 $\downarrow$ 0.03 $\downarrow$ 0.02 $\downarrow$ 0.02 $\downarrow$ 0.01 $\downarrow$ 0.00 $\downarrow$ -0.00 $\downarrow$ -0.01 $\downarrow$ -0.01 $\downarrow$ -0.01 $\uparrow$
+2026 -0.00 $\uparrow$ 0.00 $\uparrow$ 0.01 $\uparrow$ 0.01 $\uparrow$ 0.02 $\uparrow$ 0.03 $\uparrow$ 0.04 $\uparrow$ 0.05 $\uparrow$ 0.06 $\uparrow$ 0.08 $\uparrow$ 0.09 $\uparrow$ 0.10 $\uparrow$
+2027 0.11 $\uparrow$ 0.12 $\uparrow$ 0.12 $\uparrow$ 0.13 $\uparrow$ 0.13 $\uparrow$ 0.13 $\uparrow$ 0.13 $\uparrow$ 0.13 $\downarrow$ 0.13 $\downarrow$ 0.12 $\downarrow$ 0.12 $\downarrow$ 0.11 $\downarrow$
+2028 0.10 $\downarrow$ 0.09 $\downarrow$ 0.09 $\downarrow$ 0.08 $\downarrow$ 0.07 $\downarrow$ 0.06 $\downarrow$ 0.05 $\downarrow$ 0.05 $\downarrow$ 0.04 $\downarrow$ 0.04 $\downarrow$ 0.04 $\uparrow$ 0.04 $\uparrow$
+2029 0.05 $\uparrow$ 0.06 $\uparrow$ 0.07 $\uparrow$ 0.08 $\uparrow$ 0.10 $\uparrow$ 0.12 $\uparrow$ 0.14 $\uparrow$ 0.17 $\uparrow$ 0.19 $\uparrow$ 0.22 $\uparrow$ 0.25 $\uparrow$ 0.28 $\uparrow$
+2030 0.31 $\uparrow$ 0.34 $\uparrow$ 0.37 $\uparrow$ 0.40 $\uparrow$ 0.42 $\uparrow$ 0.45 $\uparrow$ 0.47 $\uparrow$ 0.49 $\uparrow$ 0.50 $\uparrow$ 0.52 $\uparrow$ 0.52 $\uparrow$ 0.53 $\uparrow$
+2031 0.53 $\uparrow$ 0.52 $\downarrow$ 0.52 $\downarrow$ 0.50 $\downarrow$ 0.49 $\downarrow$ 0.47 $\downarrow$ 0.45 $\downarrow$ 0.42 $\downarrow$ 0.39 $\downarrow$ 0.36 $\downarrow$ 0.33 $\downarrow$ 0.30 $\downarrow$
+2032 0.27 $\downarrow$ 0.23 $\downarrow$ 0.20 $\downarrow$ 0.17 $\downarrow$ 0.14 $\downarrow$ 0.12 $\downarrow$ 0.09 $\downarrow$ 0.07 $\downarrow$ 0.05 $\downarrow$ 0.03 $\downarrow$ 0.02 $\downarrow$ 0.01 $\downarrow$
+2033 0.01 $\downarrow$ 0.01 $\downarrow$ 0.01 $\uparrow$ 0.02 $\uparrow$ 0.02 $\uparrow$ 0.03 $\uparrow$ 0.05 $\uparrow$ 0.06 $\uparrow$ 0.08 $\uparrow$ 0.09 $\uparrow$ 0.11 $\uparrow$ 0.13 $\uparrow$
+2034 0.14 $\uparrow$ 0.15 $\uparrow$ 0.17 $\uparrow$ 0.17 $\uparrow$ 0.18 $\uparrow$ 0.19 $\uparrow$ 0.19 $\uparrow$ 0.18 $\downarrow$ 0.18 $\downarrow$ 0.16 $\downarrow$ 0.15 $\downarrow$ 0.13 $\downarrow$
+2035 0.11 $\downarrow$ 0.09 $\downarrow$ 0.06 $\downarrow$ 0.03 $\downarrow$ 0.00 $\downarrow$ -0.03 $\downarrow$ -0.06 $\downarrow$ -0.10 $\downarrow$ -0.13 $\downarrow$ -0.16 $\downarrow$ -0.19 $\downarrow$ -0.22 $\downarrow$
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S5: January data model $\mathcal{M}$=4 in Table S8. Red dots are -0.76 and -0.54 for 2025 and 2026. Otherwise as in Figure S2.
-Table S8: Periods in January data. Note that we the pre-assigned sigficance level $\gamma=0.05$ (Equation 48) because monthly SST data are so noisy. Otherwise as in Table S1.
+Figure S6: January data model $\mathcal{M}$=4 in Table S11. Red dots from Table S7 (Column 3). Otherwise as in Figure S2.
+Table S11: Periods in January data. Note that we use the pre-assigned sigficance level $\gamma=0.05$ (Equation 48) because data of separate months data are so noisy. Otherwise as in Table S1.
 Period analysis Fisher-test
@@
 $\mathcal{M}$1,1,1,R 5 $0.46\pm 0.12$ - - - 3.33 3.40 3.26 JanR14K111.dat
-64.0 $1872.14\pm 0.45$ - - - 0.021 0.0035 0.0012
+64.0 $1872.18\pm 0.46$ - - - 0.021 0.0035 0.0012
 Two signals
@@
 $\mathcal{M}$2,1,1,R 8 $0.45\pm 0.11$ $0.46\pm 0.13$ - - - 3.14 3.09 JanR14K211.dat
-59.9 $1872.14\pm 0.41$ $1870.98\pm 0.59$ - - - 0.027 0.0071
+59.9 $1872.18\pm 0.44$ $1871.02\pm 0.59$ - - - 0.027 0.0071
 Three signals
@@
 $\mathcal{M}$3,1,1,R 11 $0.45\pm 0.10$ $6\pm 0.11 0.44±0$$4\pm 0.11$ - - - 2.91 JanR14K311.dat
-56.2 $1872.15\pm 0.43$ $1870.98\pm 0.43$ $1871.6\pm 1.2$ - - - 0.037
+56.2 $1872.19\pm 0.44$ $1871.02\pm 0.57$ $1871.7\pm 1.4$ - - - 0.037
 Four signals
-$\mathcal{M}$=4 $3.765\pm 0.018$ $4.737\pm 0.048$ $5.643\pm 0.15$ $12.68\pm 0.30$ - - -
+$\mathcal{M}$=4 $3.768\pm 0.018$ $4.737\pm 0.048$ $5.643\pm 0.15$ $12.68\pm 0.30$ - - -
 $\mathcal{M}$4,1,1,R 14 $5\pm 0.12 0.44±0$$4\pm 0.12$ $0.41\pm 0.10$ $0.43\pm 0.11$ - - - JanR14K411.dat
-52.9 $1872.17\pm 0.45$ $1870.90\pm 0.67$ $1870.80\pm 0.66$ $1871.9\pm 1.4.$ - - -
+52.9 $1872.21\pm 0.44$ $1870.94\pm 0.62$ $1870.84\pm 0.73$ $1871.9\pm 1.4.$ - - -
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S6: February data model $\mathcal{M}$=4 in Table S9. Red dots are -0.35 and -0.40 for 2025 and 2026. Otherwise as in Figure S3
+Figure S7: February data model $\mathcal{M}$=4 in Table S12. Otherwise as in Figure S6.
 Table S12: Periods in February data. Otherwise as in Table S1.
@@
 $\mathcal{M}$1,1,1,R 5 $0.48\pm 0.13$ - - - 3.55 3.36 3.23 FebR14K111.dat
-55.8 $1870.60\pm 0.76$ - - - 0.016 0.0039 0.0013
+55.8 $1870.73\pm 0.46$ - - - 0.016 0.0039 0.0013
 Two signals
 $\mathcal{M}$=2 $5.656\pm 0.025$ $12.74\pm 0.22$ - - - $\uparrow$ $\uparrow$
-$\mathcal{M}$2,1,1,R 8 $0.47\pm 0.12$ $0.442\pm 0.086$ - - - 3.02 2.92 FebR14K211.dat
-52.0 $1870.55\pm 0.40$ $1871.2\pm 1.2$ - - - 0.032 0.010
+$\mathcal{M}$2,1,1,R 8 $0.47\pm 0.12$ $0.441\pm 0.086$ - - - 3.02 2.92 FebR14K211.dat
+52.0 $1870.68\pm 0.75$ $1871.3\pm 1.6$ - - - 0.032 0.010
 Three signals
-$\mathcal{M}$=3 $4.7314PM0.050$ $5.672\pm 0.032$ $12.68\pm 0.44$ - - - $\uparrow$
-$\mathcal{M}$3,1,1,R 11 $0.40\pm 0.11$ $0.450\pm 0.098$ $0.42\pm 0.10$ - - - 2.72 FebR14K311.dat
-48.9 $1870.840\pm 0.69$ $1870.37\pm 0.54$ $1871.6\pm 1.6$ - - - 0.046
+$\mathcal{M}$=3 $4.741\pm 0.052$ $5.644\pm 0.089$ $12.73\pm 0.69$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.40\pm 0.11$ $0.45\pm 0.11$ $0.43\pm 0.101$ - - - 2.72 FebR14K311.dat
+48.9 $1870.960\pm 0.60$ $1870.88\pm 0.48$ $1871.37\pm 1.8$ - - - 0.046
 Four signals
-$\mathcal{M}$=4 $3.771\pm 0.035$ $4.741\pm 0.058$ $5.647\pm 0.031$ $12.73\pm 0.22$ - - -
-$\mathcal{M}$4,1,1,R 14 $0.3\pm 1.8$ $0.40\pm 0.10$ $0.45\pm 0.13$ $0.436\pm 0.094$ - - - FebR14K411.dat
-46.2 $1873.8\pm 1.2$ $1870.8\pm 0.53$ $1870.68\pm 0.040$ $1871.32\pm 1.7$ - - -
+$\mathcal{M}$=4 $3.770\pm 0.034$ $4.742\pm 0.046$ $5.647\pm 0.031$ $12.73\pm 0.22$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.37\pm 0.11$ $0.398\pm 0.097$ $0.45\pm 0.13$ $0.436\pm 0.094$ - - - FebR14K411.dat
+46.2 $1872.12\pm 0.44$ $1870.93\pm 0.67$ $1870.80\pm 0.47$ $1871.31\pm 1.5$ - - -
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S7: March data model $\mathcal{M}$=3 in Table S10. Red dot is -0.01 for 2025. Otherwise as in Figure S3
+Figure S8: March data model $\mathcal{M}$=3 in Table S13. Otherwise as in Figure S6.
 Table S13: Periods in March data. Otherwise as in Table S1.
@@
 $\mathcal{M}$=1 $5.665\pm 0.029$ - - - $\uparrow$ $\uparrow$ $\uparrow$
-$\mathcal{M}$1,1,1,R 5 $0.451\pm 0.086$ - - - 3.94 4.27 3.71 MarR14K111.dat
-44.1 $1870.46\pm 0.45$ - - - 0.0097 0.00054 0.00033
+$\mathcal{M}$1,1,1,R 5 $0.45\pm 0.13$ - - - 3.94 4.27 3.71 MarR14K111.dat
+44.1 $1870.67\pm 0.64$ - - - 0.0097 0.00054 0.00033
 Two signals
 $\mathcal{M}$=2 $5.669\pm 0.042$ $12.68\pm 0.56$ - - - $\uparrow$ $\uparrow$
-$\mathcal{M}$2,1,1,R 8 $0.44\pm 0.11$ $0.412\pm 0.085$ - - - 4.33 3.41 MarR14K211.dat
-40.8 $1870.41\pm 0.72$ $1871.7\pm 1.6$ - - - 0.0059 0.0036
+$\mathcal{M}$2,1,1,R 8 $0.444\pm 0.092$ $0.412\pm 0.085$ - - - 4.33 3.41 MarR14K211.dat
+40.8 $1870.61\pm 0.53$ $1871.9\pm 1.1$ - - - 0.0059 0.0036
 Three signals
-$\mathcal{M}$=3 $3.633\pm 0.019$ $5.644\pm 0.052$ $12.73\pm 0.44$ - - - $\leftarrow$
-$\mathcal{M}$3,1,1,R 11 $0.418\pm 0.090$ $0.45\pm 0.10$ $0.43\pm 0.13$ - - - 2.36 MarR14K311.dat
-37.4 $1872.96\pm 0.37$ $1870.74\pm 0.81$ $1871.2\pm 1.2$ - - - 0.074
+$\mathcal{M}$=3 $3.633\pm 0.019$ $5.671\pm 0.0289$ $12.68\pm 0.13$ - - - $\leftarrow$
+$\mathcal{M}$3,1,1,R 11 $0.42\pm 0.12$ $0.45\pm 0.10$ $0.418\pm 0.090$ - - - 2.36 MarR14K311.dat
+37.4 $1873.16\pm 0.26$ $1870.58\pm 0.50$ $1871.85\pm 0.77$ - - - 0.074
 Four signals
@@
 $\mathcal{M}$4,1,1,R 14 $0.419\pm 0.093$ $0.311\pm 0.084$ $0.45\pm 0.10$ $0.418\pm 0.080$ - - - MarR14K411.dat
-35.6 $1872.94\pm 0.37$ $1870.30\pm 0.92$ $1870.42\pm 0.51$ $1871.7\pm 1.1$ - - -
+35.6 $1873.14\pm 0.17$ $1870.51\pm 0.48$ $1870.63\pm 0.65$ $1871.9\pm 1.5$ - - -
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S8: April data model $\mathcal{M}$=4 in Table S11. Red dot is -0.20 for 2025. Otherwise as in Figure S3
+Figure S9: April data model $\mathcal{M}$=4 in Table S14. Otherwise as in Figure S6.
 Table S14: Periods in April data. Otherwise as in Table S1.
@@
 One signal
-$\mathcal{M}$=1 $12.68\pm 0.13$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$=1 $12.69\pm 0.13$ - - - $\uparrow$ $\uparrow$ $\uparrow$
 $\mathcal{M}$1,1,1,R 5 $0.42\pm 0.10$ - - - 5.06 4.29 3.92 AprR14K111.dat
-30.8 $1871.6\pm 1.0$ - - - 0.0023 0.00052 0.00018
+30.8 $1871.90\pm 0.92$ - - - 0.0023 0.00052 0.00018
 Two signals
@@
 $\mathcal{M}$2,1,1,R 8 $0.388\pm 0.073$ $0.416\pm 0.068$ - - - 3.29 3.13 AprR14K211.dat
-27.9 $1870.44\pm 0.62$ $1871.8\pm 1.0$ - - - 0.022 0.0065
+27.9 $1870.73\pm 0.44$ $1872.1\pm 1.3$ - - - 0.022 0.0065
 Three signals
-$\mathcal{M}$=3 $5.665\pm 0.017$ $12.73\pm 0.13$ $18.84\pm 0.84$ - - - $\uparrow$
-$\mathcal{M}$3,1,1,R 11 $0.382\pm 0.097$ $0.416\pm 0.068$ $0.308\pm 0.065$ - - - 2.82 AprR14K311.dat
-26.1 $1870.42\pm 0.29$ $1871.36\pm 0.85$ $1873.7\pm 2.0$ - - - 0.040
+$\mathcal{M}$=3 $5.665\pm 0.017$ $12.73\pm 0.13$ $18.83\pm 0.84$ - - - $\uparrow$
+$\mathcal{M}$3,1,1,R 11 $0.382\pm 0.097$ $0.416\pm 0.068$ $0.31\pm 0.11$ - - - 2.82 AprR14K311.dat
+26.1 $1870.71\pm 0.48$ $1871.65\pm 0.89$ $1874.0\pm 1.6$ - - - 0.040
 Four signals
-$\mathcal{M}$=4 $3.760\pm 0.039$ $5.669\pm 0.069$ $12.74\pm 0.12$ $18.87\pm 0.37$ - - -
+$\mathcal{M}$=4 $3.760\pm 0.039$ $5.669\pm 0.069$ $12.74\pm 0.12$ $18.85\pm 0.37$ - - -
 $\mathcal{M}$4,1,1,R 14 $0.276\pm 0.074$ $0.382\pm 0.096$ $0.42\pm 0.13$ $0.310\pm 0.090$ - - - AprR14K411.dat
-24.6 $1871.98\pm 0.44$ $1870.36\pm 0.47$ $1871.30\pm 0.84$ $1873.6\pm 2.2$ - - -
+24.6 $1872.27\pm 0.38$ $1870.65\pm 0.56$ $1871.59\pm 0.89$ $1873.9\pm 1.6$ - - -
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S9: May data model $\mathcal{M}$=4 in Table S12. Red dot is -0.04 for 2025. Otherwise as in Figure S3
+Figure S10: May data model $\mathcal{M}$=4 in Table S15. Otherwise as in Figure S6.
 Table S15: Periods in May data. Otherwise as in Table S1.
@@
 $\mathcal{M}$=1 $12.67\pm 0.24$ - - - $\uparrow$ $\uparrow$ $\uparrow$
-$\mathcal{M}$1,1,1,R 5 $0.36\pm 0.13$ - - - 5.67 3.39 3.38 MayR14K111.dat
-29.7 $1872.0\pm 1.7$ - - - 0.0010 0.0037 0.00088
+$\mathcal{M}$1,1,1,R 5 $0.358\pm 0.070$ - - - 5.67 3.39 3.38 MayR14K111.dat
+29.7 $1872.4\pm 1.2$ - - - 0.0010 0.0037 0.00088
 Two signals
@@
 $\mathcal{M}$2,1,1,R 8 $0.329\pm 0.073$ $0.353\pm 0.075$ - - - 2.93 3.37 MayR14K211.dat
-27.6 $1870.38\pm 0.54$ $1872.1\pm 1.2$ - - - 0.036 0.0060
+27.6 $1870.76\pm 0.56$ $1872.51\pm 0.96$ - - - 0.036 0.0060
 Three signals
@@
 $\mathcal{M}$3,1,1,R 11 $0.288\pm 0.090$ $0.330\pm 0.057$ $0.356\pm 0.080$ - - - 3.26 MayR14K311.dat
-26.0 $1873.05\pm 0.51$ $1870.38\pm 0.41$ $1872.18\pm 0.92$ - - - 0.023
+26.0 $1873.42\pm 0.40$ $1870.76\pm 0.66$ $1872.55\pm 0.39$ - - - 0.023
 Four signals
@@
 $\mathcal{M}$4,1,1,R 14 $0.304\pm 0.098$ $0.30\pm 0.10$ $0.332\pm 0.067$ $0.356\pm 0.081$ - - - MayR14K411.dat
-24.3 $1873.15\pm 0.50$ $1871.66\pm 0.54$ $1870.37\pm 0.40$ $1872.14\pm 0.90$ - - -
+24.3 $1873.53\pm 0.33$ $1872.03\pm 0.37$ $1870.7\pm 0.47$ $1872.5\pm 1.4$ - - -
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S10: June data model $\mathcal{M}$=4 in Table S13. Red dot is -0.02 for 2025. Otherwise as in Figure S3
+Figure S11: June data model $\mathcal{M}$=4 in Table S16. Otherwise as in Figure S6.
 Table S16: Periods in June data. Otherwise as in Table S1.
@@
 $\mathcal{M}$1,1,1,R 5 $0.42\pm 0.10$ - - - 4.67 4.07 4.00 JunR14K111.dat
-27.4 $1871.5\pm 1.1$ - - - 0.0038 0.00084 0.00014
+27.4 $1872.00\pm 0.86$ - - - 0.0038 0.00084 0.00014
 Two signals
 $\mathcal{M}$=2 $4.740\pm 0.022$ $12.72\pm 0.14$ - - - $\uparrow$ $\uparrow$
-$\mathcal{M}$2,1,1,R 8 $0.344\pm 0.076$ $0.424\pm 0.082$ - - - 3.26 3.42 JunR14K211.dat
-25.0 $1870.28\pm 0.52$ $1871.60\pm 0.97$ - - - 0.023 0.0035
+$\mathcal{M}$2,1,1,R 8 $0.34\pm 0.10$ $0.42\pm 0.11$ - - - 3.26 3.42 JunR14K211.dat
+25.0 $1870.74\pm 0.33$ $1872.05\pm 0.79$ - - - 0.023 0.0035
 Three signals
@@
 $\mathcal{M}$3,1,1,R 11 $0.342\pm 0.072$ $0.411\pm 0.072$ $0.297\pm 0.068$ - - - 3.42 JunR14K311.dat
-23.4 $1870.27\pm 0.30$ $1871.40\pm 0.92$ $1891.0\pm 2.1$ - - - 0.019
+23.4 $1870.73\pm 0.47$ $1871.9\pm 1.0$ $1891.5\pm 1.6$ - - - 0.019
 Four signals
-$\mathcal{M}$=4 $3.647\pm 0.012$ $4.741PM0.037$ $12.76\pm 0.12$ $22.49\pm 1.7$ - - -
-$\mathcal{M}$4,1,1,R 14 $0.289\pm 0.069$ $0.338\pm 0.094$ $0.415\pm 0.077$ $0.302\pm 0.086$ - - - JunR14K411.dat
-21.8 $1872.07\pm 0.35$ $1870.23\pm 0.65$ $1871.35\pm 0.85$ $1891.0\pm 1.8$ - - -
+$\mathcal{M}$=4 $3.647\pm 0.012$ $4.741\pm 0.037$ $12.76\pm 0.12$ $22.44\pm 1.7$ - - -
+$\mathcal{M}$4,1,1,R 14 $0.289\pm 0.069$ $0.338\pm 0.09$ $0.415\pm 0.077$ $0.302\pm 0.086$ - - - JunR14K411.dat
+21.8 $1872.53\pm 0.35$ $1870.8\pm 0.36$ $1871.81\pm 0.99$ $1891.5\pm 1.8$ - - -
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S11: July data model $\mathcal{M}$=4 in Table S14. Red dot is -0.10 for 2025. Otherwise as in Figure S3
+Figure S12: July data model $\mathcal{M}$=4 in Table S17. Otherwise as in Figure S6.
 Table S17: Periods in July data. Otherwise as in Table S1.
@@
 $\mathcal{M}$1,1,1,R 5 $0.372\pm 0.049$ - - - 4.07 4.23 4.24 JulR14K111.dat
-28.5 $1870.8\pm 1.3$ - - - 0.0082 0.00059 $7.1\times 10^{-5}$
+28.5 $1871.3\pm 1.4$ - - - 0.0082 0.00059 $7.1\times 10^{-5}$
 Two signals
 $\mathcal{M}$=2 $12.77\pm 0.18$ $20.0\pm 1.0$ - - - $\uparrow$ $\uparrow$
-$\mathcal{M}$2,1,1,R 8 $71\pm 0.078 0.330$$30\pm 0.081$ - - - 4.14 4.06 JulR14K211.dat
-26.3 $1870.6\pm 1.5$ $1872.50\pm 2.2$ - - - 0.0076 0.00087
+$\mathcal{M}$2,1,1,R 8 $72\pm 0.093 0.330$$30\pm 0.081$ - - - 4.14 4.06 JulR14K211.dat
+26.3 $1871.1\pm 1.3$ $1873.0\pm 2.1$ - - - 0.0076 0.00087
 Three signals
@@
 $\mathcal{M}$3,1,1,R 11 $0.331\pm 0.085$ $0.369\pm 0.078$ $0.327\pm 0.069$ - - - 3.75 JulR14K311.dat
-24.2 $1870.23\pm 0.44$ $1870.7\pm 1.0$ $1872.7\pm 1.8$ - - - 0.012
+24.2 $1870.78\pm 0.40$ $1871.2\pm 1.3$ $1873.2\pm 1.6$ - - - 0.012
 Four signals
@@
 $\mathcal{M}$4,1,1,R 14 $0.302\pm 0.070$ $0.333\pm 0.08$ $0.370\pm 0.081$ $0.329\pm 0.047$ - - - JulR14K411.dat
-22.4 $1871.68\pm 0.34$ $1870.19\pm 0.42$ $1870.62\pm 0.80$ $1872.5\pm 1.7$ - - -
+22.4 $1872.25\pm 0.34$ $1870.73\pm 0.42$ $1871.2\pm 1.2$ $1873.1\pm 2.1$ - - -
 Refer to caption Refer to caption
@@
 $\mathcal{M}$1,1,1,R 5 $0.38\pm 0.14$ - - - 3.86 37.4 3.56 AugR14K111.dat
-35.4 $1871.87\pm 0.44$ - - - 0.011 0.0017 0.00051
+35.4 $1872.50\pm 0.36$ - - - 0.011 0.0017 0.00051
 Two signals
@@
 $\mathcal{M}$2,1,1,R 8 $0.38\pm 0.10$ $0.368\pm 0.071$ - - - 3.43 3.26 Aug14K211.dat
-32.8 $1871.88\pm 0.26$ $1871.1\pm 1.7$ - - - 0.019 0.0050
+32.8 $1872.51\pm 0.23$ $1871.8\pm 2.7$ - - - 0.019 0.0050
 Three signals
@@
 $\mathcal{M}$3,1,1,R 11 $0.38\pm 0.10$ $0.331\pm 0.081$ $0.367\pm 0.071$ - - - 2.92 AugR14K311.dat
-30.6 $1871.88\pm 0.41$ $1872.4\pm 1.7$ $1871.2\pm 1.9$ - - - 0.036
+30.6 $1872.51\pm 0.32$ $1873.0\pm 1.5$ $1871.9\pm 2.0$ - - - 0.036
 Four signals
@@
 $\mathcal{M}$4,1,1,R 14 $0.382\pm 0.084$ $0.311\pm 0.064$ $0.334\pm 0.084$ $0.353\pm 0.092$ - - - AugR14K411.dat
-28.8 $1871.88\pm 0.26$ $1872.31\pm 0.99$ $1872.9\pm 1.2$ $1871.3\pm 2.4$ - - -
+28.8 $1872.51\pm 0.36$ $1872.94\pm 0.64$ $1873.6\pm 1.1$ $1871.9\pm 1.9$ - - -
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S13: September data model $\mathcal{M}$=3 in Table S16. Red dot is -0.40 for 2025. Otherwise as in Figure S3
+Figure S14: September data model $\mathcal{M}$=3 in Table S19. Otherwise as in Figure S6.
 Table S19: Periods in September data. Otherwise as in Table S1.
@@
 $\mathcal{M}$1,1,1,R 5 $0.350\pm 0.078$ - - - 2.84 2.81 2.78 SepR14K111.dat
-43.5 $1872.5\pm 1.3$ - - - 0.040 0.013 0.0050
+43.5 $1873.23\pm 0.99$ - - - 0.040 0.013 0.0050
 Two signals
-$\mathcal{M}$=2 $3.546\pm 0.022$ $20.10\pm 0.68$ - - - $\uparrow$ $\uparrow$
-$\mathcal{M}$2,1,1,R 8 $0.349\pm 0.087$ $0.3550\pm 0.076$ - - - 2.70 2.66 Sep14K211.dat
-41.1 $1871.60\pm 0.43$ $1871.6\pm 2.6$ - - - 0.048 0.018
+$\mathcal{M}$=2 $3.546\pm 0.022$ $20.1\pm 1.3$ - - - $\uparrow$ $\uparrow$
+$\mathcal{M}$2,1,1,R 8 $0.35\pm 0.13$ $0.350\pm 0.076$ - - - 2.70 2.66 Sep14K211.dat
+41.1 $1872.32\pm 0.34$ $1872.3\pm 2.9$ - - - 0.048 0.018
 Three signals
-$\mathcal{M}$=3 $3.546\pm 0.018$ $9.13\pm 0.21$ $20.0\pm 1.4$ - - - $\leftarrow$
+$\mathcal{M}$=3 $3.546\pm 0.018$ $9.13\pm 0.21$ $19.99\pm 0.89$ - - - $\leftarrow$
 $\mathcal{M}$3,1,1,R 11 $0.35\pm 0.11$ $0.340\pm 0.094$ $0.341\pm 0.083$ - - - 2.53 SepR14K311.dat
-38.9 $1871.61\pm 0.37$ $1872.47\pm 0.96$ $1871.9\pm 3.5$ - - - 0.060
+38.9 $1872.31\pm 0.47$ $1873.18\pm 0.81$ $1872.6\pm 2.3$ - - - 0.060
 Four signals
-$\mathcal{M}$=4 $3.545\pm 0.025$ $4.749\pm 0.071$ $9-12\pm 0.11$ $19.92\pm 0.93$ - - -
+$\mathcal{M}$=4 $3.545\pm 0.025$ $4.749\pm 0.071$ $9.12\pm 0.11$ $19.92\pm 0.93$ - - -
 $\mathcal{M}$4,1,1,R 14 $0.35\pm 0.12$ $0.32\pm 0.11$ $0.343\pm 0.079$ $0.340\pm 0.082$ - - - SepR14K411.dat
-36.9 $1871.64\pm 0.30$ $1874.44\pm 0.63$ $1872.5\pm 1.1$ $1872.0\pm 3.1$ - - -
+36.9 $1872.35\pm 0.27$ $1875.18\pm 0.74$ $1873.2PM1.4$ $1872.7\pm 2.3$ - - -
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S14: October data model $\mathcal{M}$=4 in Table S17. Red dot is -0.48 for 2025. Otherwise as in Figure S3
+Figure S15: October data model $\mathcal{M}$=4 in Table S20. Otherwise as in Figure S6.
 Table S20: Periods in October data. Otherwise as in Table S1.
@@
 $\mathcal{M}$1,1,1,R 5 $0.38\pm 0.13$ - - - 2.51 2.70 2.78 OctR14K111.dat
-56.0 $1871.59\pm 0.53$ - - - 0.039 0.016 0.0049
+56.0 $1872.40\pm 0.54$ - - - 0.039 0.016 0.0049
 Two signals
@@
 $\mathcal{M}$2,1,1,R 8 $0.38\pm 0.11$ $0.40\pm 0.12$ - - - 2.46 2.65 Oct14K211.dat
-52.9 $1871.62\pm 0.41$ $1874.13\pm 0.64$ - - - 0.065 0.018
+52.9 $1871.62\pm 0.41$ $1874.92\pm 0.88$ - - - 0.065 0.018
 Three signals
@@
 $\mathcal{M}$3,1,1,R 11 $0.381\pm 0.093$ $0.419\pm 0.11$ $0.367\pm 0.093$ - - - 2.75 OctR14K311.dat
-50.3 $1871.63\pm 0.52$ $1874.09\pm 0.44$ $1882.6PM1.3$ - - - 0.045
+50.3 $1872.42\pm 0.40$ $1874.88\pm 0.48$ $1883.4\pm 2.0$ - - - 0.045
 Four signals
@@
 $\mathcal{M}$4,1,1,R 14 $0.38\pm 0.10$ $0.40\pm 0.15$ $0.38\pm 0.13$ $0.38\pm 0.11$ - - - OctR14K411.dat
-47.5 $1871.64\pm 0.50$ $1874.07\pm 0.52$ $1882.2\pm 1.1$ $1871.3\pm 3.0$ - - -
+47.5 $1872.44\pm 0.46$ $1874.86\pm 0.47$ $1883.0\pm 1.4$ $1872.0\pm 2.7$ - - -
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S15: Novermber data model $\mathcal{M}$=3 in Table S18. Red dot is -0.80 for 2025. Otherwise as in Figure S3
+Figure S16: November data model $\mathcal{M}$=3 in Table S21. Otherwise as in Figure S6.
 Table S21: Periods in November data. Otherwise as in Table S1.
@@
 $\mathcal{M}$1,1,1,R 5 $0.47\pm 0.10$ - - - 2.86 3.02 3.08 NovR14K111.dat
-64.8 $1875.17\pm 0.75$ - - - 0.039 0.0081 0.0021
+64.8 $1876.05\pm 0.40$ - - - 0.039 0.0081 0.0021
 Two signals
@@
 $\mathcal{M}$2,1,1,R 8 $0.43\pm 0.12$ $0.47\pm 0.14$ - - - 3.07 3.06 Nov14K211.dat
-61.2 $1871.67\pm 0.42$ $1875.13\pm 0.59$ - - - 0.039 0.0076
+61.2 $1872.54\pm 0.48$ $1876.01\pm 0.57$ - - - 0.039 0.0076
 Three signals
@@
 $\mathcal{M}$3,1,1,R 11 $0.44\pm 0.14$ $0.46\pm 0.12$ $0.43\pm 0.13$ - - - 2.67 NovR14K311.dat
-57.5 $1871.68\pm 0.42$ $1875.09\pm 0.57$ $1882.2\pm 2.2$ - - - 0.050
+57.5 $1872.56\pm 0.44$ $187596\pm 0.59$ $1883.1\pm 1.2$ - - - 0.050
 Four signals
-$\mathcal{M}$=4 $3.647\pm 0.029$ $5.67\pm 0.11$ $9.10\pm 0.20$ $12.90\pm 0.31$ - - -
+$\mathcal{M}$=4 $3.647\pm 0.029$ $5.667\pm 0.072$ $9.095\pm 0.090$ $12.90\pm 0.31$ - - -
 $\mathcal{M}$4,1,1,R 14 $0.43\pm 0.13$ $0.47\pm 0.15$ $0.42\pm 0.12$ $0.432\pm 0.096$ - - - NovR14K411.dat
-54.1 $1871.67\pm 0.53$ $1875.14\pm 0.67$ $72.67\pm 0.77 1881$$81.9\pm 1.2$ - - -
+54.1 $1872.55\pm 0.41$ $1876.01\pm 0.67$ $1873.54\pm 0.78$ $1882.8\pm 1.7$ - - -
 Refer to caption Refer to caption
@@
             (c)             (d)
-Figure S16: December data model $\mathcal{M}$=4 in Table S19. Red dot is -0.92 for 2025. Otherwise as in Figure S3
+Figure S17: December data model $\mathcal{M}$=4 in Table S22. Otherwise as in Figure S6.
 Table S22: Periods in December data. Otherwise as in Table S1.
@@
 One signal
-$\mathcal{M}$=1 $5.645\pm 0.033$ - - - $\uparrow$ $\uparrow$ $\uparrow$
-$\mathcal{M}$1,1,1,R 5 $0.48\pm 0.13$ - - - 4.97 4.20 3.94 DecR14K111.dat
-64.8 $1871.70\pm 0.44$ - - - 0.0026 0.00064 0.00017
+$\mathcal{M}$=1 $3.645\pm 0.019$ - - - $\uparrow$ $\uparrow$ $\uparrow$
+$\mathcal{M}$1,1,1,R 5 $0.48\pm 0.12$ - - - 4.97 4.20 3.94 DecR14K111.dat
+64.8 $1872.66\pm 0.45$ - - - 0.0026 0.00064 0.00017
 Two signals
@@
 $\mathcal{M}$2,1,1,R 8 $0.49\pm 0.15$ $0.50\pm 0.16$ - - - 3.20 3.20 Dec14K211.dat
-58.8 $1871.71\pm 0.36$ $1870.3\pm 1.5$ - - - 0.025 0.0056
+58.8 $1872.67\pm 0.36$ $1871.29\pm 0.95$ - - - 0.025 0.0056
 Three signals
@@
 $\mathcal{M}$3,1,1,R 11 $0.49\pm 0.10$ $0.43\pm 0.10$ $0.50\pm 0.11$ - - - 3.06 DecR14K311.dat
-55.1 $1871.73\pm 0.31$ $1875.23\pm 0.67$ $1870.4\pm 1.1$ - - - 0.030
+55.1 $1872.69\pm 0.38$ $1876.19\pm 0.61$ $1871.35\pm 0.98$ - - - 0.030
 Four signals
-$\mathcal{M}$=4 $3.644\pm 0.040$ $5.566\pm 0.020$ $5.664\pm 0.058$ $12.74\pm 0.32$ - - -
+$\mathcal{M}$=4 $3.644\pm 0.040$ $4.566\pm 0.020$ $5.664\pm 0.058$ $12.74\pm 0.32$ - - -
 $\mathcal{M}$4,1,1,R 14 $0.49\pm 0.10$ $0.42\pm 0.10$ $0.428\pm 0.092$ $0.50\pm 0.12$ - - - DecR14K411.dat
-51.7 $1871.76\pm 0.38$ $1874.40\pm 0.38$ $1875.26\pm 0.50$ $1870.4\pm 2.0$ - - -
-Table S20: “Big Wave” monthly forecast 2025-2036.
+51.7 $1872.71\pm 0.38$ $1875.35\pm 0.59$ $1876.22\pm 0.80$ $1871.4\pm 1.2$ - - -
+Table S23: Separate months. “Big Wave” forecast (Figures S6-S17: open circles). Otherwise as in Table S10.
 Year Jan Feb Mar Apr May Jun Jul Aug Sep Oct Nov Dec
 ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$) ($\mathrm{C}^{o}$)
-2025 $0.3\pm 0.2$ $0.4\pm 0.2$ $0.2\pm 0.2$ $0.1\pm 0.2$ $0.2\pm 0.2$ $0.3\pm 0.2$ $2\pm 0.2 0.0$$0\pm 0.2$ $2\pm 0.2 0.0$$\pm 0.3 0.0$$0\pm 0.2$ $0.3\pm 0.2$
-2026 $0.0\pm 0.3$ $1\pm 0.2 0.0$$0\pm 0.2$ $0.1\pm 0.1$ $0.3\pm 0.1$ $0.3\pm 0.2$ $0.1\pm 0.2$ $\pm 0.2 0.2±0.2$$\pm 0.2$ $0.2\pm 0.2$ $0.3\pm 0.3$ $0.2\pm 0.3$
-2027 $0.3\pm 0.2$ $0.2\pm 0.3$ $0.4\pm 0.2$ $0.1\pm 0.2$ $0.4\pm 0.1$ $0.1\pm 0.1$ $0.1\pm 0.2$ $0.4\pm 0.1$ $0.2\pm 0.2$ $0.3\pm 0.2$ $0.3\pm 0.2$ $0.5\pm 0.2$
-2028 $0\pm 0.2 0.0$$0\pm 0.2$ $\pm 0.2 0.2±0.2$$\pm 0.2$ $0.0\pm 0.2$ $0.1\pm 0.2$ $0.1\pm 0.2$ $0.2\pm 0.2$ $\pm 0.2 0.3±0.3$$\pm 0.3$ $1\pm 0.2 0.0$$0\pm 0.3$
-2029 $0.3\pm 0.3$ $0.3\pm 0.2$ $0.2\pm 0.2$ $1\pm 0.1 0.0$$0\pm 0.2$ $0.2\pm 0.2$ $0.5\pm 0.2$ $1\pm 0.2 0.0$$0\pm 0.2$ $3\pm 0.2 0.0$$0\pm 0.3$ $0.1\pm 0.2$
-2030 $0.2\pm 0.2$ $0.3\pm 0.2$ $\pm 0.2 0.1±0.1$$\pm 0.1$ $0.5\pm 0.2$ $0.4\pm 0.2$ $0.3\pm 0.2$ $4\pm 0.2 0.0$$0\pm 0.2$ $0.1\pm 0.2$ $0.6\pm 0.2$ $0.4\pm 0.2$
-2031 $0.3\pm 0.3$ $0.3\pm 0.2$ $0.6\pm 0.2$ $0.6\pm 0.2$ $0.4\pm 0.2$ $0.3\pm 0.1$ $0.1\pm 0.1$ $0.5\pm 0.1$ $1\pm 0.2 0.0$$0\pm 0.2$ $0.7\pm 0.2$ $0.6\pm 0.2$
-2032 $0.4\pm 0.2$ $0.4\pm 0.2$ $0.4\pm 0.2$ $0.7\pm 0.2$ $0.2\pm 0.1$ $1\pm 0.1 0.0$$0\pm 0.2$ $1\pm 0.1 0.0$$0\pm 0.3$ $0.3\pm 0.3$ $0.2\pm 0.3$ $0.3\pm 0.2$
-2033 $0.1\pm 0.3$ $2\pm 0.2 0.0$$0\pm 0.2$ $0.4\pm 0.2$ $0.4\pm 0.2$ $0.3\pm 0.2$ $3\pm 0.1 0.0$$0\pm 0.2$ $0.2\pm 0.2$ $2\pm 0.3 0.0$$\pm 0.3 0.0$$0\pm 0.2$
-2034 $0.2\pm 0.3$ $2\pm 0.3 0.0$$\pm 0.2 0.0$$0\pm 0.1$ $0.2\pm 0.2$ $0.5\pm 0.2$ $0.1\pm 0.2$ $1\pm 0.2 0.0$$0\pm 0.2$ $0.4\pm 0.3$ $0.1\pm 0.3$ $0.2\pm 0.3$
-2035 $0.1\pm 0.3$ $0.1\pm 0.3$ $\pm 0.3 0.1±0.1$$\pm 0.1$ $0.4\pm 0.2$ $0.3\pm 0.2$ $\pm 0.1 0.2±0.2$-$\pm 0.2$ $0.2\pm 0.3$ $0.5\pm 0.3$ $0.1\pm 0.2$ $0.4\pm 0.3$
-2036 $0\pm 0.3 0.0$$0\pm 0.2$ $0.3\pm 0.3$ $0.1\pm 0.2$ $0.1\pm 0.2$ $0.2\pm 0.1$ $0.2\pm 0.1$ $5\pm 0.1 0.0$$\pm 0.2 0.0$$0\pm 0.2$ $0.1\pm 0.2$ $0.1\pm 0.2$
+2025 $0.3\pm 0.3$ $\uparrow$ $0.4\pm 0.3$ $\uparrow$ $0.2\pm 0.2 \downarrow$ $0.1\pm 0.2 \downarrow$ $0.2\pm 0.2 \downarrow$ $0.3\pm 0.1 \downarrow$ $2\pm 0.2 \uparrow 0.0$$0\pm 0.2 \downarrow$ $2\pm 0.2 \uparrow 0.0$$\pm 0.2 \downarrow 0.0$$0\pm 0.2 \downarrow$ $0.3\pm 0.2 \downarrow$
+2026 $0.0\pm 0.2 \uparrow$ $1\pm 0.2 \uparrow 0.0$$0\pm 0.2 \uparrow$ $0.1\pm 0.1 \downarrow$ $0.3\pm 0.2$ $\uparrow$ $0.3\pm 0.1 \downarrow$ $0.1\pm 0.2$ $\uparrow$ $\pm 0.2 \uparrow 0.2±0.2$$\pm 0.2$ $\uparrow$ $0.2\pm 0.2$ $\uparrow$ $0.3\pm 0.2$ $\uparrow$ $0.2\pm 0.2$ $\downarrow$
+2027 $0.3\pm 0.3 \downarrow$ $0.2\pm 0.2 \downarrow$ $0.4\pm 0.2$ $\uparrow$ $0.1\pm 0.2$ $\uparrow$ $0.4\pm 0.2$ $\uparrow$ $0.1\pm 0.1 \downarrow$ $0.1\pm 0.2 \uparrow$ $0.4\pm 0.2$ $\uparrow$ $0.2\pm 0.2 \downarrow$ $0.3\pm 0.3$ $\uparrow$ $0.3\pm 0.2$ $\uparrow$ $0.5\pm 0.2$ $\uparrow$
+2028 $0.1\pm 0.4$ $\downarrow$ $0.1\pm 0.2$ $\downarrow$ $0.2\pm 0.2$ $\uparrow$ $2\pm 0.1 \uparrow 0.0$$0\pm 0.2 \downarrow$ $\pm 0.2 \uparrow 0.1±0.1$$\pm 0.1$ $\uparrow$ $0.2\pm 0.2$ $\uparrow$ $0.3\pm 0.2 \downarrow$ $0.3\pm 0.2$ $\uparrow$ $1\pm 0.3 \downarrow 0.0$$0\pm 0.2 \uparrow$
+2029 $0.3\pm 0.2$ $\uparrow$ $0.3\pm 0.2$ $\uparrow$ $0.2\pm 0.2 \downarrow$ $1\pm 0.2 \downarrow 0.0$$0\pm 0.2 \downarrow$ $0.2\pm 0.2$ $\uparrow$ $0.5\pm 0.1$ $\uparrow$ $1\pm 0.1 \downarrow 0.0$$0\pm 0.2 \downarrow$ $3\pm 0.3 \uparrow 0.0$$0\pm 0.2 \downarrow$ $0.1\pm 0.3 \downarrow$
+2030 $0.2\pm 0.3$ $\uparrow$ $0.3\pm 0.2$ $\uparrow$ $0.1\pm 0.2$ $\downarrow$ $0.1\pm 0.2$ $\downarrow$ $0.5\pm 0.2$ $\uparrow$ $0.4\pm 0.1$ $\downarrow$ $0.3\pm 0.2$ $\downarrow$ $4\pm 0.2 \uparrow 0.0$$0\pm 0.2 \downarrow$ $0.1\pm 0.2$ $\uparrow$ $0.6\pm 0.2$ $\uparrow$ $0.4\pm 0.2$ $\downarrow$
+2031 $0.3\pm 0.3$ $\downarrow$ $0.4\pm 0.2$ $\downarrow$ $0.6\pm 0.1$ $\uparrow$ $0.6\pm 0.2$ $\uparrow$ $0.4\pm 0.1$ $\downarrow$ $0.3\pm 0.1$ $\downarrow$ $0.1\pm 0.1 \downarrow$ $0.5\pm 0.2$ $\uparrow$ $1\pm 0.2 \downarrow 0.0$$0\pm 0.2 \uparrow$ $0.7\pm 0.1$ $\uparrow$ $0.6\pm 0.3$ $\downarrow$
+2032 $0.4\pm 0.2$ $\downarrow$ $0.4\pm 0.2$ $\uparrow$ $0.4\pm 0.2$ $\uparrow$ $0.7\pm 0.2$ $\uparrow$ $0.2\pm 0.2$ $\downarrow$ $1\pm 0.2 \downarrow 0.0$$0\pm 0.1 \downarrow$ $1\pm 0.2 \uparrow 0.0$$0\pm 0.2 \downarrow$ $0.3\pm 0.2$ $\uparrow$ $0.2\pm 0.2$ $\downarrow$ $0.3\pm 0.2$ $\uparrow$
+2033 $0.1\pm 0.3$ $\downarrow$ $1\pm 0.2 \uparrow 0.0$$0\pm 0.2 \downarrow$ $0.4\pm 0.2$ $\uparrow$ $0.4\pm 0.2$ $\downarrow$ $0.3\pm 0.2$ $\downarrow$ $3\pm 0.1 \downarrow 0.0$$0\pm 0.2 \downarrow$ $0.2\pm 0.2$ $\uparrow$ $2\pm 0.3 \downarrow 0.0$$\pm 0.3 \downarrow 0.0$$0\pm 0.3 \downarrow$
+2034 $\pm 0.3 \downarrow 0.2±0.2$-$\pm 0.2$ $\downarrow$ $0\pm 0.2 \uparrow 0.0$$0\pm 0.2 \uparrow$ $0.2\pm 0.2$ $\uparrow$ $0.5\pm 0.2$ $\uparrow$ $0.1\pm 0.2$ $\downarrow$ $1\pm 0.2 \uparrow 0.0$$0\pm 0.2 \downarrow$ $0.4\pm 0.3 \downarrow$ $0.1\pm 0.2$ $\uparrow$ $0.2\pm 0.2 \downarrow$
+2035 $0.1\pm 0.3 \uparrow$ $0.1\pm 0.2 \uparrow$ $0.1\pm 0.2 \downarrow$ $0.1\pm 0.2$ $\uparrow$ $0.4\pm 0.2 \downarrow$ $0.3\pm 0.2$ $\uparrow$ $\pm 0.2 \downarrow 0.2±0.2$-$\pm 0.2$ $\uparrow$ $0.2\pm 0.2 \uparrow$ $0.5\pm 0.2 \downarrow$ $0.1\pm 0.2 \uparrow$ $0.4\pm 0.2 \downarrow$
+2036 $0\pm 0.3 \uparrow 0.0$$0\pm 0.3 \uparrow$ $0.3\pm 0.2 \downarrow$ $0.1\pm 0.2$ $\uparrow$ $0.1\pm 0.2 \downarrow$ $\pm 0.2 \downarrow 0.2±0.2$-$\pm 0.2$ $\uparrow$ $5\pm 0.1 \downarrow 0.0$$\pm 0.2 \uparrow 0.0$$0\pm 0.3 \uparrow$ $0.1\pm 0.2 \downarrow$ $0.1\pm 0.2 \uparrow$
+[…]
+Line $P_{1}$ $P_{2}=10$ $P_{2}=11$ $P_{2}=11.86$
+(-) (y) (y) (y) (y)
+(1) (2) (3) (4) (5)
+(1) 10 - 110 63.8
+(2) - 5.23 5.42
+(3) 10/2 - 9.16 8.64
+(4) - 3.43 3.51
+(5) 10/3 - 4.78 4.63
+(6) - 2.55 2.60
+(7) 11 110 - 152
+(8) 5.23 - 5.70
+(9) 11/2 12.22 - 10.25
+(10) 3.55 - 3.75
+(11) 11/3 5.79 - 5.31
+(12) 2.68 - 2.80
+(13) 11.86 63.8 152 -
+(14) 5.42 5.71 -
+(15) 11.86/2 14.57 12.86 -
+(16) 3.72 3.85 -
+(17) 11.86/3 6.53 6.17 -
+(18) 2.83 2.90 -